### v1

# Two-Player Zero-Sum Games with Bandit Feedback

Abstract
We study a two-player zero-sum game (TPZSG) in which the row player aims to maximize their payoff against an adversarial column player, under an unknown payoff matrix estimated through bandit feedback. We propose and analyze two algorithms: ETC-TPZSG, which directly applies ETC to the TPZSG setting and ETC-TPZSG-AE, which improves upon it by incorporating an action pair elimination (AE) strategy that leverages the $\varepsilon$-Nash Equilibrium property to efficiently select the optimal action pair. Our objective is to demonstrate the applicability of ETC in a TPZSG setting by focusing on learning pure strategy Nash Equilibrium. A key contribution of our work is a derivation of instance-dependent upper bounds on the expected regret for both algorithms, has received limited attention in the literature on zero-sum games. Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC-TPZSG and $O(\frac{\log(T\Delta^{2})}{\Delta})$ for ETC-TPZSG-AE, where $\Delta$ denotes the suboptimality gap. Therefore, our results indicate that ETC-based algorithms perform effectively in adversarial game settings, achieving regret bounds comparable to existing methods while providing insights through instance-dependent analysis.
1 Introduction
This paper focuses on finite two-player zero-sum games (TPZSGs), where two players interact in an adversarial setting with strictly opposing objectives, as first analyzed in v1928theorie ; vonneumanandmorgenstern . These games play a foundational role in game theory and online learning, capturing adversarial interactions in various applications such as economic competition and adversarial machine learning. In such settings, players repeatedly select actions from finite sets, with payoffs determined by a fixed but unknown game matrix.
Studies on learning in games typically use either full information feedback where players observe the opponent’s actions or the entire payoff matrix after each round or partial feedback, also called bandit feedback, where each player observes only their own payoffs without access to the opponent’s strategy or the full matrix blum2007external . In this paper, we focus specifically on a TPZSG setting with bandit feedback and propose two algorithms, ETC-TPZSG and ETC-TPZSG-AE, designed to minimize regret by learning pure strategies for both players through balancing of exploration and exploitation.
Several existing studies have advanced our understanding of how players learn to play optimally in zero-sum games with bandit feedback. The variant of UCB algorithm for adversarial games under bandit feedback have been analyzed using worst-case regret bounds in o2021matrix . In particular, worst-case analyses lead to all games as equally hard by ignoring the specific structure of the game. Instance-dependent analysis adapts the regret guarantees to the specific properties of a game to provide more practical performance guarantees. More recently, instance-dependent regret bounds have been investigated for the Tsallis-INF algorithm in TPZSGs ito2025instance . However an understanding of instance-dependent regret in TPZSGs under bandit feedback remains incomplete.
Contributions.
This paper investigates the effectiveness of the Explore-Then-Commit (ETC) algorithm and extensions in the context of TPZSGs. While ETC has been widely studied in standard multi-armed bandit (MAB) problems, there has been no prior work specifically analyzing its performance to adversarial game settings, which makes this the first work to provide its detailed analysis in such a setting. A key motivation for studying ETC is its algorithmic simplicity allowing for a clear structure.
In particular, we propose two algorithms that adapt the ETC approach to TPZSG setting and present their instance-dependent regret analyses. The first one, ETC-TPZSG, follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure strategy, achieves an upper bound of $\tilde{O}(\Delta+\sqrt{T})$ The second algorithm, ETC-TPZSG-AE, incorporates an adaptive elimination approach, similar to auer2010ucb , by sequentially eliminating actions which, with high probability, are not part of an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), gives us an upper regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$. Intuitively, this design is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action spaces of the players. Consequently, despite the simplicity of ETC framework, our results demonstrate that ETC-based approaches can be highly effective in adversarial settings. In both cases, the regret bounds depend on a suitable $\Delta$ notion, which varies depending on the type of the regret we are analyzing.
Paper structure.
This paper is organized as follows. Section 2 reviews related work. Section 3 introduces preliminaries and formalizes the two-player zero-sum game (TPZSG) setting. In Section 4, we present the ETC in a TPZSG setting, ETC-TPZSG, and analyze its regret. Section 5 introduces an elimination based algorithm, ETC-TPZSG-AE, that leverages the $\varepsilon$-NE property, and provides its regret analysis. Section 6 offers empirical results validating their theoretical performances. Finally, Section 7 discusses our conclusions and future research directions.
2 Related Work
The multi-armed bandit (MAB) problem, has been introduced by thompson1933likelihood and originally described by robbins1952some , has been widely explored and has become a crucial framework for online decision-making under uncertainty, as explored in works such as audibert2009exploration ; auer2002finite ; auer1995gambling ; auer2010ucb ; blum2007learning ; bubeck2012regret . The MAB problem models scenarios in which a player must repeatedly choose from a set of actions to maximize their own rewards while balancing exploration-exploitation tradeoff, has formalized by lai1985asymptotically and further explored in studies such as audibert2007tuning ; maes2012learning .
A central challenge is identifying the action with the highest expected reward while minimizing suboptimal choices, requiring efficient exploration and statistical confidence. Several approaches guarantee low error probabilities within finite trials, aiming to minimize regret audibert2010best ; gabillon2012best . Since the player has no prior knowledge of expected rewards, a dedicated exploration phase is essential, has motivated numerous studies such as bubeck2009pure ; neu2015explore .
The Explore-Then-Commit (ETC) algorithm, which has been introduced by perchet2016batched , is one of the fundamental approaches in the MAB setting. It explores each action a fixed number of times, then commits to the best-performing action for the remaining rounds. Due to its simplicity and effectiveness, variants of the ETC algorithm are widely used in various decision-making scenarios garivier2016explore ; yekkehkhany2019etc .
Zero-sum games, are a fundamental concept in game theory. They describe competitive scenarios where one player’s gain exactly equals the other’s loss. First highlighted by v1928theorie ; vonneumanandmorgenstern , their theory and applications have been extensively explored blum2007learning ; cesa2006prediction ; washburn2014two . The Minimax Theorem in v1928theorie ; vonneumanandmorgenstern guarantees equilibrium existence in TPZSGs, but finding equilibria becomes significantly harder when the payoff matrix is unknown daskalakis2011near ; daskalakis2009complexity ; mazumdar2019finding ; zinkevich2007new .
A more general concept are Nash Equilibria nash1950equilibrium ; nash1951non . While these are identical to the minimax equilibria in TZPSGs, they are not as easy to compute in general games daskalakis2009complexity . The concept of $\varepsilon$-Nash Equilibrium provides an approximate solution where each player’s strategy is within an $\varepsilon$ tolerance level. daskalakis2009note used the $\varepsilon$-NE concept to efficiently approximate equilibrium in two-player games.
In a zero-sum game setting with bandit feedback, players engage in repeated interactions without full knowledge of the payoff matrix, making it essential to estimate the unknown payoffs by previous observations. Various approaches, such as UCB o2021matrix and Tsallis-INF ito2025instance , have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret. Learning the game matrix is crucial for converging to equilibrium strategies and ensuring optimal decision-making in adversarial environments. o2021matrix focuses on worst-case regret analysis, while ito2025instance provides an instance-dependent regret analysis which achieves an upper bound of $O(\frac{\log T}{\Delta})$ in the case of a pure strategy equilibrium. Additionally, maiti2023instance investigates instance-dependent sample complexity bounds for zero-sum games under bandit feedback while cai2023uncoupled examines convergence rates to NE.
3 Preliminaries
In this section, we define the basic concept of a TPZSG such as payoff matrix, the regret definitions we focus on and the gaps $\Delta$ that we use for the analysis.
In a TPZSG, the goals of the players are strictly opposed, which implies that any gain by one player corresponds to an equal loss by the other. This framework provides a fundamental model for adversarial interactions where strategic decision-making under competition is central. The outcome of such a game is fully determined by the strategies chosen by both players and the goal of each player is to maximize their own payoff while minimizing that of their opponent.
Notation.
Throughout this paper, $O(\cdot)$ denotes an upper bound up to constant factors.
3.1 Two-Player Zero-Sum Games
In a two-player zero-sum game, consisting of a row $(x)$ player and a column $(y)$ player , the row player aims to maximize their own payoff while the column player attempts to take an action to minimize this payoff.
$S_{x}$ and $S_{y}$ are the action sets of the row player and column player, respectively. Thus, the action space of the game is $S_{A}=S_{x}\times S_{y}$, so that the action pairs $(i,j)\in S_{A}$ where $i\in S_{x},j\in S_{y}$. We also define $m=|S_{x}|$ as the number of actions for the row player, and $l=|S_{y}|$ as the number of actions for the column player, while finally $N=ml=|S_{A}|$ is the total number of action pairs. The game itself is defined through a game matrix $A$, so that if the players choose the pair $(i,j)$ then the row player obtains a payoff of $A(i,j)$ and the column player $-A(i,j)$.
The row player has a strategy (i.e. a probability distribution) $p$ to select an action $i\in S_{x}$, while the column player selects an action $j\in S_{y}$ according to a strategy $q$.
Definition 3.1.
A pair of mixed strategies $(p^{*},q^{*})$ is a Nash equilibrium (NE) if for all strategies $p$ and $q$, it satisfies
$V^{*}$ $\geq{p}^{\top}Aq^{*},$ $V^{*}$ $\leq{p^{*}}^{\top}Aq$ (1)
where $V^{*}={p^{*}}^{\top}Aq^{*}$ is the value of the game nash1950equilibrium . It means that $p^{*}$ and $q^{*}$ are optimal strategies for row and column players, respectively.
There always exists an optimal mixed strategy for both players that guarantees the best possible outcome by the Minimax Theorem v1928theorie ; vonneumanandmorgenstern . When a mixed strategy assigns a probability of one to a single action and zero to all others, it is referred to as a pure strategy. In our setting, we focus on a TPZSG where both players adopt pure strategies.
Definition 3.2 (Pure Nash Equilibria).
A pair $(i^{*},j^{*})$ is a pure NE if:
$i^{*}=\operatorname*{argmax}_{i\in S_{x}}\min_{j\in S_{y}}A(i,j)\quad\text{ and }\quad j^{*}=\operatorname*{argmin}_{j\in S_{y}}\max_{i\in S_{x}}A(i,j),$
or equivalently
$A(i,j^{*})\leq A(i^{*},j^{*})\leq A(i^{*},j),\qquad\forall i\in S_{x},j\in S_{y}.$ (2)
The condition in (2) ensures that no player has an incentive to deviate to another action. We use $V^{*}=A(i^{*},j^{*})$ to denote the value of the game.
3.2 Games with Bandit Feedback
The game proceeds in rounds. At time $t$, the row player draws an action $i_{t}\in S_{x}$ from a strategy $p_{t}$, and the column player draws $j_{t}\in S_{y}$ from a strategy $q_{t}$. The players then observe a noisy payoff $r_{t}$ with expected value $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, with the maximizing player obtaining $r_{t}$ and the minimizing player $-r_{t}$. We assume that both players observe each other’s action.
Since the algorithms do not know $A$, we use the notation $\hat{A}(i,j)$ to denote the estimated payoff when the action pair $(i,j)$ is played. We express the action pair played in round $t$ by $(i^{t},j^{t})$. After round $t$, we estimate each entry of the estimated game matrix by computing the average payoff for each action pair $(i,j)$ as
$\hat{A}_{t}(i,j)=\frac{1}{n_{ij,t}}\sum_{s=1}^{t}\mathbb{I}\{(i^{s},j^{s})=(i,j)\}r_{s}$ (3)
where $n_{ij,t}$ is the number of times the action pair $(i,j)$ played until round $t$, $r_{s}$’s are independently and identically distributed (iid) $\sigma$-subgaussian payoffs. We assume that the estimator is unbiased, meaning that $\mathbb{E}[\hat{A}]=A$.
3.3 Gap Notions
To quantify how close a chosen action pair is to the NE, we define suboptimality gaps based on deviations from the equilibrium and the best responses of the players. These gaps capture the extent to which each player could improve their outcome by deviating from their current strategy. For any action $i\in S_{x},j\in S_{y}$, we define the following suboptimality gaps:
$\Delta_{ij}^{\text{max}}=\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)-A(i,j)$ (4)
$\Delta_{ij}^{\text{min}}=A(i,j)-\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})$ (5)
$\Delta_{ij}=\Delta_{ij}^{\text{max}}+\Delta_{ij}^{\text{min}}$ (6)
$\Delta_{ij}^{*}=A(i^{*},j^{*})-A(i,j)=V^{*}-A(i,j)$ (7)
We note that $\Delta_{ij}^{\text{max}}\geq 0$ and $\Delta_{ij}^{\text{min}}\geq 0$, thus $\Delta_{ij}\geq 0$ for any action pair $(i,j)$. However, $\Delta_{ij}^{*}$ might be negative.
Using the NE property in (2), for all $i\in S_{x}$ and $j\in S_{y}$ we can write the following:
$A(i^{*},j^{*})\leq A(i^{*},j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)\implies A(i^{*},j^{*})-A(i,j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)-A(i,j)$ (8)
$A(i^{*},j^{*})\geq A(i,j^{*})\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})\implies A(i^{*},j^{*})-A(i,j)\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})-A(i,j)$ (9)
Then, the inequality (8) implies that $\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$ and similarly from (9), we have $\Delta_{ij}^{*}\geq-\Delta_{ij}^{\text{min}}$. Therefore, we have $-\Delta_{ij}^{\text{min}}\leq\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$. On the other hand, since $\Delta_{ij}^{\text{min}}\geq 0$, we can write
$\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}+\Delta_{ij}^{\text{min}}=\Delta_{ij}.$ (10)
3.4 Regret Notions
To characterize the performance of our algorithms, we consider several regret notions. Let begin with the following formulations, which are the external regret of the max and min player respectively:
$R_{T}^{\text{max}}=\max_{i\in S_{x}}\mathbb{E}\Big{[}\sum_{t=1}^{T}A(i,j^{t})-A(i^{t},j^{t})\Big{]}$
$R_{T}^{\text{min}}=\max_{j\in S_{y}}\mathbb{E}\Big{[}\sum_{t=1}^{T}A(i^{t},j^{t})-A(i^{t},j)\Big{]}$
where $i^{t}$ and $j^{t}$ are the actions selected by the row and column player in round $t$, respectively. $R_{T}^{\text{max}}$ represents the expected loss due to not selecting the best possible action of row player against column player’s action $j$ while $R_{T}^{\text{min}}$ captures the expected regret from not choosing the column player’s best action for a given action $i$.
We can now rewrite the regret definitions in terms of the suboptimality gaps:
$R_{T}^{\text{max}}$ $=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}],$ $R_{T}^{\text{min}}$ $=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}],$ (11)
where $\mathbb{E}[n_{ij,T}]$ denotes the expected number of times the action pair $(i,j)$ is played over $T$ rounds.
In this paper, we analyze two regret notions. The external regret and the Nash regret.
Definition 3.3 (External regret).
The (combined) external regret is given by the following:
$R_{T}=R_{T}^{\text{max}}+R_{T}^{\text{min}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}\mathbb{E}[n_{ij,T}]$ (12)
External regret quantifies how much worse the players have performed compared to their best actions in hindsight. An alternative regret notion is the Nash regret, which measures the deviation of value from the Nash equilibrium:
Definition 3.4 (Nash regret).
The Nash regret is expressed as
$R_{T}^{*}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}].$ (13)
The Nash regret, as defined here, can be negative, since the minimizing player might play suboptimally. However, if we are able to bound the Nash regret for the maximizing player, then by symmetry, we can also bound it for the minimizing player, thus effectively bounding its absolute value.
4 ETC-TPZSG Algorithm and Regret Analysis
In this section, we extend ETC algorithm to a two-person zero-sum game setting as ETC-TPZSG. Each player selects action from their corresponding finite action sets, which are known. The game is repeated over given time horizon $T$, and the true payoff matrix $A$ is unknown to both players. Instead, they receive observations of the payoffs based on their chosen actions, which are iid subgaussian random variables.
Let $k$ represent the number of times to explore each action pair $(i,j)\in S_{A}$. During the exploration phase, each player samples actions to estimate the expected payoffs, which enables us to construct an empirical payoff matrix using the observed payoffs. Then, in commit phase, they play according to the pure NE strategy derived from the estimated payoff matrix. After presenting the ETC-TPZSG algorithm, finally, we analyze its expected regret, measuring the performance due to limited information and suboptimal action pairs during the exploration phase.
As a result, the unknown payoff matrix is approximated by the ETC approach, as in Lattimore_Szepesvari_2020 for a bandit setting. The ETC algorithm consists of two phases: an exploration phase, where the player randomly samples actions to gather information with given exploration time $k$, and a commit phase, where the player selects the empirically best action based on the gathered data. Algorithm 1 presents the implementation of ETC-TPZSG, where each player only observes the payoffs of the action pair they choose during the exploration period, then in the committing phase the algorithm converges to an optimal action pair which refers to the pure NE.
Applying ETC in a zero-sum game setting presents some challenges compared to the standard MAB scenario. In bandit problems, after the exploration phase, a player simply commits to the arm with the highest estimated reward, which is typically effective if exploration is sufficient. However, in a zero-sum game with the pure NE, the objective is not to identify an action with high reward but to find a possibly true NE. Inadequate exploration can lead to poor estimates of payoffs, making it difficult to correctly identify such an equilibrium. In fact, due to inaccurate estimates, a pure NE might not even appear to exist, even when one does in the true game. As a result, the design of ETC approach in zero-sum game setting requires careful balancing of exploration phase.
1:Input:
2:$S_{A}$: set of action pairs in the game matrix $A$triangleright $S_{A}$ includes action pairs $(i,j)$
3:$m$: number of actions for row player
4:$l$: number of actions for column player
5:$k$: exploration time
6:$T$: time horizon $\triangleright$ $1\leq mlk\leq T$
7:$\sigma^{2}$: subgaussian variance factor
8:Initialize: $\hat{A}=[0]^{m\times l}$triangleright $\hat{A}$ is estimated game matrix
9:In round $t=0,1,2,...,mlk$; $\triangleright$ Exploration Phase
10:        Explore each action pair $(i,j)$ in $S_{A}$ $k$ times
11:        Update $\hat{A}$ using 3
12:In round $t=mlk+1,mlk+2,...,T$; $\triangleright$ Committing Phase
13:        Play an equilibrium $(i^{*},j^{*})$ satisfying the following property:
$\hat{A}(i,j^{*})\leq\hat{A}(i^{*},j^{*})\leq\hat{A}(i^{*},j)$
Algorithm 1 ETC-TPZSG
Theorem 4.1.
The Nash regret of Algorithm 1, when interacting with $\sigma$-subgaussian payoffs, is upper bounded as follows:
$R_{T}^{*}\leq k\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}+(T-Nk)\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}\exp{\Big{(}-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big{)}}$ (14)
where $k$ is the exploration time per action pair and $N$ is the total number of action pairs.
We provide the proof in Appendix B. The regret comprises two main components, the exploration phase and the loss due to misidentifying the NE. Thus, it is crucial to choose an appropriate exploration time $k$ that balances sufficient exploration and efficient exploitation. Choosing $k$ too large leads to unnecessary exploration while setting it too small increases the risk of making suboptimal decisions. Hence, careful tuning of $k$ is essential to minimize overall regret.
Let assume there are two action pairs $(i_{1},j)$ and $(i_{2},j)$ in the game such that row player has two actions and column player has one action. In addition, let suppose the NE is at $(i_{1},j)$. Then, we have $\Delta=A(i_{1},j)-A(i_{2},j)$ and we can write the regret simply as
$R_{T}^{*}\leq k\Delta+(T-2k)\Delta\exp{\Big{(}-\frac{k\Delta^{2}}{16\sigma^{2}}\Big{)}}\leq k\Delta+T\Delta\exp{\Big{(}-\frac{k\Delta^{2}}{16\sigma^{2}}\Big{)}}.$ (15)
Taking the first derivative with respect to $k$ and solve it, we get the following for the exploration time:
$k=\max\Big{\{}1,\Big{\lceil}\frac{16\sigma^{2}}{\Delta^{2}}\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big{\rceil}\Big{\}}$ (16)
If $\Delta$ is known, we can easily find the necessary exploration time. Therefore, when we put it into the regret bound, it becomes
$R_{T}^{*}\leq\min\Big{\{}T\Delta,\Delta+\frac{16\sigma^{2}}{\Delta}\Big{(}1+\max\Big{\{}0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big{\}}\Big{)}\Big{\}}$ (17)
where the first term $T\Delta$ is worst-case regret and for the other, we combine the regret we obtain in (15) with the value of exploration time in (16). Then, focusing on $\max\Big{\{}0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big{\}}$, we obtain that $\Delta=\frac{4\sigma}{\sqrt{T}}$ is a critical point. If we put it in the regret bound (17), we have the following:
$R_{T}^{*}\leq\Delta+c\sqrt{T}$
where $c>0$ is a some constant. If $\Delta\leq 1$, it becomes
$R_{T}^{*}\leq 1+c\sqrt{T}.$ (18)
Bounds like (18) are referred to as instance-independent since it only depends on the time horizon $T$ not on the game instance. In the zero-sum game setting, our analysis shows that we get a regret bound comparable to the standard bandit setting in Lattimore_Szepesvari_2020 , indicating that the ETC-based algorithm performs effectively in adversarial environments and aligns with known results in the literature.
5 ETC-TPZSG-AE Algorithm and Regret Analysis
If a pure NE exists at $(i^{*},j^{*})$, then it must satisfy the condition given in (2). For the algorithm in the section, we make use of the concept of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), which approximately satisfies the standard NE condition. Specifically, an $\varepsilon$-NE satisfies the NE conditions within a tolerance level $\varepsilon$, making it a suitable criterion in learning settings where exact equilibrium identification is hard.
Based on this criterion, we introduce an elimination strategy in Algorithm 2, inspired by auer2010ucb , to reduce unnecessary exploration of clearly suboptimal action pairs. Specifically, action pairs that do not satisfy the $\varepsilon$-NE property are eliminated from further play, which enables a more efficient learning process by concentrating exploration on approximately optimal action pairs.
Definition 5.1.
The action pair $(i^{*},j^{*})$ is an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE) if the following condition holds:
$A(i,j^{*})-\varepsilon\leq A(i^{*},j^{*})\leq A(i^{*},j)+\varepsilon,\quad\forall i\in S_{x},j\in S_{y}.$ (19)
For the exploration time $k$ of the algorithm, let just go back to the regret bound in (14) and assume that there exists some $\Delta^{*}$ and $\Delta$ such that $\Delta^{*}\geq\Delta_{ij}^{*}$ and $\Delta\leq\Delta_{ij}$ for all $(i,j)\in S_{A}$. Then, when we put them into the regret bound and take its derivative with respect to $k$, we obtain a similar result as in (16); thus, as the exploration time, clearly we can use the following:
$k=\Big{\lceil}\frac{16\sigma^{2}}{\Delta^{2}}\ln{\Big{(}\frac{\Delta^{2}T}{16\sigma^{2}}\Big{)}}\Big{\rceil}$ (20)
Since the true suboptimality gaps $\Delta$ are unknown, the algorithm adopts a decreasing $\Delta$ approach across rounds. It is motivated by the intuition that the remaining action pairs become closer to the true NE as suboptimal action pairs are progressively eliminated. Consequently, the suboptimality gaps among the remaining action pairs are expected to decrease over time. In other words, as the rounds progress, we expect that the players get closer to true NE, which means they tend to repeat better action pairs more often. We can afford to use smaller $\Delta$ values to explore action pairs near the NE more thoroughly, since smaller $\Delta$ leads to a longer exploration time $k$.
The algorithm schedules $\Delta_{ij}$ across rounds using $\hat{\Delta}_{t}$, which guides both exploration time and the tolerance level for the $\varepsilon$-NE condition. When $\hat{\Delta}_{t}$ is large, action pairs are played fewer times and more action pairs are kept during elimination. As $\hat{\Delta}_{t}$ decreases, the algorithm focuses exploration on pairs closer to equilibrium, having already eliminated clearly suboptimal ones.
In order to perform updates over rounds, an initial estimate is required. When the payoffs are bounded within a known interval, such as $[0,1]$, it is common practice to initialize the estimated suboptimality gap $\hat{\Delta}_{t}$ with value 1, representing the maximum possible difference between arm rewards as in auer2010ucb . However, in our setting, the payoffs are assumed to be $\sigma$-subgaussian, which implies that there is no strict upper bound on the suboptimality gaps, we only have $\Delta_{ij}\geq 0$. Since no empirical estimates are available at the initialization step, it is not feasible to use a bound for the expected value of the maximum of $\sigma$-subgaussian random variables to guide the choice of $\hat{\Delta}_{t}$. To address this, we initialize $\hat{\Delta}_{t}$ with $4\sigma$, which provides a practical starting point for the analysis.
Moreover, we use $\hat{\Delta}_{t}=2^{-t+2}\sigma$ in each round $t$ which takes values from 00 to $\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$. Regarding $\Delta\approx\sqrt{(16\sigma^{2}/k)\ln{(\Delta^{2}T/16\sigma^{2})}}$ from (20) so $\ln{(\Delta^{2}T/16\sigma^{2})}\geq 0$, which gives us $\Delta\geq\sqrt{16\sigma^{2}/T}$. Then, using $2^{-t+2}\sigma\geq\sqrt{16\sigma^{2}/T}$ and after some calculations, we obtain $t\leq\frac{1}{2}\log_{2}T$. The division by $e$ is a technical adjustment, which makes the bound safer in the analysis; otherwise the number of play would be zero in the last round.
If we assume that a unique pure NE exists in the game as before, a well estimated game matrix $\hat{A}$ should also exhibit the properties necessary to support the equilibrium. Specifically, $\hat{A}$ must satisfy the conditions required for the $\varepsilon$-NE in (19) to hold. Based on this principle, the Algorithm 2 systematically eliminates action pairs that fail to satisfy this condition, as such pairs cannot be near to the equilibrium. By reducing the set of action pairs to include only the pairs that meet this criterion, it ensures that the remaining action pairs are the only ones that could potentially be near the NE; thus, it enables to focus on the most relevant action pairs. Consequently, we expect that the algorithm converges to an accurate approximation of the NE, even in the presence of estimation errors in $\hat{A}$.
1:Input:
2:$S_{A}$: set of action pairs in the game matrix $A$triangleright $S_{A}$ includes action pairs $(i,j)$
3:$m$: number of actions for row player
4:$l$: number of actions for column player
5:$T$: time horizon $\triangleright$ $1\leq ml\ll T$
6:$\sigma^{2}$: subgaussian variance factor
7:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$triangleright $\hat{A}$ is estimated game matrix
8:In round $t=0,1,2,...,\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$;
9:        Action Pair Selection:
10:            If $\mathbf{|S_{t}|>1}:$triangleright Exploration Phase
11:              Explore each action pair $(i,j)$ in $S_{t}$ $k_{t}=\Big{\lceil}\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}\Big{\rceil}$ times
12:              Update $\hat{A}$ using (3).
13:            Else: $\triangleright$ Committing Phase
14:              Play with a unique action pair in $S_{t}$ until step $T$
15:        Action Pair Elimination:
16:            Remove the action pairs $(i,j)$ NOT satisfying the following for $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}}$ :
$\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
17:        Update:
18:            Reset $S_{t+1}$ by "Action Pair Elimination" step
19:            $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
Algorithm 2 ETC-TPZSG-AE
Theorem 5.1.
The Nash regret of Algorithm 2, interacting with $\sigma$-subgaussian payoffs, is upper bounded by
$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big{(}1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big{(}\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big{)}}\Big{)}+\sum_{(i,j)\in S_{A_{2}}}\Big{(}\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}+\Delta_{ij}^{*}T\Big{)}$ (21)
where $\lambda\geq\sqrt{\frac{16\sigma^{2}e}{T}}$ and, $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$ are two subsets of the action pairs.
Moreover, if we consider another regret approach in (12), we obtain the following regret bound.
Theorem 5.2.
The external regret incurred by Algorithm 2 when interacting with $\sigma$-subgaussian payoffs is bounded as following:
$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Big{(}\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\Big{(}\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big{)}}\Big{)}+\sum_{(i,j)\in S_{A_{2}}}\Big{(}\frac{512\sigma^{2}}{\lambda}+\lambda T\Big{)}$ (22)
We simplify the regret bound to $O(\frac{\log(T\Delta^{2})}{\Delta})$, as the logarithmic term dominates since by setting $\lambda=\sqrt{\frac{16\sigma^{2}e}{T}}$, we ensure that the term $\lambda T$ is bounded by $\sqrt{16\sigma^{2}eT}$ which is at most $\frac{e}{\Delta_{ij}}$ when $\Delta_{ij}\leq\lambda$. Our results are consistent with the regret bound derived by auer2010ucb . The proofs of the theorems are provided in Appendix C.
In this section, we analyze two distinct regret formulations for Algorithm 2, Nash regret, defined in (3.4), and external regret, defined in (3.3). The first, denoted by $R_{T}^{*}$, evaluates performance against the NE, capturing the cumulative loss relative to the equilibrium strategy. In contrast, the regret analysis by $R_{T}$ allows us to consider the losses of both players with respect to their individual best responses at each round, providing a broader and more flexible notion of regret. By the inequality $\Delta{ij}^{*}\leq\Delta{ij}$ for all action pairs, it follows that $R_{T}^{*}\leq R_{T}$. This outcome aligns with the interpretation that the expected regret against the NE establishes a more specific comparison and thus, it naturally provides a tighter upper bound on the expected regret.
6 Experiments
In this section, we present a set of simulations to support our theoretical findings and demonstrate the performance of the proposed algorithms. The experiments are conducted with fixed time horizons of $T=10^{3}$ and $T=10^{4}$. Additional details are provided in Appendix D.
(a) The expected regret of ETC-TPZSG with $k$ in (16) and the upper bound in (17)
(b) The cumulative regrets of ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF from ito2025instance
(c) The theoretical regret bounds for ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF from ito2025instance
Figure 1: Regret performance comparison of the algorithms
Figure 1(a) shows the expected regret bound, averaged over $10^{5}$ simulation runs, of ETC-TPZSG using $k$ in (16) and the upper bound in (17). We consider a setting with $N=2$, where the row player has two actions, the column player has one, and the first row corresponds to the NE. The suboptimality gap $\Delta$ varies from 0 to 1 and the subgaussian parameter is set to $\sigma=0.5$. The results align with the bound in Theorem (4.1).
In Figure 1(b), we compare ETC-TPZSG, ETC-TPZSG-AE, and Tsallis-INF ito2025instance , the only existing method with instance-dependent bounds for TPZSGs, based on cumulative regret, computed as the absolute difference between the game value and the payoff from the action played to avoid negative values, averaged over $10^{3}$ runs. The results show that ETC-TPZSG-AE achieves lower regret than ETC-TPZSG, highlighting the effectiveness of its action pair elimination strategy. We use a $2\times 2$ game matrix by Gaussian payoffs and $k$ exploration rounds per action pair for ETC-TPZSG is randomly selected from a predefined list. On the other hand, Figure 1(c) compares their theoretical regret bounds ($O(\cdot)$ notation) for $\Delta=0.5$, which aligns with the results in Figure 1(b).
7 Discussion
We investigate a TPZSG with bandit feedback, where the payoff matrix is unknown and must be learned through player interactions. This setting is challenging, as players must estimate payoffs while making strategic decisions in an adversarial environment. We adopt the ETC algorithm due to its simplicity, widespread use and lack of prior analysis in this context. We also integrate an elimination based method, allowing the systematic removal of suboptimal action pairs, thereby improving convergence to the equilibrium and reducing unnecessary exploration on suboptimal ones. A key contribution of our work is the derivation of instance-dependent upper bounds on the expected regret for both algorithms, which has received limited attention in the literature on zero-sum games.
While our study focuses on pure strategy learning in a zero-sum game with bandit feedback and provides a theoretical expected regret bounds, several directions remain open for future study. One interesting extension is to consider games where the equilibrium is mixed. While the algorithms provided can be used to identify the support of the equilibrium, another approach should be used to efficiently converge to a mixed equilibrium.
Finally, an important and relatively unexplored direction is fairness in zero-sum games. For example, introducing mechanisms that ensure similar action pairs are explored equally can promote fairness in strategy estimation. This type of fair play mechanism could be essential in TPZSG settings, thus it can promote balanced exploration and improve overall strategy estimation. On the other hand, by integrating fairness based algorithms or constraints, it may be possible to generate game environments that ensure balanced opportunities for all players.
Appendix A Subgaussian Properties
This section collects some supplementary results from [20] that are used repeatedly throughout our proofs. These results concern standard properties of subgaussian random variables such as tail inequalities. They are included here for completeness and easy reference since they play a critical role in our analyses. The statements are not original, we include only the parts that are directly useful for our purposes.
Theorem A.1 ([20, Theorem 5.3]).
If $X$ is $\sigma$-subgaussian random variable, for any $\epsilon\geq 0$,
$\mathbb{P}(X\geq\epsilon)\leq\exp{\Big{(}-\frac{\epsilon^{2}}{2\sigma^{2}}\Big{)}}.$ (23)
Lemma A.1 ([20, Lemma 5.4]).
If $X$ is $\sigma$-subgaussian and $X_{1}$ and $X_{2}$ are independent with $\sigma_{1}$ and $\sigma_{2}$ subgaussian parameter, respectively, then we can write the followings:
  1. (a)
    $cX$ is $|c|\sigma$-subgaussian for all $c\in\mathbb{R}.$
  2. (b)
    $X_{1}+X_{2}$ is $\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}}$-subgaussian.
Appendix B Proof of Theorem 4.1
To analyze the Nash regret of the ETC-TPZSG algorithm, we begin by decomposing the total expected regret into two phases: the exploration phase and the committing phase. During exploration, each action pair is played a fixed number of times $k$ to estimate their mean rewards, potentially incurring regret if suboptimal action pairs are played. Once the algorithm commits to the empirically best action pair, which is the pure NE, the regret accumulates only if this action pair is not the true optimal one. Our goal is to bound the expected regret by quantifying the probability of committing to a suboptimal action pair, which we can call it as a misidentified NE, based on the estimates of their payoffs obtained during exploration phase.
Let $(i^{\prime},j^{\prime})$ denote a misidentified NE, so we aim to identify an action pair $(i^{\prime},j^{\prime})$, which appears better than $(i^{*},j^{*})$ based on the estimated matrix. To analyze the probability of playing with a misidentified NE, we consider whether the following two events, $E_{1}$ and $E_{2}$, occur:
$E_{1}:\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime}),\quad E_{2}:\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j),\quad\forall i\in S_{x},\forall j\in S_{y}$
and we want to find a bound for $\mathbb{P}(E_{1}\cap E_{2})$ because these two conditions must be satisfied to consider $(i^{\prime}.j^{\prime})$ as a NE.
Since the events $E_{1}$ and $E_{2}$ are not independent, we consider two different approaches to find an upper bound for $\mathbb{P}(E_{1}\cap E_{2})$. We then select the approach that gives the tighter bound, as it provides a more accurate estimate of this probability. The first approach uses the fact that $\mathbb{P}(E_{1}\cap E_{2})\leq\mathbb{P}(E_{1})$ and $\mathbb{P}(E_{1}\cap E_{2})\leq\mathbb{P}(E_{2})$, which together imply $\mathbb{P}(E_{1}\cap E_{2})\leq\sqrt{\mathbb{P}(E_{1})\mathbb{P}(E_{2})}$. Alternatively, we can use the inequality $\mathbb{P}(E_{1}\cap E_{2})\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}$, which may provide a tighter bound depending on the values of $\mathbb{P}(E_{1})$ and $\mathbb{P}(E_{2})$. The choice between these approaches depends on which enables us the smaller upper bound.
Let us start by considering the first approach. We can write it as
$\mathbb{P}(E_{1}\cap E_{2})$ $=\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j),\forall i,j)$ (24)
$\leq\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))$ (25)
$=\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime}))\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))}$ (26)
$=\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)}$ (27)
$\leq\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq\Delta_{i^{\prime}j^{\prime}}^{\text{max}})\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})}$ (28)
$\leq\sqrt{\exp{\Big{(}-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\Big{)}}\exp{\Big{(}-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\Big{)}}}$ (29)
$=\sqrt{\exp{\Big{(}-\frac{k[(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}]}{4\sigma^{2}}\Big{)}}}$ (30)
$\leq\sqrt{\exp{\Big{(}-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{8\sigma^{2}}\Big{)}}}$ (31)
$=\sqrt{\exp{\Big{(}-\frac{k(\Delta_{i^{\prime}j^{\prime}})^{2}}{8\sigma^{2}}\Big{)}}}$ (32)
$=\exp{\Big{(}-\frac{k(\Delta_{i^{\prime}j^{\prime}})^{2}}{16\sigma^{2}}\Big{)}}$ (33)
where $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}=\max_{i}A(i,j^{\prime})-A(i^{\prime},j^{\prime})$, $\Delta_{i^{\prime}j^{\prime}}^{\text{min}}=A(i^{\prime},j^{\prime})-\min_{j}A(i^{\prime},j)$ and $\Delta_{i^{\prime}j^{\prime}}=\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}}$.
To get (26), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, that is, for any collection of events $e_{1},e_{2},\ldots,e_{n}$, we have $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{n})\leq\mathbb{P}(e_{i})$ for all $i\in{1,\ldots,n}$. The inequality (28) holds since $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}\geq 0,\Delta_{i^{\prime}j^{\prime}}^{\text{min}}\geq 0$, which are true from their definitions. For the step (29), we apply Theorem A.1 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable.
At this point, we need to show that the differences $\hat{A}(i,j)-\hat{A}(x,y)$ are $\sqrt{2\sigma^{2}/k}$-subgaussian where $(i,j)$ and $(x,y)$ are any action pairs in $S_{A}$. To establish this, we utilize Lemma A.1, which provides key properties of subgaussian random variables. Specifically, both $\hat{A}(i,j)$ and $\hat{A}(x,y)$ are $\sqrt{\sigma^{2}/k}$-subgaussian as the average payoffs are computed by the equation (3) and during exploration phase, each action pair is played $k$ times. By the lemma, the difference of two independent subgaussian random variables with parameters $\sqrt{\sigma^{2}/k}$ is subgaussian with parameter $\sqrt{2\sigma^{2}/k}$, which follows that $\hat{A}(i,j)-\hat{A}(x,y)$ is $\sqrt{2\sigma^{2}/k}$-subgaussian for any action pairs $(i,j)$ and $(x,y)$.
Then, to get the last inequality (31), we observe that
$\frac{\Delta_{i^{\prime}j^{\prime}}^{2}}{2}=\frac{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{2}\leq(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}$
where the inequality follows from the fact that $(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}-\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\geq 0$.
As we mention before, there is an alternative approach to bound $\mathbb{P}(E_{1}\cap E_{2})$, which means to find a bound for the probability of a misidentified NE. It involves an inequality based on the minimum of the individual probabilities. Similarly, let follow this:
$\mathbb{P}(E_{1}\cap E_{2})$ $=\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j),\forall i,j)$ (34)
$\leq\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))$ (35)
$=\min\{\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})),\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))\}$ (36)
$=\min\{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)\}$ (37)
$\leq\min\{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq\Delta_{i^{\prime}j^{\prime}}^{\text{max}}),\mathbb{P}(\hat{A}(i^{\prime}j^{\prime})-\hat{A}(i^{\prime}j)\leq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})\}$ (38)
$\leq\min\Big{\{}\exp{\Big{(}-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\Big{)}},\exp{\Big{(}-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\Big{)}}\Big{\}}$ (39)
$=\exp{\Big{(}-\frac{k\max\{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2},(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\}}{4\sigma^{2}}\Big{)}}$ (40)
$\leq\exp{\Big{(}-\frac{k[(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}]}{8\sigma^{2}}\Big{)}}$ (41)
$\leq\exp{\Big{(}-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{16\sigma^{2}}\Big{)}}$ (42)
$\leq\exp{\Big{(}-\frac{k\Delta_{i^{\prime}j^{\prime}}^{2}}{16\sigma^{2}}\Big{)}}$ (43)
where to get the inequality (41), we use the following:
$\max\{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2},(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\}\geq\frac{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{2}$ (44)
If we assume that their maximum is $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}$, then it will imply $(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}\geq(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}$. Similarly, if $\max\{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2},(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\}=(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}$, it will give us $(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\geq(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}$. They are true by assumptions, then, (44) holds.
Thus, both approaches provide the same upper bound for the probability of misidentified NE. In the regret analysis, we will focus on the loss due to playing with suboptimal (or not true NE) action pairs. Thus, after $T$ times playing, $1\leq Nk\leq T$ where $N=ml$ as the number of action pairs, we can write the Nash regret as
$R_{T}^{*}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]$
where $\mathbb{E}[n_{ij,T}]$ is the expected number of playing the action pair $(i,j)$. We already know that each action pair $(i,j)\in S_{A}$ is played at least $k$ times for exploration. After the exploration step, we will consider playing the misidentified NE $(i^{\prime},j^{\prime})$ instead of real one $(i^{*},j^{*})$ because we expect the players to select NE action pair as optimal after exploration phase. Since we already have the probability of playing a misidentified NE from (33) or (43), we can write
$\mathbb{E}[n_{ij,T}]\leq k+(T-Nk)\exp{\Big{(}-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big{)}}.$
Hence, the proof is concluded.
$\Box$
Appendix C Proof of Theorem 5.1
According to the algorithm, $\sqrt{\frac{16\sigma^{2}e}{T}}$ is a critical value for $\hat{\Delta}_{t}$ since it takes this value in the last round. In this way, let define a threshold parameter as $\lambda\geq\sqrt{\frac{16\sigma^{2}e}{T}}$, which gives us two subsets of actions pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$. Since $\Delta_{ij}^{*}\leq\Delta_{ij}$ ensures the bounds are preserved, it makes sense to use these action pair sets in the analysis. Thus, we can write the Nash regret as
$R_{T}^{*}=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]+\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]$
where $\Delta_{ij}^{*}=A(i^{*},j^{*})-A(i,j).$ and $n_{ij,T}$ is the total number of times to play action pair $(i,j)$ by time $T$.
To analyze the Nash regret of the ETC-TPZSG-AE algorithm, it is essential to consider various scenarios that may lead to suboptimal outcomes. Each of these scenarios contributes to the total Nash regret, as they either involve the incorrect elimination of the optimal action pair or the unnecessary selection of suboptimal pairs. The former results in discarding the optimal strategy because of insufficient evidence, while the latter leads to spend more time with action pairs that are unlikely to be part of any near optimal equilibrium.
The algorithm employs a scheduling mechanism for $\Delta_{ij}$, where $\hat{\Delta}_{t}$ is halved in each round. This approach is necessary because the true values of $\Delta{ij}$ are unknown to the players. We use $\hat{\Delta}_{t}$ to determine both the exploration time and the tolerance level in the $\varepsilon$-NE property. Specifically, when $\hat{\Delta}_{t}$ is large, the algorithm plays each action pair fewer times and keeps more pairs during the elimination step, since the threshold for elimination is relatively loose. As $\hat{\Delta}_{t}$ decreases, the algorithm allocates more exploration to action pairs that are closer to being part of a NE, because those that are clearly suboptimal have already been eliminated. This adaptive process enables more precise identification of near-optimal strategies by $\varepsilon$-NE condition while minimizing regret.
For each suboptimal action pair $(i,j)$, let $t_{ij}=\min\{t:\hat{\Delta}_{t}<\Delta_{ij}/2\}$ refer to the earliest round such that $\hat{\Delta}_{t}<\Delta_{ij}/2$. Using the fact that $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ and the definition of $t_{ij}$, we can write the following:
$\frac{1}{\hat{\Delta}_{t_{ij}}}\leq\frac{4}{\Delta_{ij}}<\frac{1}{\hat{\Delta}_{t_{ij}+1}}$ (45)
For simplicity of notation, we write $\hat{A}$ instead of $\hat{A}_{t}$ for the estimated payoff matrix in round $t$ throughout this section. Since we already include additional terms such as $\varepsilon_{t}$ that refers to the tolerance level in round $t$, it is clear enough which round is being referenced.
We now analyze the regret contributions by considering the following cases:
Case C.1.
A suboptimal action pair $(i,j)$ is not eliminated in round $t_{ij}$ or earlier while the optimal action pair is in the set $S_{t_{ij}}$.
Since neither some suboptimal nor the optimal action pair is not eliminated, the algorithm fails to discard suboptimal choices. It might converge to an incorrect solution without eliminating suboptimal actions, but since it still maintains the optimal action pair within the action pair selection set. The regret contribution of this case comes from the fact that the algorithm spends time with a worse choice when a better one is already available.
Let explain briefly in which situations the algorithm eliminates or keep an action pair. For action pair $(i,j)$, if $\varepsilon$-NE property does not hold in round $t=t_{ij}$ which implies that at least one of the inequalities fails, then it will be eliminated in round $t_{ij}$.
Furthermore, we have $\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\Big{(}\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\Big{)}}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{t_{ij}+1}<\frac{\Delta_{ij}}{4}$. In other words, if there exist $i^{\prime}\in S_{x}$ or $j^{\prime}\in S_{y}$, such that one of the following is true:
$I_{1}:\hat{A}(i^{\prime},j)-\hat{A}(i,j)>\varepsilon_{t_{ij}},\qquad I_{2}:\hat{A}(i,j)-\hat{A}(i,j^{\prime})>\varepsilon_{t_{ij}},$
then $(i,j)$ is eliminated. The inequality $I_{1}$ indicates that there exists an alternative action $i^{\prime}$ for the row player that offers a higher payoff than action $i$ against the column action $j$. Similarly, $I_{2}$ implies that there exists a better action $j^{\prime}$ for the column player than action $j$ when playing against the row action $i$. If either $I_{1}$ or $I_{2}$ holds, the action pair $(i,j)$ cannot be a part of the NE, as at least one player has an incentive to deviate. Hence, $(i,j)$ is not a NE with high probability.
On the other hand, to keep an action pair $(i,j)$ the following events must both hold:
$E_{1}:\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j),\forall i^{\prime}\in S_{x},\quad E_{2}:\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall j^{\prime}\in S_{y}$ (46)
To calculate the probability of keeping an action pair $(i,j)$, denoted by $\mathbb{P}(E_{1}\cap E_{2})$, we begin by analyzing the event $E_{1}$. Specifically, for an action pair $(i,j)$ which is not a NE, i.e. there exists at least one $i^{\prime}$ such that $A(i^{\prime},j)>A(i,j)$, the probability of $E_{1}$ can be expressed as follows:
$\mathbb{P}(E_{1})$ $=\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t},\forall i^{\prime}\in S_{x})$ (47)
$\leq\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t})$ (48)
$\leq\exp{\Bigg{(}-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big{(}\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big{)}}}{\frac{4\sigma^{2}}{k_{t}}}\Bigg{)}}$ (49)
$=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}$ (50)
where $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{\Big{(}\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big{)}}}.$ To derive inequality (48), we use the fact that the probability of the intersection of multiple events is always less than or equal to the probability of any of individual events. Specifically, since we have $m$ actions in $S_{x}$, for any collection of events $e_{1},e_{2},\cdots,e_{m}$, it holds that $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{m})\leq\mathbb{P}(e_{i})$ for all $i=1,2,\cdots,m$. In the last inequality (49), we apply Theorem A.1 with $\sqrt{2\sigma^{2}/k_{t}}$-subgaussian random variable. The subgaussian parameter is derived by using Lemma A.1 and we note that each action pair is played $k_{t}$ times.
Similarly, in order to calculate the probability of the event $E_{2}$, we can write
$\mathbb{P}(E_{2})=\mathbb{P}(\hat{A}(i,j)-\hat{A}(i,j^{\prime})\leq\varepsilon_{t},\forall j^{\prime}\in S_{y})\leq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (51)
Therefore, we have the probability of keeping an action pair $(i,j)$ in any round $t$ as
$\mathbb{P}(E_{1}\cap E_{2})\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (52)
It means that the probability that a suboptimal action pair $(i,j)$ is not eliminated in round $t_{ij}$ or before is bounded by $\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}T}$. Then, the regret contribution is simply bounded by a worst case which is $T\Delta_{ij}^{*}$ for any suboptimal action pair $(i,j)\in S_{A_{1}}$.
Hence, using the probability of keeping a suboptimal action pair and summing up over all action pairs, we can write their regret contribution as
$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}T\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}T}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\frac{256\sigma^{2}}{\Delta_{ij}^{2}}$ (53)
where we apply inequality (45) to bound it.
Here, we note that we do not need to consider the action pairs in $S_{A_{2}}$ because, by the design of the algorithm, the gap $\hat{\Delta}_{t}$ is guaranteed to be at least around $\lambda$ in the corresponding rounds. If the optimal action pair remains in the game, then all suboptimal action pairs should have already been eliminated by the last round or earlier. This means that suboptimal choices are progressively discarded over time as the algorithm learns. As a result, after all the rounds have been played, only one action pair should remain since we consider the case the optimal action pair $(i^{*},j^{*})$ remain in the game in addition to elimination of suboptimal ones. This final pair is expected to correspond to the pure NE, representing the best choice for both players with no incentive to change their actions.
Case C.2.
A suboptimal action pair $(i,j)$ is eliminated in round $t_{ij}$ or earlier with the optimal action pair in $S_{t_{ij}}$.
It enables us to bound the number of times it is played. Once a suboptimal action pair is eliminated and the optimal action pair $(i^{*},j^{*})$ remains in the game, it no longer contributes to the regret in next rounds. This is because regret arises only when a suboptimal action is chosen instead of the optimal one.
On the other hand, we note that there is no need to consider the action pairs $(i,j)\in S_{A_{2}}$ because we handle the elimination of suboptimal pairs before the final round. Thus, using (45), each action pair $(i,j)$ is played at most $k_{t_{ij}}$ times such that $k_{t_{ij}}=\Big{\lceil}\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big{\rceil}\leq\Big{\lceil}\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big{(}\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big{)}}\Big{\rceil}$.
Then, the regret contribution is expressed by
$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big{\lceil}\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big{(}\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big{)}}\Big{\rceil}$ $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big{(}1+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big{(}\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big{)}}\Big{)}$ (54)
$=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big{(}\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big{)}}$ (55)
Case C.3.
The optimal action pair $(i^{*},j^{*})$ is eliminated by some suboptimal one $(i,j)$ in round $t_{*}$ such that $t_{ij}\geq t_{*}$.
It leads to a misidentification of the NE, implying that an action pair $(i,j)$ should be better than the optimal one based on the current estimates. By $\varepsilon$-NE property, clearly we have
$\hat{A}(i^{*},j)-\varepsilon_{t_{*}}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{*})+\varepsilon_{t_{*}}.$
However, it is not enough to keep a suboptimal action pair $(i,j)$. On the other hand, we note that if the optimal action pair $(i^{*},j^{*})$ is eliminated in round $t_{*}$, it must fail to satisfy $\varepsilon$-NE property as the following:
$\hat{A}(i,j^{*})-\hat{A}(i^{*},j^{*})>\varepsilon_{t_{*}}\quad\text{or/and}\quad\hat{A}(i^{*},j^{*})-\hat{A}(i^{*},j)>\varepsilon_{t_{*}}.$
That is, under the estimated payoff matrix $\hat{A}$, there exists an action $i$ that is estimated to yield a higher payoff than $i^{*}$ for the row player, and an action $j$ that is estimated to be more favorable than $j^{*}$ for the column player.
The optimal action pair can only be eliminated by a suboptimal action pair $(i,j)$ in round $t_{*}$ such that $t_{ij}\geq t_{*}$ since action pair $(i,j)$ must remain under consideration at the elimination round of the optimal action pair. This condition also implies that $\hat{\Delta}_{t_{ij}}\leq\hat{\Delta}_{t_{*}}$ because of the decreasing behavior of $\hat{\Delta}_{t}$ across rounds in the algorithm. Moreover, it is important to note that the algorithm keeps this action pair $(i,j)$ for at least one additional round following the elimination of the optimal pair as it is identified as a NE. We further observe that all action pairs $(i,j)$ with $t_{ij}<t_{*}$ have already been eliminated in round $t_{ij}$ or earlier, which we already consider this condition in the previous case.
On the other hand, for the action pairs in $S_{A_{2}}$, we assume that $\hat{\Delta}_{t}<\frac{\lambda}{2}$, which implies that a suboptimal action pair remains in the game until the final round. This scenario emphasizes the regret contribution when the estimated suboptimality gap $\hat{\Delta}_{t}$ becomes sufficiently small, a case represented by the action pairs in $S_{A_{2}}$. Consequently, these suboptimal action pairs contribute to the overall regret and the analysis accounts for their effect to establish accurate performance guarantees.
If the algorithm keeps a suboptimal action pair $(i,j)$ up to round $t_{*}$, the events specified in (46) must hold with parameter $\varepsilon_{t_{*}}$. Consequently, by applying (52), the probability of keeping a suboptimal action pair can be bounded by $\frac{4N\sigma^{2}}{\hat{\Delta}_{t_{*}}^{2}T}$. Thus, we can write the regret contribution of this case as
$\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{16\sigma^{2}}{\hat{\Delta}_{t_{*}}^{2}T}$ $<\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}+1}^{2}T}$ (57)
$=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{64\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}T}$ (58)
$\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\frac{512\sigma^{2}}{\Delta_{ij}^{2}}+\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}$ (59)
where $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ and we apply (45). We use $\hat{\Delta}_{t+1}$ because the optimal action pair might be eliminated at some round $t_{*}$ such that $t_{*}\in[0,\max_{ij}t_{ij}]$. This implies that the optimal action pair can be eliminated by any $(i,j)$ no later than the last round that $(i,j)$ remains in the game. However, we must keep $(i,j)$ at least until the next round, since it is not removed in round $t_{*}$ and is selected as a misidentified NE.
Case C.4.
A suboptimal action pair $(i,j)$ in the set $S_{A_{2}}$ remains in the game as a unique action pair in $S_{t_{ij}}$.
This implies that a suboptimal action pair is played during the committing phase of Algorithm 2 up to step $T$. Thus, we need to account for an additional regret contribution term given by
$\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T.$ (60)
This regret contribution accounts for scenarios where the algorithm may eliminate all action pairs except one. If the remaining action pair is suboptimal, the algorithm will continue to play with this action pair for the remaining rounds. Since the action pair selection persists up to time step $T$, resulting in the additional term in the regret bound.
We note that $\Delta_{ij}^{*}\leq\Delta_{ij}$, which ensures that the regret does not grow excessively, and this relationship allows us to analyze the regret in terms of the defined action sets. As a result, if we sum these regret contributions as mentioned, we can conclude the proof.
$\Box$
C.1 Proof of Theorem 5.2
We consider the regret approach consisting of the losses of both players by comparing their best choices to the action played. This regret measures how much worse a player performs compared to their optimal strategy if they had known the opponent’s behavior in advance. Clearly we can write the followings:
$R_{T}$ $=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}]+\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}]$
$=\sum_{(i,j)\in S_{A}}\Delta_{ij}\mathbb{E}[n_{ij,T}]$
$=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}\mathbb{E}[n_{ij,T}]+\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}\mathbb{E}[n_{ij,T}]$
This implies that the term involving $\Delta_{ij}^{\text{max}}$ characterizes the regret incurred by the maximizing player, whereas the loss of the minimizing player is determined using $\Delta_{ij}^{\text{min}}$. In particular, we utilize the same underlying algorithm, which maintains the same uniform playing strategy and elimination strategy throughout the process. This consistency allows us to to leverage many elements of the previous regret analysis. Consequently, we can apply the same case scenarios from the proof of Theorem 5.1 as the elimination of action pairs follows the same strategy and the selection process of action pairs remains the same across rounds.
Although our current analysis involves a different notion of regret, consisting of $\Delta_{ij}$, this does not affect the underlying structure of the regret analysis. This ensures that the previous theoretical bounds can be extended by using $\Delta_{ij}$ instead of $\Delta_{ij}^{*}$ and we apply $\Delta_{ij}^{*}\leq\Delta_{ij}$ and $\Delta_{ij}\leq\lambda$ for the action pairs in $S_{A_{2}}$, then the result follows.
$\Box$
Appendix D Experimental Details
In this section, we present additional experiments to compare the performances of the algorithms based on their theoretical bounds and cumulative regret. To calculate the cumulative regret, we use the absolute difference between the value of the game and the payoff of the action played as the regret might otherwise be negative.
In Figure 1(b), we compare the performance of the ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF [18] algorithms. In each run, the exploration time $k$ for each action pair in ETC-TPZSG is randomly selected from a predefined list ranging from 100 to 2500 in increments of 100. The total number of rounds is set to $T=10^{4}$. On the other hand, each run uses a new game matrix with a unique pure NE which is generated using Gaussian payoffs with mean zero, standard deviation $\sigma$ selected from the list $[0.25,0.5,0.75,1]$. Varying $\sigma$ not only changes the payoff distribution but also affects the exploration time and the tolerance parameter $\varepsilon$ used in the elimination phase of ETC-TPZSG-AE. The cumulative regret is averaged over $10^{3}$ simulation runs. As observed, the ETC-TPZSG-AE algorithm consistently outperforms ETC-TPZSG, demonstrating the effectiveness of its elimination strategy in reducing cumulative regret, while Tsallis-INF perform better when compared to ETC-based algorithms. Notably, the ETC-based algorithms offer the advantage of conceptual and implementational simplicity.
Figure 2: Theoretical expected regret bound comparison between ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF from [18] using different $\Delta$ values
On the other hand, Figure 2 compares the theoretical expected regret bounds (in $O(\cdot)$ notation) of the algorithms across different values of the suboptimality gap $\Delta$. The results show that Tsallis-INF tends to perform better when $\Delta$ is large while the ETC-TPZSG-AE algorithm achieves lower regret for smaller values of $\Delta$.

### v2

@@
 Abstract
-We study a two-player zero-sum game (TPZSG) in which the row player aims to maximize their payoff against an adversarial column player, under an unknown payoff matrix estimated through bandit feedback.
-We propose and analyze two algorithms: ETC-TPZSG, which directly applies ETC to the TPZSG setting and ETC-TPZSG-AE, which improves upon it by incorporating an action pair elimination (AE) strategy that leverages the $\varepsilon$-Nash Equilibrium property to efficiently select the optimal action pair.
-Our objective is to demonstrate the applicability of ETC in a TPZSG setting by focusing on learning pure strategy Nash Equilibrium.
-A key contribution of our work is a derivation of instance-dependent upper bounds on the expected regret for both algorithms, has received limited attention in the literature on zero-sum games.
-Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC-TPZSG and $O(\frac{\log(T\Delta^{2})}{\Delta})$ for ETC-TPZSG-AE, where $\Delta$ denotes the suboptimality gap.
-Therefore, our results indicate that ETC-based algorithms perform effectively in adversarial game settings, achieving regret bounds comparable to existing methods while providing insights through instance-dependent analysis.
+We study a two-player zero-sum game in which the row player aims to maximize their payoff against an adversarial column player, under an unknown payoff matrix estimated through bandit feedback.
+We propose three algorithms based on the Explore-Then-Commit framework.
+The first adapts it to zero-sum games, the second incorporates adaptive elimination that leverages the $\varepsilon$-Nash Equilibrium property to efficiently select the optimal action pair, and the third extends the elimination algorithm by employing non-uniform exploration.
+Our objective is to demonstrate the applicability of ETC in a zero-sum game setting by focusing on learning pure strategy Nash Equilibria.
+A key contribution of our work is a derivation of instance-dependent upper bounds on the expected regret of our proposed algorithms, which has received limited attention in the literature on zero-sum games.
+Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC in zero-sum game setting and $O(\log(T\Delta^{2})/\Delta)$ for the adaptive elimination algorithm and its variant with non-uniform exploration, where $\Delta$ denotes the suboptimality gap.
+Therefore, our results indicate that ETC-based algorithms perform effectively in adversarial game settings, achieving regret bounds comparable to existing methods while providing insight through instance-dependent analysis.
+Keywords zero-sum games, regret analysis, bandit feedback
 1 Introduction
-This paper focuses on finite two-player zero-sum games (TPZSGs), where two players interact in an adversarial setting with strictly opposing objectives, as first analyzed in v1928theorie ; vonneumanandmorgenstern .
+This paper focuses on finite two-player zero-sum games (TPZSGs), where two players interact in an adversarial setting with strictly opposing objectives, as first analyzed in [31, 32].
 These games play a foundational role in game theory and online learning, capturing adversarial interactions in various applications such as economic competition and adversarial machine learning.
 In such settings, players repeatedly select actions from finite sets, with payoffs determined by a fixed but unknown game matrix.
-Studies on learning in games typically use either full information feedback where players observe the opponent’s actions or the entire payoff matrix after each round or partial feedback, also called bandit feedback, where each player observes only their own payoffs without access to the opponent’s strategy or the full matrix blum2007external .
-In this paper, we focus specifically on a TPZSG setting with bandit feedback and propose two algorithms, ETC-TPZSG and ETC-TPZSG-AE, designed to minimize regret by learning pure strategies for both players through balancing of exploration and exploitation.
+Studies on learning in games typically use either full information feedback where players observe the opponent’s actions or the entire payoff matrix after each round or partial feedback, also called bandit feedback, where each player observes only their own payoffs without access to the opponent’s strategy or the full matrix [7].
+In this paper, we study a TPZSG setting with bandit feedback, where players observe each other’s actions, by learning pure strategies for both players through balancing of exploration and exploitation.
 Several existing studies have advanced our understanding of how players learn to play optimally in zero-sum games with bandit feedback.
-The variant of UCB algorithm for adversarial games under bandit feedback have been analyzed using worst-case regret bounds in o2021matrix .
+The variant of UCB algorithm for adversarial games under bandit feedback has been analyzed using worst-case regret bounds in [27].
 In particular, worst-case analyses lead to all games as equally hard by ignoring the specific structure of the game.
 Instance-dependent analysis adapts the regret guarantees to the specific properties of a game to provide more practical performance guarantees.
-More recently, instance-dependent regret bounds have been investigated for the Tsallis-INF algorithm in TPZSGs ito2025instance .
-However an understanding of instance-dependent regret in TPZSGs under bandit feedback remains incomplete.
+More recently, instance-dependent regret bounds have been investigated for the Tsallis-INF algorithm in TPZSGs [18].
+However, an understanding of instance-dependent regret in TPZSGs under bandit feedback remains incomplete.
 Contributions.
@@
 While ETC has been widely studied in standard multi-armed bandit (MAB) problems, there has been no prior work specifically analyzing its performance to adversarial game settings, which makes this the first work to provide its detailed analysis in such a setting.
-A key motivation for studying ETC is its algorithmic simplicity allowing for a clear structure.
-In particular, we propose two algorithms that adapt the ETC approach to TPZSG setting and present their instance-dependent regret analyses.
-The first one, ETC-TPZSG, follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure strategy, achieves an upper bound of $\tilde{O}(\Delta+\sqrt{T})$ The second algorithm, ETC-TPZSG-AE, incorporates an adaptive elimination approach, similar to auer2010ucb , by sequentially eliminating actions which, with high probability, are not part of an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), gives us an upper regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$.
-Intuitively, this design is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action spaces of the players.
-Consequently, despite the simplicity of ETC framework, our results demonstrate that ETC-based approaches can be highly effective in adversarial settings.
-In both cases, the regret bounds depend on a suitable $\Delta$ notion, which varies depending on the type of the regret we are analyzing.
+A key motivation for studying ETC is its algorithmic simplicity, which allows for a clear structure.
+In particular, we propose three algorithms that adapt the ETC approach to the TPZSG setting and present their instance-dependent regret analyses.
+The first one follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure Nash Equilibria (NE), achieves an upper bound of $O(\Delta+\sqrt{T})$ .
+The second algorithm incorporates an adaptive elimination (AE), similar to [6], by sequentially eliminating actions that, with high probability, are not part of an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), gives us an upper regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$.
+Intuitively, this design is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action sets of the players.
+Finally, the last algorithm extends the adaptive elimination approach through non-uniform exploration, which also achieves a regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$.
+Consequently, despite the simplicity of the ETC framework, our results demonstrate that ETC-based approaches can be highly effective in adversarial settings.
 Paper structure.
-This paper is organized as follows.
 Section 2 reviews related work.
-Section 3 introduces preliminaries and formalizes the two-player zero-sum game (TPZSG) setting.
-In Section 4, we present the ETC in a TPZSG setting, ETC-TPZSG, and analyze its regret.
-Section 5 introduces an elimination based algorithm, ETC-TPZSG-AE, that leverages the $\varepsilon$-NE property, and provides its regret analysis.
-Section 6 offers empirical results validating their theoretical performances.
+Section 3 introduces preliminaries and formalizes a two-player zero-sum game setting.
+In Section 4, we present the ETC in a TPZSG setting and analyze its regret.
+Section 5 introduces an adaptive elimination algorithm and provides its regret analysis.
+In Section 6, we extend the elimination algorithm by using the non-uniform exploration approach and analyze its regret bound.
+Section 7 offers empirical results to validate their theoretical performances.
 Finally, Section 8 discusses our conclusions and future research directions.
 2 Related Work
-The multi-armed bandit (MAB) problem, has been introduced by thompson1933likelihood and originally described by robbins1952some , has been widely explored and has become a crucial framework for online decision-making under uncertainty, as explored in works such as audibert2009exploration ; auer2002finite ; auer1995gambling ; auer2010ucb ; blum2007learning ; bubeck2012regret .
-The MAB problem models scenarios in which a player must repeatedly choose from a set of actions to maximize their own rewards while balancing exploration-exploitation tradeoff, has formalized by lai1985asymptotically and further explored in studies such as audibert2007tuning ; maes2012learning .
-A central challenge is identifying the action with the highest expected reward while minimizing suboptimal choices, requiring efficient exploration and statistical confidence.
-Several approaches guarantee low error probabilities within finite trials, aiming to minimize regret audibert2010best ; gabillon2012best .
-Since the player has no prior knowledge of expected rewards, a dedicated exploration phase is essential, has motivated numerous studies such as bubeck2009pure ; neu2015explore .
-The Explore-Then-Commit (ETC) algorithm, which has been introduced by perchet2016batched , is one of the fundamental approaches in the MAB setting.
-It explores each action a fixed number of times, then commits to the best-performing action for the remaining rounds.
-Due to its simplicity and effectiveness, variants of the ETC algorithm are widely used in various decision-making scenarios garivier2016explore ; yekkehkhany2019etc .
-Zero-sum games, are a fundamental concept in game theory.
-They describe competitive scenarios where one player’s gain exactly equals the other’s loss.
-First highlighted by v1928theorie ; vonneumanandmorgenstern , their theory and applications have been extensively explored blum2007learning ; cesa2006prediction ; washburn2014two .
-The Minimax Theorem in v1928theorie ; vonneumanandmorgenstern guarantees equilibrium existence in TPZSGs, but finding equilibria becomes significantly harder when the payoff matrix is unknown daskalakis2011near ; daskalakis2009complexity ; mazumdar2019finding ; zinkevich2007new .
-A more general concept are Nash Equilibria nash1950equilibrium ; nash1951non .
-While these are identical to the minimax equilibria in TZPSGs, they are not as easy to compute in general games daskalakis2009complexity .
-The concept of $\varepsilon$-Nash Equilibrium provides an approximate solution where each player’s strategy is within an $\varepsilon$ tolerance level.
-daskalakis2009note used the $\varepsilon$-NE concept to efficiently approximate equilibrium in two-player games.
-In a zero-sum game setting with bandit feedback, players engage in repeated interactions without full knowledge of the payoff matrix, making it essential to estimate the unknown payoffs by previous observations.
-Various approaches, such as UCB o2021matrix and Tsallis-INF ito2025instance , have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret.
+The multi-armed bandit (MAB) problem, has been introduced by [30] and originally described by [29], has been widely explored and has become a crucial framework for online decision-making under uncertainty, as explored in works such as [3, 4, 5, 6, 8, 9].
+The MAB problem models scenarios in which a player must repeatedly choose from a set of actions to maximize their own rewards while balancing exploration-exploitation tradeoff, has been formalized by [19] and further explored in studies such as [2, 21].
+A central challenge is identifying the action with the highest expected reward while minimizing suboptimal choices, which require efficient exploration and statistical confidence.
+Several approaches guarantee low error probabilities within finite trials, in order to minimize regret [1, 16].
+Since the player has no prior knowledge of the expected rewards, a dedicated exploration phase is essential, which has motivated numerous studies such as [10, 26].
+The Explore-Then-Commit (ETC) algorithm, which has been introduced by [28], is one of the fundamental approaches in the MAB setting.
+It explores each action a fixed number of times, then commits to the best performing action for the remaining rounds.
+Due to its simplicity and effectiveness, variants of the ETC algorithm are widely used in various decision-making scenarios [17, 34].
+Zero-sum games represent a core concept in game theory.
+They describe competitive scenarios in which the gain of one player is exactly equal to the loss of the other.
+First highlighted by [31, 32], their theory and applications have been extensively explored [8, 12, 33].
+The Minimax Theorem in [31, 32] guarantees the existence of equilibrium in TPZSGs, but finding equilibrium becomes significantly harder when the payoff matrix is unknown [13, 14, 23, 35].
+Furthermore, the concept of Nash Equilibrium, which is introduced by [24, 25], describes optimal strategies of players in which no player is incentive to deviate.
+The study [14] shows that converging to such an equilibrium is challenging.
+The concept of $\varepsilon$-Nash Equilibrium provides an approximate solution where the strategy of each player is within a tolerance level $\varepsilon$.
+[15] uses this property to efficiently approximate equilibrium in two-player games.
+In a zero-sum game setting with bandit feedback, players engage in repeated interactions without full knowledge of the payoff matrix, which makes it essential to estimate the unknown payoffs by previous observations.
+Various approaches, such as UCB [27] and Tsallis-INF [18], have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret.
 Learning the game matrix is crucial for converging to equilibrium strategies and ensuring optimal decision-making in adversarial environments.
-o2021matrix focuses on worst-case regret analysis, while ito2025instance provides an instance-dependent regret analysis which achieves an upper bound of $O(\frac{\log T}{\Delta})$ in the case of a pure strategy equilibrium.
-Additionally, maiti2023instance investigates instance-dependent sample complexity bounds for zero-sum games under bandit feedback while cai2023uncoupled examines convergence rates to NE.
+Additionally, [22] investigates instance-dependent sample complexity bounds for zero-sum games under bandit feedback while [11] examines the convergence rates to NE.
+While [27] focuses on the worst-case regret analysis, [18] provides an instance-dependent regret analysis which achieves an upper bound of $O(\log_{+}(T\Delta^{2})/\Delta)$ in the case of a pure strategy equilibrium.
+Compared to [18], our work introduces more refined notions of suboptimality gaps for game instances, whereas their regret analysis focuses solely on the gap between the NE and the action chosen by a player when the opponent plays the equilibrium action.
+Since their regret bound involves the square of the sum of suboptimality gaps, our instance-dependent analysis gives a tighter bound.
 3 Preliminaries
-In this section, we define the basic concept of a TPZSG such as payoff matrix, the regret definitions we focus on and the gaps $\Delta$ that we use for the analysis.
+In this section, we define the basic concept of a TPZSG such as payoff matrix, the regret definitions we focus on, and the gaps $\Delta$ that we use for the analysis.
 In a TPZSG, the goals of the players are strictly opposed, which implies that any gain by one player corresponds to an equal loss by the other.
@@
 The outcome of such a game is fully determined by the strategies chosen by both players and the goal of each player is to maximize their own payoff while minimizing that of their opponent.
-Notation.
-Throughout this paper, $O(\cdot)$ denotes an upper bound up to constant factors.
 3.1 Two-Player Zero-Sum Games
 In a two-player zero-sum game, consisting of a row $(x)$ player and a column $(y)$ player, the row player aims to maximize their own payoff while the column player attempts to take an action to minimize this payoff.
-$S_{x}$ and $S_{y}$ are the action sets of the row player and column player, respectively.
+$S_{x}$ and $S_{y}$ are the action sets for the row player and column player, respectively.
 Thus, the action space of the game is $S_{A}=S_{x}\times S_{y}$, so that the action pairs $(i,j)\in S_{A}$ where $i\in S_{x},j\in S_{y}$.
-We also define $m=|S_{x}|$ as the number of actions for the row player, and $l=|S_{y}|$ as the number of actions for the column player, while finally $N=ml=|S_{A}|$ is the total number of action pairs.
-The game itself is defined through a game matrix $A$, so that if the players choose the pair $(i,j)$ then the row player obtains a payoff of $A(i,j)$ and the column player $-A(i,j)$.
+We also define $m=|S_{x}|$ as the number of actions for the row player and $l=|S_{y}|$ as the number of actions for the column player, while finally $N=ml=|S_{A}|$ is the total number of action pairs.
+The game itself is defined through a game matrix $A$, therefore if the players choose the pair $(i,j)$ then the row player obtains a payoff of $A(i,j)$ and the column player A(i,j)-A(i,j).
 The row player has a strategy (i.e. a probability distribution) $p$ to select an action $i\in S_{x}$, while the column player selects an action $j\in S_{y}$ according to a strategy $q$.
-Definition 3.1.
-A pair of mixed strategies $(p^{*},q^{*})$ is a Nash equilibrium (NE) if for all strategies $p$ and $q$, it satisfies
+Definition 1.
+A pair of mixed strategies $(p^{*},q^{*})$ is a Nash Equilibrium if for all strategies $p$ and $q$, it satisfies
 $V^{*}$ $\geq{p}^{\top}Aq^{*},$ $V^{*}$ $\leq{p^{*}}^{\top}Aq$ (1)
-where $V^{*}={p^{*}}^{\top}Aq^{*}$ is the value of the game nash1950equilibrium . It means that $p^{*}$ and $q^{*}$ are optimal strategies for row and column players, respectively.
-There always exists an optimal mixed strategy for both players that guarantees the best possible outcome by the Minimax Theorem v1928theorie ; vonneumanandmorgenstern .
-When a mixed strategy assigns a probability of one to a single action and zero to all others, it is referred to as a pure strategy.
+where $V^{*}={p^{*}}^{\top}Aq^{*}$ is the value of the game [24]. This means that $p^{*}$ and $q^{*}$ are optimal strategies for row and column players, respectively.
+There always exists an optimal mixed strategy for both players that guaranties the best possible outcome according to the Minimax Theorem [31, 32].
+When a mixed strategy assigns a probability of one to a single action and zero to all others, it refers to a pure strategy.
 In our setting, we focus on a TPZSG where both players adopt pure strategies.
-Definition 3.2 (Pure Nash Equilibria).
-A pair $(i^{*},j^{*})$ is a pure NE if:
-$i^{*}=\operatorname*{argmax}_{i\in S_{x}}\min_{j\in S_{y}}A(i,j)\quad\text{ and }\quad j^{*}=\operatorname*{argmin}_{j\in S_{y}}\max_{i\in S_{x}}A(i,j),$
+Definition 2 (Pure Nash Equilibria).
+A pair $(i^{*},j^{*})$ is a unique pure NE if:
+$i^{*}=\operatorname*{argmax}_{i\in S_{x}}\min_{j\in S_{y}}A(i,j)\quad\text{ and }\quad j^{*}=\operatorname*{argmin}_{j\in S_{y}}\max_{i\in S_{x}}A(i,j),$ (2)
 or equivalently
@@
 At time $t$, the row player draws an action $i_{t}\in S_{x}$ from a strategy $p_{t}$, and the column player draws $j_{t}\in S_{y}$ from a strategy $q_{t}$.
-The players then observe a noisy payoff $r_{t}$ with expected value $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, with the maximizing player obtaining $r_{t}$ and the minimizing player $-r_{t}$.
+The players then observe a noisy payoff $r_{t}$ with expected value $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, with the maximizing player obtaining $r_{t}$ and the minimizing player $r_{t}$.
 We assume that both players observe each other’s action.
 Since the algorithms do not know $A$, we use the notation $\hat{A}(i,j)$ to denote the estimated payoff when the action pair $(i,j)$ is played.
-We express the action pair played in round $t$ by $(i^{t},j^{t})$.
-After round $t$, we estimate each entry of the estimated game matrix by computing the average payoff for each action pair $(i,j)$ as
-$\hat{A}_{t}(i,j)=\frac{1}{n_{ij,t}}\sum_{s=1}^{t}\mathbb{I}\{(i^{s},j^{s})=(i,j)\}r_{s}$ (3)
+We express the action pair played in round $t$ by $(i_{t},j_{t})$.
+After round $t$, we estimate each entry in the estimated game matrix by computing the average payoff for each action pair $(i,j)$ as
+$\hat{A}_{t}(i,j)=\frac{1}{n_{ij,t}}\sum_{s=1}^{t}\mathbb{I}\{(i_{s},j_{s})=(i,j)\}r_{s}$ (4)
 where $n_{ij,t}$ is the number of times the action pair $(i,j)$ played until round $t$, $r_{s}$’s are independently and identically distributed (iid) $\sigma$-subgaussian payoffs.
-We assume that the estimator is unbiased, meaning that $\mathbb{E}[\hat{A}]=A$.
+We assume that the estimator is unbiased, which means that $\mathbb{E}[\hat{A}]=A$.
 3.3 Gap Notions
@@
 $\Delta_{ij}^{*}=A(i^{*},j^{*})-A(i,j)=V^{*}-A(i,j)$ (8)
-We note that $\Delta_{ij}^{\text{max}}\geq 0$ and $\Delta_{ij}^{\text{min}}\geq 0$, thus $\Delta_{ij}\geq 0$ for any action pair $(i,j)$. However, $\Delta_{ij}^{*}$ might be negative.
+We note that $\Delta_{ij}^{\text{max}}\geq 0$ and $\Delta_{ij}^{\text{min}}\geq 0$, therefore $\Delta_{ij}\geq 0$ for any action pair $(i,j)$. However, $\Delta_{ij}^{*}$ might be negative.
 Using the NE property in (3), for all $i\in S_{x}$ and $j\in S_{y}$ we can write the following:
-$A(i^{*},j^{*})\leq A(i^{*},j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)\implies A(i^{*},j^{*})-A(i,j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)-A(i,j)$ (8)
-$A(i^{*},j^{*})\geq A(i,j^{*})\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})\implies A(i^{*},j^{*})-A(i,j)\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})-A(i,j)$ (9)
+$A(i^{*},j^{*})\leq A(i^{*},j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)$
+which implies
+$A(i^{*},j^{*})-A(i,j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)-A(i,j).$ (9)
+Moreover,
+$A(i^{*},j^{*})\geq A(i,j^{*})\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})$
+implies the following:
+$A(i^{*},j^{*})-A(i,j)\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})-A(i,j).$ (10)
 Then, the inequality (9) implies that $\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$ and similarly from (10), we have $\Delta_{ij}^{*}\geq-\Delta_{ij}^{\text{min}}$.
-Therefore, we have $-\Delta_{ij}^{\text{min}}\leq\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$.
+Therefore, we have $\Delta_{ij}^{\text{min}}\leq\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$.
 On the other hand, since $\Delta_{ij}^{\text{min}}\geq 0$, we can write
@@
 3.4 Regret Notions
-To characterize the performance of our algorithms, we consider several regret notions. Let begin with the following formulations, which are the external regret of the max and min player respectively:
-$R_{T}^{\text{max}}=\max_{i\in S_{x}}\mathbb{E}\Big{[}\sum_{t=1}^{T}A(i,j^{t})-A(i^{t},j^{t})\Big{]}$
-$R_{T}^{\text{min}}=\max_{j\in S_{y}}\mathbb{E}\Big{[}\sum_{t=1}^{T}A(i^{t},j^{t})-A(i^{t},j)\Big{]}$
+To characterize the performance of our algorithms, we consider several regret notions.
+Let begin with the following formulations, which are the external regret of the max and min players, respectively:
+$R_{T}^{\text{max}}=\max_{i\in S_{x}}\mathbb{E}\Big[\sum_{t=1}^{T}A(i,j_{t})-A(i_{t},j_{t})\Big]$
+$R_{T}^{\text{min}}=\max_{j\in S_{y}}\mathbb{E}\Big[\sum_{t=1}^{T}A(i_{t},j_{t})-A(i_{t},j)\Big]$
 where $i^{t}$ and $j^{t}$ are the actions selected by the row and column player in round $t$, respectively.
-$R_{T}^{\text{max}}$ represents the expected loss due to not selecting the best possible action of row player against column player’s action $j$ while $R_{T}^{\text{min}}$ captures the expected regret from not choosing the column player’s best action for a given action $i$.
+$R_{T}^{\text{max}}$ represents the expected loss due to not selecting the best possible action of the row player against column player’s action $j$ while $R_{T}^{\text{min}}$ captures the expected regret of not choosing the column player’s best action for a given action $i$.
 We can now rewrite the regret definitions in terms of the suboptimality gaps:
-$R_{T}^{\text{max}}$ $=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}],$ $R_{T}^{\text{min}}$ $=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}],$ (11)
+$R_{T}^{\text{max}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}],$
+$R_{T}^{\text{min}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}],$
 where $\mathbb{E}[n_{ij,T}]$ denotes the expected number of times the action pair $(i,j)$ is played over $T$ rounds.
 In this paper, we analyze two regret notions. The external regret and the Nash regret.
-Definition 3.3 (External regret).
+Definition 3 (External regret).
 The (combined) external regret is given by the following:
@@
 External regret quantifies how much worse the players have performed compared to their best actions in hindsight.
-An alternative regret notion is the Nash regret, which measures the deviation of value from the Nash equilibrium:
-Definition 3.4 (Nash regret).
+An alternative regret notion is Nash regret, which measures the deviation of value from the Nash equilibrium:
+Definition 4 (Nash regret).
 The Nash regret is expressed as
@@
 However, if we are able to bound the Nash regret for the maximizing player, then by symmetry, we can also bound it for the minimizing player, thus effectively bounding its absolute value.
-4 ETC-TPZSG Algorithm and Regret Analysis
-In this section, we extend ETC algorithm to a two-person zero-sum game setting as ETC-TPZSG.
+4 ETC-TPZSG Algorithm
+In this section, we extend the ETC algorithm to a two-person zero-sum game setting, called ETC-TPZSG.
 Each player selects action from their corresponding finite action sets, which are known.
-The game is repeated over given time horizon $T$, and the true payoff matrix $A$ is unknown to both players.
+The game is repeated over a given time horizon $T$, and the true payoff matrix $A$ is unknown to both players.
 Instead, they receive observations of the payoffs based on their chosen actions, which are iid subgaussian random variables.
@@
 During the exploration phase, each player samples actions to estimate the expected payoffs, which enables us to construct an empirical payoff matrix using the observed payoffs.
-Then, in commit phase, they play according to the pure NE strategy derived from the estimated payoff matrix.
-After presenting the ETC-TPZSG algorithm, finally, we analyze its expected regret, measuring the performance due to limited information and suboptimal action pairs during the exploration phase.
-As a result, the unknown payoff matrix is approximated by the ETC approach, as in Lattimore_Szepesvari_2020 for a bandit setting.
-The ETC algorithm consists of two phases: an exploration phase, where the player randomly samples actions to gather information with given exploration time $k$, and a commit phase, where the player selects the empirically best action based on the gathered data.
+Then, in the commit phase, they play according to the pure NE strategy derived from the estimated payoff matrix.
+Finally, we analyze the expected regret of the algorithm, measuring the performance due to limited information and suboptimal action pairs during the exploration phase.
+As a result, the unknown payoff matrix is approximated by the ETC approach, as in [20] for a bandit setting.
+The ETC algorithm consists of two phases: an exploration phase, in which the player randomly samples actions to gather information with a given exploration time $k$, and a commit phase, where the player selects the empirically best action based on the observations.
 Algorithm 1 presents the implementation of ETC-TPZSG, where each player only observes the payoffs of the action pair they choose during the exploration period, then in the committing phase the algorithm converges to an optimal action pair which refers to the pure NE.
@@
 In bandit problems, after the exploration phase, a player simply commits to the arm with the highest estimated reward, which is typically effective if exploration is sufficient.
-However, in a zero-sum game with the pure NE, the objective is not to identify an action with high reward but to find a possibly true NE.
+However, in a zero-sum game with pure NE, the objective is not to identify an action with high reward but to find a possibly true NE.
 Inadequate exploration can lead to poor estimates of payoffs, making it difficult to correctly identify such an equilibrium.
-In fact, due to inaccurate estimates, a pure NE might not even appear to exist, even when one does in the true game.
-As a result, the design of ETC approach in zero-sum game setting requires careful balancing of exploration phase.
+As a result, the design of the ETC approach in the zero-sum game setting requires careful balance of the exploration phase.
 1:Input:
-2:$S_{A}$: set of action pairs in the game matrix $A$triangleright $S_{A}$ includes action pairs $(i,j)$
+[…]
 3:$m$: number of actions for row player
@@
 6:$T$: time horizon $\triangleright$ $1\leq mlk\leq T$
-7:$\sigma^{2}$: subgaussian variance factor
-8:Initialize: $\hat{A}=[0]^{m\times l}$triangleright $\hat{A}$ is estimated game matrix
-9:In round $t=0,1,2,...,mlk$; $\triangleright$ Exploration Phase
+[…]
+8:Initialize: $\hat{A}=[0]^{m\times l}$ $\triangleright$ $\hat{A}$ is estimated game matrix
+9:In round t=0,1,2,,mlkt=0,1,2,...,mlk; $\triangleright$ Exploration Phase
 10:    Explore each action pair $(i,j)$ in $S_{A}$ $k$ times
-11:        Update $\hat{A}$ using 3
-12:In round $t=mlk+1,mlk+2,...,T$; $\triangleright$ Committing Phase
-13:        Play an equilibrium $(i^{*},j^{*})$ satisfying the following property:
-$\hat{A}(i,j^{*})\leq\hat{A}(i^{*},j^{*})\leq\hat{A}(i^{*},j)$
+11:    Update $\hat{A}$ using (4)
+12:In round t=mlk+1,mlk+2,,Tt=mlk+1,mlk+2,...,T; $\triangleright$ Committing Phase
+13:    Play an equilibrium $(i^{*},j^{*})$ satisfying the following
+14:    property:
+$\hat{A}(i,j^{*})\leq\hat{A}(i^{*},j^{*})\leq\hat{A}(i^{*},j),\quad\forall i\in S_{x},j\in S_{y}$
 Algorithm 1 ETC-TPZSG
-Theorem 4.1.
+Theorem 1.
 The Nash regret of Algorithm 1, when interacting with $\sigma$-subgaussian payoffs, is upper bounded as follows:
@@
 where $k$ is the exploration time per action pair and $N$ is the total number of action pairs.
-We provide the proof in Appendix B.
-The regret comprises two main components, the exploration phase and the loss due to misidentifying the NE.
+Proof Sketch.
+The total regret consists of two components: the regret from exploration and the regret from exploitation.
+During the exploration phase, each action pair is played $k$ times.
+In the exploitation phase, the algorithm plays according to the pure equilibria, where we account for the probability that the pure NE is not correctly identified.
+In other words, in the exploitation phase, we analyze the loss incurred from playing with a misidentified NE.
+Thus, the expected number of plays for any suboptimal action pair $(i,j)$ can be written by the following:
+$\mathbb{E}[n_{ij,T}]=k+(T-Nk)\ \mathbb{P}((i,j)\text{ is identified as NE})$
+Once this probability is bounded, the proof is complete. A detailed derivation is provided in Appendix B. ∎
+As mentioned above, the regret comprises two main components, the exploration phase and the loss due to misidentifying the NE.
 Thus, it is crucial to choose an appropriate exploration time $k$ that balances sufficient exploration and efficient exploitation.
-Choosing $k$ too large leads to unnecessary exploration while setting it too small increases the risk of making suboptimal decisions.
+Choosing $k$ too large leads to unnecessary exploration, while setting it too small increases the risk of making suboptimal decisions.
 Hence, careful tuning of $k$ is essential to minimize overall regret.
-Let assume there are two action pairs $(i_{1},j)$ and $(i_{2},j)$ in the game such that row player has two actions and column player has one action.
+Let assume there are two action pairs $(i_{1},j)$ and $(i_{2},j)$ in the game such that the row player has two actions and the column player has one action.
 In addition, let suppose the NE is at $(i_{1},j)$.
 Then, we have $\Delta=A(i_{1},j)-A(i_{2},j)$ and we can write the regret simply as
-$R_{T}^{*}\leq k\Delta+(T-2k)\Delta\exp{\Big{(}-\frac{k\Delta^{2}}{16\sigma^{2}}\Big{)}}\leq k\Delta+T\Delta\exp{\Big{(}-\frac{k\Delta^{2}}{16\sigma^{2}}\Big{)}}.$ (15)
-Taking the first derivative with respect to $k$ and solve it, we get the following for the exploration time:
+$R_{T}^{*}$ $\leq k\Delta+(T-2k)\Delta\exp{\Big(-\frac{k\Delta^{2}}{16\sigma^{2}}\Big)}$ (15)
+$\leq k\Delta+T\Delta\exp{\Big(-\frac{k\Delta^{2}}{16\sigma^{2}}\Big)}.$ (16)
+Taking the first derivative with respect to $k$ and solving it, we get the following for the exploration time:
 $k=\max\Big\{1,\Big\lceil\frac{16\sigma^{2}}{\Delta^{2}}\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\rceil\Big\}$ (17)
@@
 $R_{T}^{*}\leq\min\Big\{T\Delta,\Delta+\frac{16\sigma^{2}}{\Delta}\Big(1+\max\Big\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\}\Big)\Big\}$ (18)
-where the first term $T\Delta$ is worst-case regret and for the other, we combine the regret we obtain in (15) with the value of exploration time in (16).
-Then, focusing on $\max\Big{\{}0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big{\}}$, we obtain that $\Delta=\frac{4\sigma}{\sqrt{T}}$ is a critical point.
+where the first term $T\Delta$ is the worst-case regret and for the other, we combine the regret we obtain in (16) with the value of the exploration time in (17).
+Then, focusing on $\max\Big\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\}$, we find that $\Delta=\frac{4\sigma}{\sqrt{T}}$ is a critical point.
 If we put it in the regret bound (18), we have the following:
-$R_{T}^{*}\leq\Delta+c\sqrt{T}$
-where $c>0$ is a some constant. If $\Delta\leq 1$, it becomes
-$R_{T}^{*}\leq 1+c\sqrt{T}.$ (18)
-Bounds like (18) are referred to as instance-independent since it only depends on the time horizon $T$ not on the game instance.
-In the zero-sum game setting, our analysis shows that we get a regret bound comparable to the standard bandit setting in Lattimore_Szepesvari_2020 , indicating that the ETC-based algorithm performs effectively in adversarial environments and aligns with known results in the literature.
-5 ETC-TPZSG-AE Algorithm and Regret Analysis
-If a pure NE exists at $(i^{*},j^{*})$, then it must satisfy the condition given in (2).
-For the algorithm in the section, we make use of the concept of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), which approximately satisfies the standard NE condition.
-Specifically, an $\varepsilon$-NE satisfies the NE conditions within a tolerance level $\varepsilon$, making it a suitable criterion in learning settings where exact equilibrium identification is hard.
-Based on this criterion, we introduce an elimination strategy in Algorithm 2, inspired by auer2010ucb , to reduce unnecessary exploration of clearly suboptimal action pairs.
+$R_{T}^{*}\leq\Delta+c\sqrt{T}$ (19)
+where $c>0$ is some constant.
+If $\Delta\leq 1$, it becomes $R_{T}^{*}\leq 1+c\sqrt{T}$, which refers to an instance-independent bound since it depends only on the time horizon $T$ not on the game instance.
+In the zero-sum game setting, our analysis shows that we get a regret bound similar to the standard bandit setting in [20], indicating that the ETC-based algorithm performs effectively in adversarial environments.
+5 ETC-TPZSG-AE Algorithm
+In this section, we extend Algorithm 1 by incorporating adaptive elimination, called ETC-TPZSG-AE.
+We use the concept of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), which approximately satisfies the standard NE condition, to eliminate action pairs.
+Specifically, if a pure NE exists at $(i^{*},j^{*})$, then it must satisfy the condition given in (3) while a $\varepsilon$-NE satisfies the NE conditions within a tolerance level $\varepsilon$, making it a suitable criterion in learning settings where exact equilibrium identification is challenging.
+Based on this criterion, we introduce an elimination method in Algorithm 2, inspired by [6], to reduce unnecessary exploration of clearly suboptimal action pairs.
 Specifically, action pairs that do not satisfy the $\varepsilon$-NE property are eliminated from further play, which enables a more efficient learning process by concentrating exploration on approximately optimal action pairs.
-Definition 5.1.
+Definition 5.
 The action pair $(i^{*},j^{*})$ is an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE) if the following condition holds:
 $A(i,j^{*})-\varepsilon\leq A(i^{*},j^{*})\leq A(i^{*},j)+\varepsilon,\quad\forall i\in S_{x},j\in S_{y}.$ (20)
-For the exploration time $k$ of the algorithm, let just go back to the regret bound in (14) and assume that there exists some $\Delta^{*}$ and $\Delta$ such that $\Delta^{*}\geq\Delta_{ij}^{*}$ and $\Delta\leq\Delta_{ij}$ for all $(i,j)\in S_{A}$.
-Then, when we put them into the regret bound and take its derivative with respect to $k$, we obtain a similar result as in (16); thus, as the exploration time, clearly we can use the following:
-$k=\Big{\lceil}\frac{16\sigma^{2}}{\Delta^{2}}\ln{\Big{(}\frac{\Delta^{2}T}{16\sigma^{2}}\Big{)}}\Big{\rceil}$ (20)
-Since the true suboptimality gaps $\Delta$ are unknown, the algorithm adopts a decreasing $\Delta$ approach across rounds.
+For the exploration time of the algorithm, let just go back to the regret bound in (14) and assume that there exist some $\Delta^{*}$ and $\Delta$ such that $\Delta^{*}\geq\Delta_{ij}^{*}$ and $\Delta\leq\Delta_{ij}$ for all $(i,j)\in S_{A}$.
+Then, when we put them into the regret bound and take its derivative with respect to $k$, we obtain a result similar to that of (17).
+Since the true suboptimality gaps $\Delta$ are unknown, the algorithm adopts a decreasing $\Delta$ approach over rounds.
 It is motivated by the intuition that the remaining action pairs become closer to the true NE as suboptimal action pairs are progressively eliminated.
 Consequently, the suboptimality gaps among the remaining action pairs are expected to decrease over time.
-In other words, as the rounds progress, we expect that the players get closer to true NE, which means they tend to repeat better action pairs more often.
-We can afford to use smaller $\Delta$ values to explore action pairs near the NE more thoroughly, since smaller $\Delta$ leads to a longer exploration time $k$.
-The algorithm schedules $\Delta_{ij}$ across rounds using $\hat{\Delta}_{t}$, which guides both exploration time and the tolerance level for the $\varepsilon$-NE condition.
-When $\hat{\Delta}_{t}$ is large, action pairs are played fewer times and more action pairs are kept during elimination.
-As $\hat{\Delta}_{t}$ decreases, the algorithm focuses exploration on pairs closer to equilibrium, having already eliminated clearly suboptimal ones.
+In other words, as the rounds progress, we expect that the players get closer to true NE, which means that they tend to repeat better action pairs more often.
+We can afford to use smaller $\Delta$ values to explore action pairs near the NE more thoroughly, since a smaller $\Delta$ leads to a longer exploration time $k$.
+The algorithm schedules $\Delta_{ij}$ across rounds using $\hat{\Delta}_{t}$, which guides both the exploration time and the tolerance level for the $\varepsilon$-NE condition.
+When $\hat{\Delta}_{t}$ is large, the action pairs are played fewer times and more action pairs are kept during elimination.
+As $\hat{\Delta}_{t}$ decreases, the algorithm focuses exploration on pairs closer to equilibrium and eliminates clearly suboptimal ones.
+Thus, as the exploration time, we can clearly use the following:
+$k_{t}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}\Big\rceil$ (21)
 In order to perform updates over rounds, an initial estimate is required.
-When the payoffs are bounded within a known interval, such as $[0,1]$, it is common practice to initialize the estimated suboptimality gap $\hat{\Delta}_{t}$ with value 1, representing the maximum possible difference between arm rewards as in auer2010ucb .
+When payoffs are bounded within a known interval, such as $[0,1]$, it is common practice to initialize the estimated suboptimality gap $\hat{\Delta}_{t}$ with value 1, representing the maximum possible difference between arm rewards as in [6].
 However, in our setting, the payoffs are assumed to be $\sigma$-subgaussian, which implies that there is no strict upper bound on the suboptimality gaps, we only have $\Delta_{ij}\geq 0$.
@@
 To address this, we initialize $\hat{\Delta}_{t}$ with $4\sigma$, which provides a practical starting point for the analysis.
-Moreover, we use $\hat{\Delta}_{t}=2^{-t+2}\sigma$ in each round $t$ which takes values from 00 to $\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$.
-Regarding $\Delta\approx\sqrt{(16\sigma^{2}/k)\ln{(\Delta^{2}T/16\sigma^{2})}}$ from (20) so $\ln{(\Delta^{2}T/16\sigma^{2})}\geq 0$, which gives us $\Delta\geq\sqrt{16\sigma^{2}/T}$.
+We set $\hat{\Delta}_{t}=2^{-t+2}\sigma$ in each round $t$, where $t$ ranges from 0 to $\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$.
+We have $\hat{\Delta}_{t}\approx\sqrt{(16\sigma^{2}/k_{t})\ln{(\hat{\Delta}_{t}^{2}T/16\sigma^{2})}}$ from (21), which implies $\ln{(\hat{\Delta}_{t}^{2}T/16\sigma^{2})}\geq 0$.
+It gives us $\hat{\Delta}_{t}\geq\sqrt{16\sigma^{2}/T}$ and our initialization of $\hat{\Delta}_{t}$ remains consistent with this observation.
 Then, using $2^{-t+2}\sigma\geq\sqrt{16\sigma^{2}/T}$ and after some calculations, we obtain $t\leq\frac{1}{2}\log_{2}T$.
-The division by $e$ is a technical adjustment, which makes the bound safer in the analysis; otherwise the number of play would be zero in the last round.
-If we assume that a unique pure NE exists in the game as before, a well estimated game matrix $\hat{A}$ should also exhibit the properties necessary to support the equilibrium.
+The division by $e$ is a technical adjustment, which makes the bound safer in the analysis; otherwise, the number of plays would be zero in the last round.
+Since we assume that a unique pure NE exists in the game, a well estimated game matrix $\hat{A}$ should also exhibit the properties necessary to support the equilibrium.
 Specifically, $\hat{A}$ must satisfy the conditions required for the $\varepsilon$-NE in (20) to hold.
-Based on this principle, the Algorithm 2 systematically eliminates action pairs that fail to satisfy this condition, as such pairs cannot be near to the equilibrium.
+Based on this principle, Algorithm 2 systematically eliminates action pairs that fail to satisfy this condition, as such pairs cannot be near to the equilibrium.
 By reducing the set of action pairs to include only the pairs that meet this criterion, it ensures that the remaining action pairs are the only ones that could potentially be near the NE; thus, it enables to focus on the most relevant action pairs.
 Consequently, we expect that the algorithm converges to an accurate approximation of the NE, even in the presence of estimation errors in $\hat{A}$.
+Theorem 2.
+The Nash regret of Algorithm 2, interacting with $\sigma$-subgaussian payoffs, is upper bounded by
+$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}+\Delta_{ij}^{*}T\Big)$ (22)
+where $\lambda\geq 4\sigma\sqrt{e/T}$, and $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$ are two subsets of the action pairs.
+Proof Sketch.
+To analyze the regret of Algorithm 2, at first we define $t_{ij}=\min\{t:\hat{\Delta}_{t}<\Delta_{ij}/2\}$ as the earliest round satisfying $\hat{\Delta}_{t}<\Delta_{ij}/2$ and then we consider four distinct cases.
+First, we analyze the suboptimal action pairs which are not eliminated in round $t_{ij}$ or before.
+Next, we examine the suboptimal pairs that are eliminated in the round $t_{ij}$ or earlier, allowing us to bound the number of plays with suboptimal action pairs.
+The third case examines the scenario where the optimal action pair is eliminated in $t_{ij}$.
+Finally, we consider the case in which a suboptimal action pair remains as a unique one in the game (discussed in detail in Appendix C).
+∎
 1:Input:
-2:$S_{A}$: set of action pairs in the game matrix $A$triangleright $S_{A}$ includes action pairs $(i,j)$
+[…]
 3:$m$: number of actions for row player
@@
 5:$T$: time horizon $\triangleright$ $1\leq ml\ll T$
-6:$\sigma^{2}$: subgaussian variance factor
-7:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$triangleright $\hat{A}$ is estimated game matrix
+[…]
+7:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$
 8:In round $t=0,1,2,...,\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$;
 9:    Action Pair Selection:
-10:            If $\mathbf{|S_{t}|>1}:$triangleright Exploration Phase
-11:              Explore each action pair $(i,j)$ in $S_{t}$ $k_{t}=\Big{\lceil}\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}\Big{\rceil}$ times
+10:     If $\mathbf{|S_{t}|>1}:$ $\triangleright$ Exploration Phase
+11:      Explore each $(i,j)$ in $S_{t}$ $k_{t}$ times, defined in (21)
+12:      Update $\hat{A}$ using (4)
+13:     Else: $\triangleright$ Committing Phase
+14:      Play with a unique action pair in $S_{t}$ until step $T$
+15:    Action Pair Elimination:
+16:     Remove the action pairs $(i,j)$ NOT satisfying the following for $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}}$ :
+$\hskip 35.56593pt\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
+17:    Update:
+18:     $S_{t+1}$ by "Action Pair Elimination" step
+19:     $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
+Algorithm 2 ETC-TPZSG-AE
+Moreover, if we consider another regret approach in (12), we obtain the following regret bound.
+Theorem 3.
+The external regret incurred by Algorithm 2 when interacting with $\sigma$-subgaussian payoffs is bounded as following:
+$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Big(\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}+\lambda T\Big).$ (23)
+Proof.
+Although we have a different notion of regret, this does not affect the underlying structure of the regret analysis.
+This ensures that the previous theoretical bounds can be extended using $\Delta_{ij}^{*}\leq\Delta_{ij}$, then the result follows.
+∎
+We simplify the regret bound to $O(\frac{\log(T\Delta^{2})}{\Delta})$, as the logarithmic term dominates since by setting $\lambda=\sqrt{16\sigma^{2}e/T}$, we ensure that the term $\lambda T$ is bounded by $\sqrt{16\sigma^{2}eT}$, which is at most $\frac{e}{\Delta_{ij}}$ when $\Delta_{ij}\leq\lambda$.
+Our results are consistent with the regret bound derived by [6].
+Compared to our regret bound $O(\frac{\log(T\Delta^{2})}{\Delta})$ with the bound $O(\frac{\log_{+}(T\Delta^{2})}{\Delta})$ in [18], we observe that our study introduces a more refined notion of the suboptimality gap $\Delta$.
+Specifically, in [18], $\Delta$ refers to the gap between pure NE and the action chosen when the opponent plays the action in equilibrium.
+Moreover, our regret bound is tighter, as their formulation involves the square of the sum of suboptimality gaps, while ours uses the sum of squares of suboptimality gaps within the logarithmic factor.
+Moreover, our elimination algorithm specifically incorporates the extended pure NE property through $\varepsilon$, which also allows us to explain why we achieve a tighter regret bound for a TPZSG setting where both players have a pure strategy.
+As a result, we have $R_{T}^{*}\leq R_{T}$, where $R_{T}^{*}$ evaluates its performance against the NE, capturing the cumulative loss relative to the equilibrium strategy, while the regret analysis of $R_{T}$ allows us to consider the losses of both players with respect to their best individual responses in each round, providing a broader and more flexible notion of regret.
+This outcome aligns with the interpretation that the expected regret against the NE establishes a more specific comparison, and thus, it naturally provides a tighter upper bound on the expected regret.
+6 ETC-TPZSG-AE-NUE Algorithm
+In this section, we extend the adaptive elimination algorithm by introducing a variant, ETC-TPZSG-AE-NUE.
+It uses a non-uniform exploration method The non-uniform exploration in these algorithms is based on players’ strategies, which indirectly reflect the probability of being NE.
+The key difference from the previous algorithm is that Algorithm 3 explores action pairs in proportion to the strategies of the players derived from the estimated payoff matrix.
+Thus, it is employed to reduce the number of plays for suboptimal action pairs more effectively.
+As mentioned in Section 3, the strategies of the row and column players are defined as probability distributions $p$ and $q$, respectively, over their corresponding action sets.
+In round $t$, using the estimated game matrix, we denote the estimated strategy of the row player by $\hat{p}_{t}$ and the estimated strategy of the column player by $\hat{q}_{t}$.
+Since the true game matrix is assumed to have the pure NE, we expect these estimated strategies to converge to the pure strategies over time.
+The estimated strategies of the players serve as a basis to determine the exploration time in our proposed algorithms.
+Furthermore, the probability of selecting an action pair $(i,j)$ is estimated by $\hat{P}_{ij,t}=\hat{p}_{i,t}\hat{q}_{j,t}$, which satisfies $\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=1$.
+Consequently, in Algorithm 3, the exploration time for each action pair $(i,j)$ is proportional to $\hat{P}_{ij,t}$, allowing action pairs with higher estimated probabilities to be played more frequently.
+We incorporate the strategies of the players $\hat{p}_{t}$ and $\hat{q}_{t}$ together with $k_{t}$.
+Specifically, in each round $t$, this involves estimating the optimal strategies for each player by solving a linear system using the estimated payoff matrix $\hat{A}$ as follows:
+$\begin{aligned} &\\ &\hat{P}_{t}=\hat{p}_{t}\hat{q}_{t}\\ &\hat{p}_{t}=\max_{p}\min_{q}p^{T}\hat{A}q\\ &\hat{q}_{t}=\min_{q}\max_{p}p^{T}\hat{A}q\\ \end{aligned}\qquad\begin{aligned} &\text{Subject to:}\\ &\sum_{i\in S_{x}}\hat{p}_{i,t}=1,\quad\hat{p}_{i,t}\geq 0,\ \forall i\in S_{x}\\ &\sum_{j\in S_{y}}\hat{q}_{j,t}=1,\quad\hat{q}_{j,t}\geq 0,\ \forall j\in S_{y}\\ &\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=\sum_{i,j}\hat{p}_{i,t}\hat{q}_{j,t}=1\end{aligned}$ (24)
+Thus, the exploration time is given by the following:
+$k_{ij,t}=k_{t}+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t}}{\hat{\Delta}_{t}^{2}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}\Big\rceil$ (25)
+where $k_{t}$ is defined in (21).
+Thus, the algorithm explores each action pair at least $k_{t}$ times.
+Since players’ strategies are expected to converge to pure ones over time, the Algorithm 3 progressively focuses on playing the optimal action pair more.
+This ensures that action pairs with higher estimated probabilities are explored more frequently, while those with lower probabilities are selected less often.
+Thus, these algorithms achieve a more efficient balance between exploration and exploitation by reducing unnecessary plays on suboptimal action pairs, which accelerates convergence toward optimal action pair.
+We note that we set $t=\left\lfloor\frac{1}{4}\log_{2}\frac{T}{e}\right\rfloor$ as the last round because unlike the Algorithm 2, which assigns an exploration time of $k_{t}$, the current algorithms allow exploration up to $2k_{t}$ times, ensuring that each action pair is explored at least $k_{t}$ times.
+As a result, due to this extended exploration phase, the number of rounds is reduced by half.
+1:Input:
+[…]
+3:$m$: number of actions for row player
+4:$l$: number of actions for column player
+5:$T$: time horizon $\triangleright$ $1\leq ml\ll T$
+[…]
+7:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$, $\hat{A}=[0]^{m\times l}$ and $P_{ij,0}=\frac{1}{ml},\forall i,j$
+8:In round $t=0,1,2,...,\left\lfloor\frac{1}{4}\log_{2}\frac{T}{e}\right\rfloor$;
+9:    Action Pair Selection:
+10:     If $\mathbf{|S_{t}|>1}:$ $\triangleright$ Exploration Phase
+11:      Explore each $(i,j)$ in $S_{t}$ $k_{ij,t}$ times, defined in (25)
 12:      Update $\hat{A}$ using (4).
@@
 16:     Remove the action pairs $(i,j)$ NOT satisfying the following for $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}}$:
-$\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
+$\hskip 35.56593pt\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
 17:    Update:
-18:            Reset $S_{t+1}$ by "Action Pair Elimination" step
-19:            $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
-Algorithm 2 ETC-TPZSG-AE
-Theorem 5.1.
-The Nash regret of Algorithm 2, interacting with $\sigma$-subgaussian payoffs, is upper bounded by
-$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big{(}1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big{(}\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big{)}}\Big{)}+\sum_{(i,j)\in S_{A_{2}}}\Big{(}\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}+\Delta_{ij}^{*}T\Big{)}$ (21)
-where $\lambda\geq\sqrt{\frac{16\sigma^{2}e}{T}}$ and, $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$ are two subsets of the action pairs.
-Moreover, if we consider another regret approach in (12), we obtain the following regret bound.
-Theorem 5.2.
-The external regret incurred by Algorithm 2 when interacting with $\sigma$-subgaussian payoffs is bounded as following:
-$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Big{(}\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\Big{(}\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big{)}}\Big{)}+\sum_{(i,j)\in S_{A_{2}}}\Big{(}\frac{512\sigma^{2}}{\lambda}+\lambda T\Big{)}$ (22)
-We simplify the regret bound to $O(\frac{\log(T\Delta^{2})}{\Delta})$, as the logarithmic term dominates since by setting $\lambda=\sqrt{\frac{16\sigma^{2}e}{T}}$, we ensure that the term $\lambda T$ is bounded by $\sqrt{16\sigma^{2}eT}$ which is at most $\frac{e}{\Delta_{ij}}$ when $\Delta_{ij}\leq\lambda$.
-Our results are consistent with the regret bound derived by auer2010ucb .
-The proofs of the theorems are provided in Appendix C.
-In this section, we analyze two distinct regret formulations for Algorithm 2, Nash regret, defined in (3.4), and external regret, defined in (3.3).
-The first, denoted by $R_{T}^{*}$, evaluates performance against the NE, capturing the cumulative loss relative to the equilibrium strategy.
-In contrast, the regret analysis by $R_{T}$ allows us to consider the losses of both players with respect to their individual best responses at each round, providing a broader and more flexible notion of regret.
-By the inequality $\Delta{ij}^{*}\leq\Delta{ij}$ for all action pairs, it follows that $R_{T}^{*}\leq R_{T}$.
-This outcome aligns with the interpretation that the expected regret against the NE establishes a more specific comparison and thus, it naturally provides a tighter upper bound on the expected regret.
+18:     $\hat{P}_{ij,t}$ by solving the linear system in (24).
+19:     $S_{t+1}$ by "Action Pair Elimination" step
+20:     $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
+Algorithm 3 ETC-TPZSG-AE-NUE
+Theorem 4.
+The Nash regret of Algorithm 3 is upper bounded by
+$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}+\Delta_{ij}^{*}T\Big)$ (26)
+where $\lambda\geq 4\sigma\sqrt[4]{e/T}$ and, $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$ are two subsets of the action pairs.
+Proof Sketch.
+Having the same structure as Algorithm 2 and ensuring that each action pair is explored at least $k_{t}$ times enables us to use the same proof steps; however, this modification affects the bound on the number of plays when the algorithm eliminates a suboptimal action pair in round $t_{ij}$.
+Detailed explanation is given in Appendix D.
+We note that since we have larger $\lambda$ here, it makes this bound tighter compared to the results of Algorithm 2.
+∎
+In conclusion, Algorithm 3 aims that the strategies of players converge to the corresponding pure strategies over time, which allows more focused exploration on actions near the pure NE and thus enables to eliminate suboptimal action pairs more effectively.
+In addition, the regret analysis of Algorithm 3 provides a tighter upper bound, highlighting the improved efficiency of the algorithm in learning pure NE.
 7 Experiments
 In this section, we present a set of simulations to support our theoretical findings and demonstrate the performance of the proposed algorithms.
-The experiments are conducted with fixed time horizons of $T=10^{3}$ and $T=10^{4}$.
-Additional details are provided in Appendix D.
-(a) The expected regret of ETC-TPZSG with $k$ in (16) and the upper bound in (17)
-(b) The cumulative regrets of ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF from ito2025instance
-(c) The theoretical regret bounds for ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF from ito2025instance
-Figure 1: Regret performance comparison of the algorithms
-Figure 1(a) shows the expected regret bound, averaged over $10^{5}$ simulation runs, of ETC-TPZSG using $k$ in (16) and the upper bound in (17).
-We consider a setting with $N=2$, where the row player has two actions, the column player has one, and the first row corresponds to the NE.
-The suboptimality gap $\Delta$ varies from 0 to 1 and the subgaussian parameter is set to $\sigma=0.5$.
-The results align with the bound in Theorem (4.1).
-In Figure 1(b), we compare ETC-TPZSG, ETC-TPZSG-AE, and Tsallis-INF ito2025instance , the only existing method with instance-dependent bounds for TPZSGs, based on cumulative regret, computed as the absolute difference between the game value and the payoff from the action played to avoid negative values, averaged over $10^{3}$ runs.
-The results show that ETC-TPZSG-AE achieves lower regret than ETC-TPZSG, highlighting the effectiveness of its action pair elimination strategy.
-We use a $2\times 2$ game matrix by Gaussian payoffs and $k$ exploration rounds per action pair for ETC-TPZSG is randomly selected from a predefined list.
-On the other hand, Figure 1(c) compares their theoretical regret bounds ($O(\cdot)$ notation) for $\Delta=0.5$, which aligns with the results in Figure 1(b).
+The experiments are conducted with a fixed time horizon $T=10^{3}$ for Figure 1 and $T=10^{4}$ for other simulations.
+Additional details are provided in Appendix E.
+Figure 1: The expected regret of ETC-TPZSG with $k$ in (17) and the upper bound in (18)
+Figure 2: The cumulative regrets of proposed algorithms and Tsallis-INF with large $\Delta$
+Figure 3: The cumulative regrets of proposed algorithms and Tsallis-INF with small $\Delta$
+Figure 1 shows the expected regret, averaged over $10^{5}$ runs, of ETC-TPZSG using $k$ in (17) and the upper bound in (18).
+We consider a setting with $N=2$, where the row player has two actions, the first corresponds to NE with mean zero and the second with $\Delta$, and the column player has one action.
+The suboptimality gap $\Delta$ varies from 0 to 1 and $\sigma=0.5$.
+The fluctuation around $\Delta=0.1$ arises since, up to this point, the exploration time $k$ equals one, after which it increases suddenly.
+The results align with the bound in Theorem 1.
+In Figures 3 and 3, we compare the proposed algorithms and Tsallis-INF [18], the only existing method with instance-dependent bounds for TPZSGs, based on cumulative regret, computed as the absolute difference between the game value and the payoff from the action played to avoid negative values, averaged over $10^{2}$ runs.
+We use a $2\times 2$ game matrix by Gaussian payoffs with variance $\sigma^{2}$ and $k$ exploration rounds per action pair for ETC-TPZSG is randomly selected from a predefined list.
+We note that a smaller $\sigma$ leads to a smaller $\Delta$.
+We evaluate the algorithms using the payoff matrices with large and small $\Delta$ in Figures 3 and 3, respectively, to examine how $\Delta$ affects the performance of the algorithms.
+This comparison highlights the importance of instance-dependent regret analysis.
+The results show that the algorithms incorporating elimination methods achieve lower regret than the ETC-TPZSG and Tsallis-INF algorithms, highlighting the effectiveness of adaptive elimination and exploration.
+On the other hand, when $\Delta$ is small, the Tsallis-INF algorithm fails to correctly identify pure NE as we see an increase in its cumulative regret in Figure 3.
+In contrast, Algorithms 2 and 3 perform better.
+Thus, we also emphasize that our algorithms preserve the pure NE property, which partly explains why they perform better than Tsallis-INF in games that admit pure NE.
+Experimental details and further results with different game matrices are provided in Appendix E.
 8 Discussion
-We investigate a TPZSG with bandit feedback, where the payoff matrix is unknown and must be learned through player interactions.
+We investigate a TPZSG setting with bandit feedback, where the true payoff matrix is unknown and must be learned through player interactions.
 This setting is challenging, as players must estimate payoffs while making strategic decisions in an adversarial environment.
 We adopt the ETC algorithm due to its simplicity, widespread use and lack of prior analysis in this context.
-We also integrate an elimination based method, allowing the systematic removal of suboptimal action pairs, thereby improving convergence to the equilibrium and reducing unnecessary exploration on suboptimal ones.
-A key contribution of our work is the derivation of instance-dependent upper bounds on the expected regret for both algorithms, which has received limited attention in the literature on zero-sum games.
-While our study focuses on pure strategy learning in a zero-sum game with bandit feedback and provides a theoretical expected regret bounds, several directions remain open for future study.
-One interesting extension is to consider games where the equilibrium is mixed.
-While the algorithms provided can be used to identify the support of the equilibrium, another approach should be used to efficiently converge to a mixed equilibrium.
+We integrate it to an adaptive elimination method, allowing the systematic removal of suboptimal action pairs, thereby improving convergence to the equilibrium and reducing unnecessary exploration on suboptimal ones.
+Finally, we combine the adaptive elimination algorithm with a non-uniform exploration approach to explore action pairs more effectively.
+A key contribution of our work is the derivation of instance-dependent upper bounds on the expected regret for proposed algorithms, which has received limited attention in the literature on zero-sum games.
+Although our study focuses on pure strategy learning in a zero-sum game with bandit feedback and provides instance-dependent regret bounds, several directions remain open for future study.
+An interesting extension is to explore games with a mixed strategy equilibrium, which is more challenging because of the convergence dynamics to true strategies, and also more complex to analyze regret compared to a pure strategy setting.
+Another direction is to analyze general-sum games, where the player interactions are not strictly adversarial, which requires to consider additional factors to learn optimal strategies.
 Finally, an important and relatively unexplored direction is fairness in zero-sum games.
@@
 This type of fair play mechanism could be essential in TPZSG settings, thus it can promote balanced exploration and improve overall strategy estimation.
-On the other hand, by integrating fairness based algorithms or constraints, it may be possible to generate game environments that ensure balanced opportunities for all players.
+On the other hand, by integrating fairness based algorithms or constraints, it may be possible to generate game environments that ensure equal opportunities for all players.
+Acknowledgments.
+The authors would like to thank Debabrota Basu, Thomas Kleine Buening and Ronald Ortner for fruitful discussions on this topic.
 Appendix A Subgaussian Properties
@@
 These results concern standard properties of subgaussian random variables such as tail inequalities.
-They are included here for completeness and easy reference since they play a critical role in our analyses.
+They are included here for completeness and easy reference as they play a critical role in our analyses.
 The statements are not original, we include only the parts that are directly useful for our purposes.
-Theorem A.1 ([20, Theorem 5.3]).
+Theorem 5 ([20, Theorem 5.3]).
 If $X$ is $\sigma$-subgaussian random variable, for any $\epsilon\geq 0$,
 $\mathbb{P}(X\geq\epsilon)\leq\exp{\Big(-\frac{\epsilon^{2}}{2\sigma^{2}}\Big)}.$ (27)
-Lemma A.1 ([20, Lemma 5.4]).
+Lemma 1 ([20, Lemma 5.4]).
 If $X$ is $\sigma$-subgaussian and $X_{1}$ and $X_{2}$ are independent with $\sigma_{1}$ and $\sigma_{2}$ subgaussian parameter, respectively, then we can write the followings:
   1. (a)
-    $cX$ is $|c|\sigma$-subgaussian for all $c\in\mathbb{R}.$
+    cXcX is $|c|\sigma$-subgaussian for all $c\in\mathbb{R}.$
   2. (b)
     $X_{1}+X_{2}$ is $\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}}$-subgaussian.
-Appendix B Proof of Theorem 4.1
-To analyze the Nash regret of the ETC-TPZSG algorithm, we begin by decomposing the total expected regret into two phases: the exploration phase and the committing phase.
+Appendix B Proof of Theorem 1
+To analyze the Nash regret of the Algorithm 1, we begin by decomposing the total expected regret into two phases: the exploration phase and the committing phase.
 During exploration, each action pair is played a fixed number of times $k$ to estimate their mean rewards, potentially incurring regret if suboptimal action pairs are played.
 Once the algorithm commits to the empirically best action pair, which is the pure NE, the regret accumulates only if this action pair is not the true optimal one.
-Our goal is to bound the expected regret by quantifying the probability of committing to a suboptimal action pair, which we can call it as a misidentified NE, based on the estimates of their payoffs obtained during exploration phase.
+Our goal is to bound the expected regret by quantifying the probability of committing to a suboptimal action pair, which we can call it as a misidentified NE, based on the estimates of their payoffs obtained during the exploration phase.
 Let $(i^{\prime},j^{\prime})$ denote a misidentified NE, so we aim to identify an action pair $(i^{\prime},j^{\prime})$, which appears better than $(i^{*},j^{*})$ based on the estimated matrix.
@@
 The choice between these approaches depends on which enables us the smaller upper bound.
-Let us start by considering the first approach. We can write it as
+Let us start by considering the first approach. We can write it as follows:
 $\mathbb{P}(E_{1}\cap E_{2})$ $=\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j),\forall i,j)$ (28)
@@
 $=\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)}$ (31)
-$\leq\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq\Delta_{i^{\prime}j^{\prime}}^{\text{max}})\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})}$ (28)
+$\leq\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})\leq\Delta_{i^{\prime}j^{\prime}}^{\text{max}})}$
+$\quad\times\sqrt{\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)\leq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})}$ (32)
 $\leq\sqrt{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\Big)}\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\Big)}}$ (33)
@@
 To get (30), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, that is, for any collection of events $e_{1},e_{2},\ldots,e_{n}$, we have $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{n})\leq\mathbb{P}(e_{i})$ for all $i\in{1,\ldots,n}$.
-The inequality (28) holds since $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}\geq 0,\Delta_{i^{\prime}j^{\prime}}^{\text{min}}\geq 0$, which are true from their definitions.
-For the step (29), we apply Theorem A.1 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable.
-At this point, we need to show that the differences $\hat{A}(i,j)-\hat{A}(x,y)$ are $\sqrt{2\sigma^{2}/k}$-subgaussian where $(i,j)$ and $(x,y)$ are any action pairs in $S_{A}$.
-To establish this, we utilize Lemma A.1, which provides key properties of subgaussian random variables.
-Specifically, both $\hat{A}(i,j)$ and $\hat{A}(x,y)$ are $\sqrt{\sigma^{2}/k}$-subgaussian as the average payoffs are computed by the equation (3) and during exploration phase, each action pair is played $k$ times.
+The inequality (32) holds by using the followings:
+$\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)=\mathbb{P}(\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})\leq A(i,j^{\prime})-A(i^{\prime},j^{\prime}))$ (38)
+$\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)=\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)\leq A(i^{\prime},j^{\prime})-A(i^{\prime},j))$ (39)
+Then, we apply $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}\geq A(i,j^{\prime})-A(i^{\prime},j^{\prime}),\Delta_{i^{\prime}j^{\prime}}^{\text{min}}\geq A(i^{\prime},j^{\prime})-A(i^{\prime},j)$, which are true from their definitions.
+For step (33), we apply Theorem 5 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable.
+At this point, we need to show that the differences $\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})$ and $\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)$ are $\sqrt{2\sigma^{2}/k}$-subgaussian but it is enough to check that for any action pairs $(i,j)$ and $(x,y)$ in $S_{A}$, $\hat{A}(i,j)-\hat{A}(x,y)$ are $\sqrt{2\sigma^{2}/k}$-subgaussian.
+To establish this, we utilize Lemma 1, which provides key properties of subgaussian random variables.
+Specifically, both $\hat{A}(i,j)$ and $\hat{A}(x,y)$ are $\sqrt{\sigma^{2}/k}$-subgaussian as the average payoffs are computed by the equation (4) and during the exploration phase, each action pair is played $k$ times.
 By the lemma, the difference of two independent subgaussian random variables with parameters $\sqrt{\sigma^{2}/k}$ is subgaussian with parameter $\sqrt{2\sigma^{2}/k}$, which follows that $\hat{A}(i,j)-\hat{A}(x,y)$ is $\sqrt{2\sigma^{2}/k}$-subgaussian for any action pairs $(i,j)$ and $(x,y)$.
-Then, to get the last inequality (31), we observe that
+To get the last inequality (35), we observe that
 $\frac{\Delta_{i^{\prime}j^{\prime}}^{2}}{2}=\frac{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{2}\leq(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}$
 where the inequality follows from the fact that $(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}-\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\geq 0$.
-As we mention before, there is an alternative approach to bound $\mathbb{P}(E_{1}\cap E_{2})$, which means to find a bound for the probability of a misidentified NE.
-It involves an inequality based on the minimum of the individual probabilities.
+As we mention before, there is an alternative approach to bound $\mathbb{P}(E_{1}\cap E_{2})$ to find a bound for the probability of a misidentified NE.
+It involves an inequality based on the minimum of individual probabilities.
 Similarly, let follow this:
@@
 $=\min\{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)\}$ (43)
-$\leq\min\{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq\Delta_{i^{\prime}j^{\prime}}^{\text{max}}),\mathbb{P}(\hat{A}(i^{\prime}j^{\prime})-\hat{A}(i^{\prime}j)\leq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})\}$ (38)
 $\leq\min\Big\{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\Big)},\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\Big)}\Big\}$ (44)
@@
 We already know that each action pair $(i,j)\in S_{A}$ is played at least $k$ times for exploration.
-After the exploration step, we will consider playing the misidentified NE $(i^{\prime},j^{\prime})$ instead of real one $(i^{*},j^{*})$ because we expect the players to select NE action pair as optimal after exploration phase.
+After the exploration step, we will consider playing the misidentified NE $(i^{\prime},j^{\prime})$ instead of true one $(i^{*},j^{*})$ because we expect the players to select NE action pair as optimal after the exploration phase.
 Since we already have the probability of playing a misidentified NE from (37) or (48), we can write
 $\mathbb{E}[n_{ij,T}]\leq k+(T-Nk)\exp{\Big(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big)}.$
-Hence, the proof is concluded.
-$\Box$
-Appendix C Proof of Theorem 5.1
-According to the algorithm, $\sqrt{\frac{16\sigma^{2}e}{T}}$ is a critical value for $\hat{\Delta}_{t}$ since it takes this value in the last round.
-In this way, let define a threshold parameter as $\lambda\geq\sqrt{\frac{16\sigma^{2}e}{T}}$, which gives us two subsets of actions pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$.
+Hence, the proof is concluded. $\Box$
+Appendix C Proof of Theorem 2
+According to the algorithm, $4\sigma\sqrt{e/T}$ is a critical value for $\hat{\Delta}_{t}$ since it takes this value in round $\frac{1}{2}\log_{2}\frac{T}{e}$.
+In this way, let define a threshold parameter as $\lambda\geq 4\sigma\sqrt{e/T}$, which gives us two subsets of actions pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$.
 Since $\Delta_{ij}^{*}\leq\Delta_{ij}$ ensures the bounds are preserved, it makes sense to use these action pair sets in the analysis.
@@
 where $\Delta_{ij}^{*}=A(i^{*},j^{*})-A(i,j).$ and $n_{ij,T}$ is the total number of times to play action pair $(i,j)$ by time $T$.
-To analyze the Nash regret of the ETC-TPZSG-AE algorithm, it is essential to consider various scenarios that may lead to suboptimal outcomes.
+To analyze the Nash regret of Algorithm 2, it is essential to consider various scenarios that may lead to suboptimal outcomes.
 Each of these scenarios contributes to the total Nash regret, as they either involve the incorrect elimination of the optimal action pair or the unnecessary selection of suboptimal pairs.
@@
 The algorithm employs a scheduling mechanism for $\Delta_{ij}$, where $\hat{\Delta}_{t}$ is halved in each round.
-This approach is necessary because the true values of $\Delta{ij}$ are unknown to the players.
+This approach is necessary because the true values of $\Delta_{ij}$ are unknown to the players.
 We use $\hat{\Delta}_{t}$ to determine both the exploration time and the tolerance level in the $\varepsilon$-NE property.
@@
 As $\hat{\Delta}_{t}$ decreases, the algorithm allocates more exploration to action pairs that are closer to being part of a NE, because those that are clearly suboptimal have already been eliminated.
-This adaptive process enables more precise identification of near-optimal strategies by $\varepsilon$-NE condition while minimizing regret.
+This adaptive process enables a more precise identification of near-optimal strategies by $\varepsilon$-NE condition while minimizing regret.
 For each suboptimal action pair $(i,j)$, let $t_{ij}=\min\{t:\hat{\Delta}_{t}<\Delta_{ij}/2\}$ refer to the earliest round such that $\hat{\Delta}_{t}<\Delta_{ij}/2$.
@@
 We now analyze the regret contributions by considering the following cases:
-Case C.1.
+Case 1.
 A suboptimal action pair $(i,j)$ is not eliminated in round $t_{ij}$ or earlier while the optimal action pair is in the set $S_{t_{ij}}$.
@@
 The regret contribution of this case comes from the fact that the algorithm spends time with a worse choice when a better one is already available.
-Let explain briefly in which situations the algorithm eliminates or keep an action pair.
+Let explain briefly in which situations the algorithm eliminates or keeps an action pair.
 For action pair $(i,j)$, if $\varepsilon$-NE property does not hold in round $t=t_{ij}$ which implies that at least one of the inequalities fails, then it will be eliminated in round $t_{ij}$.
 Furthermore, we have $\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\Big(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\Big)}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{t_{ij}+1}<\frac{\Delta_{ij}}{4}$.
-In other words, if there exist $i^{\prime}\in S_{x}$ or $j^{\prime}\in S_{y}$, such that one of the following is true:
+In other words, if there exists $i^{\prime}\in S_{x}$ or $j^{\prime}\in S_{y}$, such that one of the following is true:
 $I_{1}:\hat{A}(i^{\prime},j)-\hat{A}(i,j)>\varepsilon_{t_{ij}},\qquad I_{2}:\hat{A}(i,j)-\hat{A}(i,j^{\prime})>\varepsilon_{t_{ij}},$
@@
 Specifically, since we have $m$ actions in $S_{x}$, for any collection of events $e_{1},e_{2},\cdots,e_{m}$, it holds that $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{m})\leq\mathbb{P}(e_{i})$ for all $i=1,2,\cdots,m$.
-In the last inequality (49), we apply Theorem A.1 with $\sqrt{2\sigma^{2}/k_{t}}$-subgaussian random variable.
-The subgaussian parameter is derived by using Lemma A.1 and we note that each action pair is played $k_{t}$ times.
+In the last inequality (54), we apply Theorem 5 with $\sqrt{2\sigma^{2}/k_{t}}$-subgaussian random variable.
+The subgaussian parameter is derived by using Lemma 1 and we note that each action pair is played $k_{t}$ times.
 Similarly, in order to calculate the probability of the event $E_{2}$, we can write
@@
 This final pair is expected to correspond to the pure NE, representing the best choice for both players with no incentive to change their actions.
-Case C.2.
+Case 2.
 A suboptimal action pair $(i,j)$ is eliminated in round $t_{ij}$ or earlier with the optimal action pair in $S_{t_{ij}}$.
@@
 On the other hand, we note that there is no need to consider the action pairs $(i,j)\in S_{A_{2}}$ because we handle the elimination of suboptimal pairs before the final round.
-Thus, using (45), each action pair $(i,j)$ is played at most $k_{t_{ij}}$ times such that $k_{t_{ij}}=\Big{\lceil}\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big{\rceil}\leq\Big{\lceil}\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big{(}\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big{)}}\Big{\rceil}$.
+Thus, using (50), each action pair $(i,j)$ is played at most $k_{t_{ij}}$ times such that
+$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq\Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil.$
 Then, the regret contribution is expressed by
@@
 $=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}$ (60)
-Case C.3.
+Case 3.
 The optimal action pair $(i^{*},j^{*})$ is eliminated by some suboptimal one $(i,j)$ in round $t_{*}$ such that $t_{ij}\geq t_{*}$.
@@
 If the algorithm keeps a suboptimal action pair $(i,j)$ up to round $t_{*}$, the events specified in (51) must hold with parameter $\varepsilon_{t_{*}}$.
-Consequently, by applying (52), the probability of keeping a suboptimal action pair can be bounded by $\frac{4N\sigma^{2}}{\hat{\Delta}_{t_{*}}^{2}T}$.
+Consequently, by applying (57), the probability of keeping a suboptimal action pair can be bounded by $\frac{16\sigma^{2}}{\hat{\Delta}_{t_{*}}^{2}T}$.
 Thus, we can write the regret contribution of this case as
@@
 However, we must keep $(i,j)$ at least until the next round, since it is not removed in round $t_{*}$ and is selected as a misidentified NE.
-Case C.4.
-A suboptimal action pair $(i,j)$ in the set $S_{A_{2}}$ remains in the game as a unique action pair in $S_{t_{ij}}$.
+Case 4.
+A suboptimal action pair $(i,j)$ from the set $S_{A_{2}}$ is selected as the unique action pair in $S_{t_{ij}}$ for remaining rounds.
 This implies that a suboptimal action pair is played during the committing phase of Algorithm 2 up to step $T$. Thus, we need to account for an additional regret contribution term given by
 $\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T.$ (65)
-This regret contribution accounts for scenarios where the algorithm may eliminate all action pairs except one.
-If the remaining action pair is suboptimal, the algorithm will continue to play with this action pair for the remaining rounds.
-Since the action pair selection persists up to time step $T$, resulting in the additional term in the regret bound.
+This regret contribution accounts for scenarios where the algorithm selects suboptimal action pair as the pure NE.
+In other words, if the selected action pair is suboptimal, the algorithm will continue to play with this action pair for the remaining rounds.
+Since the action pair selection persists up to time step $T$, resulting in an additional term in the regret bound.
 We note that $\Delta_{ij}^{*}\leq\Delta_{ij}$, which ensures that the regret does not grow excessively, and this relationship allows us to analyze the regret in terms of the defined action sets.
@@
 $\Box$
-C.1 Proof of Theorem 5.2
-We consider the regret approach consisting of the losses of both players by comparing their best choices to the action played.
-This regret measures how much worse a player performs compared to their optimal strategy if they had known the opponent’s behavior in advance.
-Clearly we can write the followings:
-$R_{T}$ $=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}]+\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}]$
-$=\sum_{(i,j)\in S_{A}}\Delta_{ij}\mathbb{E}[n_{ij,T}]$
-$=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}\mathbb{E}[n_{ij,T}]+\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}\mathbb{E}[n_{ij,T}]$
-This implies that the term involving $\Delta_{ij}^{\text{max}}$ characterizes the regret incurred by the maximizing player, whereas the loss of the minimizing player is determined using $\Delta_{ij}^{\text{min}}$.
-In particular, we utilize the same underlying algorithm, which maintains the same uniform playing strategy and elimination strategy throughout the process.
-This consistency allows us to to leverage many elements of the previous regret analysis.
-Consequently, we can apply the same case scenarios from the proof of Theorem 5.1 as the elimination of action pairs follows the same strategy and the selection process of action pairs remains the same across rounds.
-Although our current analysis involves a different notion of regret, consisting of $\Delta_{ij}$, this does not affect the underlying structure of the regret analysis.
-This ensures that the previous theoretical bounds can be extended by using $\Delta_{ij}$ instead of $\Delta_{ij}^{*}$ and we apply $\Delta_{ij}^{*}\leq\Delta_{ij}$ and $\Delta_{ij}\leq\lambda$ for the action pairs in $S_{A_{2}}$, then the result follows.
-$\Box$
-Appendix D Experimental Details
+Appendix D Proof of Theorem 4
+We follow exactly the same steps as the proof of Theorem 2.
+We update the threshold parameter as $\lambda\geq 4\sigma\sqrt[4]{e/T}$, thus we have two subsets of action pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$.
+We can use the same probability of keeping an action pair as in (57) since the algorithm each action pair at least $k_{t}$ times.
+Then, if we keep the same events from the proof of Theorem 2, Appendix C, we have the following:
+$\mathbb{P}(E_{1})$ $=\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t},\forall i^{\prime}\in S_{x})$ (66)
+$\leq\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t})$ (67)
+$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2(\frac{\sigma^{2}}{k_{i^{\prime}j,t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)}$ (68)
+$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{i^{\prime}j,t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)}$ (69)
+$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2\sigma^{2}\Big(\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)}$ (70)
+$=\exp{\Bigg(-\frac{2(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}\Bigg)}$ (71)
+$\leq\exp{\Bigg(-\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}\Bigg)$ (72)
+$=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}$ (73)
+Since the term $\frac{2(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}$ is greater than one, the inequality (72) follows.
+In addition, since we obtain the same result for $\mathbb{P}(E_{2})$ as $\mathbb{P}(E_{2})$, which leads to the same probability of keeping an action pair as in the proof of Theorem 2 and also the same results for Case 1, Case 3.
+On the other hand, since we explore each action pair non-uniformly, we need to update Case 2 by using
+$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t_{ij}}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$
+since $\hat{P}_{ij,t_{ij}}\leq 1$ for all $i,j$. Then, the regret contribution becomes
+$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\ 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$ $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}$ (74)
+Then, it concludes the proof.
+Theorem 6.
+The external regret of Algorithm 3 is bounded by
+$R_{T}\leq$ $\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{512\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}+\lambda T\Big).$
+Since we have $\Delta_{ij}^{*}\leq\Delta_{ij}$, it concludes the proof.
+Appendix E Experimental Details
 In this section, we present additional experiments to compare the performances of the algorithms based on their theoretical bounds and cumulative regret.
 To calculate the cumulative regret, we use the absolute difference between the value of the game and the payoff of the action played as the regret might otherwise be negative.
-In Figure 1(b), we compare the performance of the ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF [18] algorithms.
-In each run, the exploration time $k$ for each action pair in ETC-TPZSG is randomly selected from a predefined list ranging from 100 to 2500 in increments of 100.
+In Figures 3 and 3 , we compare the performance of the ETC-TPZSG, ETC-TPZSG-AE, ETC-TPZSG-AE-NUE and Tsallis-INF [18] algorithms.
+In each run, the exploration time $k$ for each action pair in ETC-TPZSG is randomly selected from a predefined list ranging from 100 to 2500 in increments of 25.
 The total number of rounds is set to $T=10^{4}$.
-On the other hand, each run uses a new game matrix with a unique pure NE which is generated using Gaussian payoffs with mean zero, standard deviation $\sigma$ selected from the list $[0.25,0.5,0.75,1]$.
-Varying $\sigma$ not only changes the payoff distribution but also affects the exploration time and the tolerance parameter $\varepsilon$ used in the elimination phase of ETC-TPZSG-AE.
-The cumulative regret is averaged over $10^{3}$ simulation runs.
-As observed, the ETC-TPZSG-AE algorithm consistently outperforms ETC-TPZSG, demonstrating the effectiveness of its elimination strategy in reducing cumulative regret, while Tsallis-INF perform better when compared to ETC-based algorithms.
-Notably, the ETC-based algorithms offer the advantage of conceptual and implementational simplicity.
-Figure 2: Theoretical expected regret bound comparison between ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF from [18] using different $\Delta$ values
-On the other hand, Figure 2 compares the theoretical expected regret bounds (in $O(\cdot)$ notation) of the algorithms across different values of the suboptimality gap $\Delta$.
-The results show that Tsallis-INF tends to perform better when $\Delta$ is large while the ETC-TPZSG-AE algorithm achieves lower regret for smaller values of $\Delta$.
+On the other hand, each run uses a new game matrix with a unique pure NE which is generated using Gaussian payoffs with mean zero, standard deviation $\sigma$ selected from the list $[0.1,0.2]$ for Figure 3 and from the list $[0.25,0.5,0.75,1]$ for Figure 3.
+Here, $\sigma$ denotes the variance factor of Gaussian payoffs and setting it to a small value leads to smaller $\Delta$ values.
+Furthermore, varying $\sigma$ not only changes the payoff distribution but also affects the exploration time and the tolerance parameter $\varepsilon$ used in the elimination phase of ETC-TPZSG-AE.
+The cumulative regret is averaged over $10^{2}$ simulation runs.
+Furthermore, Figures 4 and 5 present the comparison of algorithms based on different game matrices.
+For large $\Delta$, although the Tsallis-INF algorithm performs better in the first $10^{3}$ rounds, over time, our adaptive elimination algorithms reach lower cumulative regret.
+As observed, the ETC-TPZSG-AE algorithm consistently outperforms ETC-TPZSG, demonstrating the effectiveness of its elimination strategy in reducing cumulative regret, and ETC-TPZSG-AE-NUE outperforms even for larger action sets.
+Increasing the number of actions for players makes the identification of the true NE more challenging, as reflected in the performance of ETC-TPZSG as shown in Figures 4 and 5.
+On the other hand, when $\Delta$ is smaller, Tsallis-INF [18] also has difficulty in identifying pure NE correctly, as seen in Figure 5.
+However, Algorithms 2 and 3 perform well even in these cases, as they incorporate pure NE properties during the elimination phase.
+In particular, ETC-based algorithms offer the advantage of conceptual and implementational simplicity.
+Figure 4: The cumulative regrets of the algorithms for different game matrices with large $\Delta$
+Figure 5: The cumulative regrets of the algorithms for different game matrices with small $\Delta$

### v3

@@
 We study a two-player zero-sum game in which the row player aims to maximize their payoff against an adversarial column player, under an unknown payoff matrix estimated through bandit feedback.
-We propose three algorithms based on the Explore-Then-Commit framework.
+We propose three algorithms based on the Explore-Then-Commit (ETC) framework.
 The first adapts it to zero-sum games, the second incorporates adaptive elimination that leverages the $\varepsilon$-Nash Equilibrium property to efficiently select the optimal action pair, and the third extends the elimination algorithm by employing non-uniform exploration.
@@
 Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC in zero-sum game setting and $O(\log(T\Delta^{2})/\Delta)$ for the adaptive elimination algorithm and its variant with non-uniform exploration, where $\Delta$ denotes the suboptimality gap.
-Therefore, our results indicate that ETC-based algorithms perform effectively in adversarial game settings, achieving regret bounds comparable to existing methods while providing insight through instance-dependent analysis.
+Therefore, our results indicate that ETC-based algorithms perform effectively in zero-sum game settings, achieving regret bounds comparable to existing methods while providing insight through instance-dependent analysis.
 Keywords zero-sum games, regret analysis, bandit feedback
@@
 Instance-dependent analysis adapts the regret guarantees to the specific properties of a game to provide more practical performance guarantees.
-More recently, instance-dependent regret bounds have been investigated for the Tsallis-INF algorithm in TPZSGs [18].
+More recently, instance-dependent regret bounds have been investigated in TPZSGs [18, 24].
 However, an understanding of instance-dependent regret in TPZSGs under bandit feedback remains incomplete.
@@
 This paper investigates the effectiveness of the Explore-Then-Commit (ETC) algorithm and extensions in the context of TPZSGs.
-While ETC has been widely studied in standard multi-armed bandit (MAB) problems, there has been no prior work specifically analyzing its performance to adversarial game settings, which makes this the first work to provide its detailed analysis in such a setting.
+While ETC has been widely studied in standard multi-armed bandit (MAB) problems, there has been no prior work specifically analyzing its performance to zero-sum game settings, which makes this the first work to provide its detailed analysis in such a setting.
 A key motivation for studying ETC is its algorithmic simplicity, which allows for a clear structure.
+Moreover, our main focus is to show whether ETC and action elimination algorithms is applicable to zero-sum games by offering a refined understanding of regret behavior in TPZSGs under bandit feedback.
 In particular, we propose three algorithms that adapt the ETC approach to the TPZSG setting and present their instance-dependent regret analyses.
 The first one follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure Nash Equilibria (NE), achieves an upper bound of $O(\Delta+\sqrt{T})$ .
-The second algorithm incorporates an adaptive elimination (AE), similar to [6], by sequentially eliminating actions that, with high probability, are not part of an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), gives us an upper regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$.
+The second algorithm incorporates an adaptive elimination (AE), similar to [6], by sequentially eliminating actions that, with high probability, are not part of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), gives us an upper regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$.
 Intuitively, this design is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action sets of the players.
@@
 Section 5 introduces an adaptive elimination algorithm and provides its regret analysis.
-In Section 6, we extend the elimination algorithm by using the non-uniform exploration approach and analyze its regret bound.
+In Section 6, we extend the elimination algorithm by using the non-uniform exploration approach and analyze its regret upper bound.
 Section 7 offers empirical results to validate their theoretical performances.
@@
 In a zero-sum game setting with bandit feedback, players engage in repeated interactions without full knowledge of the payoff matrix, which makes it essential to estimate the unknown payoffs by previous observations.
-Various approaches, such as UCB [27] and Tsallis-INF [18], have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret.
+Various approaches, such as UCB [29], Tsallis-INF [18] and a novel algorithm [24], have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret.
 Learning the game matrix is crucial for converging to equilibrium strategies and ensuring optimal decision-making in adversarial environments.
-Additionally, [22] investigates instance-dependent sample complexity bounds for zero-sum games under bandit feedback while [11] examines the convergence rates to NE.
-While [27] focuses on the worst-case regret analysis, [18] provides an instance-dependent regret analysis which achieves an upper bound of $O(\log_{+}(T\Delta^{2})/\Delta)$ in the case of a pure strategy equilibrium.
+Additionally, [23] and [22] investigate instance-dependent sample complexity bounds for zero-sum games under bandit feedback while [11] examines the convergence rates to NE.
+While [29] focuses on the worst-case regret analysis and [24] analyzes instance-dependent regret only for a $2\times 2$ game matrix, [18] provides an instance-dependent regret analysis which achieves an upper bound of $O(\log_{+}(T\Delta^{2})/\Delta)$ in the case of a pure strategy equilibrium.
 Compared to [18], our work introduces more refined notions of suboptimality gaps for game instances, whereas their regret analysis focuses solely on the gap between the NE and the action chosen by a player when the opponent plays the equilibrium action.
@@
 At time $t$, the row player draws an action $i_{t}\in S_{x}$ from a strategy $p_{t}$, and the column player draws $j_{t}\in S_{y}$ from a strategy $q_{t}$.
-The players then observe a noisy payoff $r_{t}$ with expected value $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, with the maximizing player obtaining $r_{t}$ and the minimizing player $r_{t}$.
+The players then observe a noisy payoff $r_{t}$ with expected value $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, where the maximizing player receives $r_{t}$ and the minimizing player receives $r_{t}$.
 We assume that both players observe each other’s action.
-Since the algorithms do not know $A$, we use the notation $\hat{A}(i,j)$ to denote the estimated payoff when the action pair $(i,j)$ is played.
+Since the players do not know $A$, we use the notation $\hat{A}(i,j)$ to denote the estimated payoff when the action pair $(i,j)$ is played.
 We express the action pair played in round $t$ by $(i_{t},j_{t})$.
-After round $t$, we estimate each entry in the estimated game matrix by computing the average payoff for each action pair $(i,j)$ as
+In each round $t$, the estimated game matrix is updated by computing the average payoff for each action pair $(i,j)$ as follows:
 $\hat{A}_{t}(i,j)=\frac{1}{n_{ij,t}}\sum_{s=1}^{t}\mathbb{I}\{(i_{s},j_{s})=(i,j)\}r_{s}$ (4)
 where $n_{ij,t}$ is the number of times the action pair $(i,j)$ played until round $t$, $r_{s}$’s are independently and identically distributed (iid) $\sigma$-subgaussian payoffs.
-We assume that the estimator is unbiased, which means that $\mathbb{E}[\hat{A}]=A$.
 3.3 Gap Notions
@@
 $R_{T}^{\text{min}}=\max_{j\in S_{y}}\mathbb{E}\Big[\sum_{t=1}^{T}A(i_{t},j_{t})-A(i_{t},j)\Big]$
-where $i^{t}$ and $j^{t}$ are the actions selected by the row and column player in round $t$, respectively.
+where $i_{t}$ and $j_{t}$ are the actions selected by the row and column player in round $t$, respectively.
 $R_{T}^{\text{max}}$ represents the expected loss due to not selecting the best possible action of the row player against column player’s action $j$ while $R_{T}^{\text{min}}$ captures the expected regret of not choosing the column player’s best action for a given action $i$.
 We can now rewrite the regret definitions in terms of the suboptimality gaps:
-$R_{T}^{\text{max}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}],$
-$R_{T}^{\text{min}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}],$
+$R_{T}^{\text{max}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}]$
+$R_{T}^{\text{min}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}]$
 where $\mathbb{E}[n_{ij,T}]$ denotes the expected number of times the action pair $(i,j)$ is played over $T$ rounds.
@@
 The ETC algorithm consists of two phases: an exploration phase, in which the player randomly samples actions to gather information with a given exploration time $k$, and a commit phase, where the player selects the empirically best action based on the observations.
-Algorithm 1 presents the implementation of ETC-TPZSG, where each player only observes the payoffs of the action pair they choose during the exploration period, then in the committing phase the algorithm converges to an optimal action pair which refers to the pure NE.
+Algorithm 1 presents the implementation of ETC-TPZSG, where each player only observes the payoffs of the action pair they choose during the exploration period, then in the commit phase, the algorithm selects an optimal action pair which refers to the pure NE.
 Applying ETC in a zero-sum game setting presents some challenges compared to the standard MAB scenario.
@@
 As a result, the design of the ETC approach in the zero-sum game setting requires careful balance of the exploration phase.
-1:Input:
-[…]
-3:$m$: number of actions for row player
-4:$l$: number of actions for column player
-5:$k$: exploration time
-6:$T$: time horizon $\triangleright$ $1\leq mlk\leq T$
-[…]
-8:Initialize: $\hat{A}=[0]^{m\times l}$ $\triangleright$ $\hat{A}$ is estimated game matrix
-9:In round t=0,1,2,,mlkt=0,1,2,...,mlk; $\triangleright$ Exploration Phase
-10:    Explore each action pair $(i,j)$ in $S_{A}$ $k$ times
-11:    Update $\hat{A}$ using (4)
-12:In round t=mlk+1,mlk+2,,Tt=mlk+1,mlk+2,...,T; $\triangleright$ Committing Phase
-13:    Play an equilibrium $(i^{*},j^{*})$ satisfying the following
-14:    property:
+Algorithm 1 ETC-TPZSG
+1:Input: $S_{A},m,l,k,\sigma,T(1\leq mlk\leq T)$
+2:Initialize: $\hat{A}=[0]^{m\times l}$
+3:for $t=1$ to mlkmlk do $\triangleright$ Exploration Phase
+4:  Explore each $(i,j)$ in $S_{A}$ $k$ times
+5:  Update $\hat{A}$ using (4)
+6:end for
+7:for $t=mlk+1$ to $T$ do $\triangleright$ Commit Phase
+8:  Play an equilibrium $(i^{*},j^{*})$ satisfying the following property:
 $\hat{A}(i,j^{*})\leq\hat{A}(i^{*},j^{*})\leq\hat{A}(i^{*},j),\quad\forall i\in S_{x},j\in S_{y}$
-Algorithm 1 ETC-TPZSG
+9:end for
 Theorem 1.
-The Nash regret of Algorithm 1, when interacting with $\sigma$-subgaussian payoffs, is upper bounded as follows:
+The Nash regret of Algorithm 1, when interacting with $\sigma$-subgaussian payoffs and $1\leq Nk\leq T$, is upper bounded as follows:
 $R_{T}^{*}\leq k\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}+(T-Nk)\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}\exp{\Big(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big)}$ (14)
@@
 where the first term $T\Delta$ is the worst-case regret and for the other, we combine the regret we obtain in (16) with the value of the exploration time in (17).
-Then, focusing on $\max\Big\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\}$, we find that $\Delta=\frac{4\sigma}{\sqrt{T}}$ is a critical point.
-If we put it in the regret bound (18), we have the following:
+Then, focusing on $\max\Big\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\}$, we obtain $\Delta\geq\frac{4\sigma}{\sqrt{T}}$.
+Thus, when we put it in the regret bound (18), we have the following:
 $R_{T}^{*}\leq\Delta+c\sqrt{T}$ (19)
@@
 We use the concept of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), which approximately satisfies the standard NE condition, to eliminate action pairs.
-Specifically, if a pure NE exists at $(i^{*},j^{*})$, then it must satisfy the condition given in (3) while a $\varepsilon$-NE satisfies the NE conditions within a tolerance level $\varepsilon$, making it a suitable criterion in learning settings where exact equilibrium identification is challenging.
+Specifically, if a pure NE exists at $(i^{*},j^{*})$, then it must satisfy the condition given in (3) while $\varepsilon$-NE satisfies the NE conditions within a tolerance level $\varepsilon$, making it a suitable criterion in learning settings where exact equilibrium identification is challenging.
 Based on this criterion, we introduce an elimination method in Algorithm 2, inspired by [6], to reduce unnecessary exploration of clearly suboptimal action pairs.
@@
 To address this, we initialize $\hat{\Delta}_{t}$ with $4\sigma$, which provides a practical starting point for the analysis.
-We set $\hat{\Delta}_{t}=2^{-t+2}\sigma$ in each round $t$, where $t$ ranges from 0 to $\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$.
+We set $\hat{\Delta}_{t}=2^{-t+2}\sigma$ in each round $t$, where $t$ ranges from 0 to $\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$, which allows a nonnegative exploration time $k_{t}$ and tolerance level $\varepsilon_{t}$.
 We have $\hat{\Delta}_{t}\approx\sqrt{(16\sigma^{2}/k_{t})\ln{(\hat{\Delta}_{t}^{2}T/16\sigma^{2})}}$ from (21), which implies $\ln{(\hat{\Delta}_{t}^{2}T/16\sigma^{2})}\geq 0$.
@@
 The division by $e$ is a technical adjustment, which makes the bound safer in the analysis; otherwise, the number of plays would be zero in the last round.
+For the property of $\varepsilon$-NE, we define the tolerance parameter as follows:
+$\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}$ (22)
+where $\varepsilon_{t}$ reflects how much players are allowed to deviate from the equilibrium and the tolerance over time gradually decreases to ensure convergence towards the NE.
+Substituting condition $k_{t}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}$ into the equation, where $k_{t}$ is defined in (21), we obtain $\varepsilon_{t}\leq\hat{\Delta}_{t}/2$.
+This bound indicates that as the number of rounds $t$ increases, the tolerance $\varepsilon_{t}$ becomes smaller, forcing the system to narrow the set of action pairs by progressively eliminating those that do not approximate a NE.
+As a result, over time, the algorithm keeps fewer action pairs, effectively focusing the exploration on action pairs that are close to the optimal one.
 Since we assume that a unique pure NE exists in the game, a well estimated game matrix $\hat{A}$ should also exhibit the properties necessary to support the equilibrium.
@@
 The Nash regret of Algorithm 2, interacting with $\sigma$-subgaussian payoffs, is upper bounded by
-$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}+\Delta_{ij}^{*}T\Big)$ (22)
+$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\Big)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (23)
 where $\lambda\geq 4\sigma\sqrt{e/T}$, and $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$ are two subsets of the action pairs.
@@
 ∎
-1:Input:
+Algorithm 2 ETC-TPZSG-AE
+1:Input: $S_{A},m,l,\sigma,T(1\leq ml\ll T)$
+2:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$
+3:for $t=0,1,2,...,\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$ do
+4:  if $\mathbf{|S_{t}|>1}:$ then $\triangleright$ Exploration Phase
+5:   Explore each $(i,j)$ in $S_{t}$ $k_{t}$ times, defined in (21)
+6:   Update $\hat{A}$ using (4)
 […]
-3:$m$: number of actions for row player
-4:$l$: number of actions for column player
-5:$T$: time horizon $\triangleright$ $1\leq ml\ll T$
-[…]
-7:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$
-8:In round $t=0,1,2,...,\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$;
-9:    Action Pair Selection:
-10:     If $\mathbf{|S_{t}|>1}:$ $\triangleright$ Exploration Phase
-11:      Explore each $(i,j)$ in $S_{t}$ $k_{t}$ times, defined in (21)
-12:      Update $\hat{A}$ using (4)
-13:     Else: $\triangleright$ Committing Phase
-14:      Play with a unique action pair in $S_{t}$ until step $T$
-15:    Action Pair Elimination:
-16:     Remove the action pairs $(i,j)$ NOT satisfying the following for $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}}$ :
+8:   Play with a unique action pair in $S_{t}$ until step $T$
+9:  end if
+10:  Remove the action pairs $(i,j)$ from $S_{t}$ NOT satisfying the following for $\varepsilon_{t}$ in (22): $\triangleright$ Elimination Phase
 $\hskip 35.56593pt\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
-17:    Update:
-18:     $S_{t+1}$ by "Action Pair Elimination" step
-19:     $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
-Algorithm 2 ETC-TPZSG-AE
+11:  $S_{t+1}\leftarrow S_{t}$
+12:  $\hat{\Delta}_{t+1}\leftarrow\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
+13:end for
 Moreover, if we consider another regret approach in (12), we obtain the following regret bound.
@@
 The external regret incurred by Algorithm 2 when interacting with $\sigma$-subgaussian payoffs is bounded as following:
-$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Big(\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}+\lambda T\Big).$ (23)
+$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Big(\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}\Big)+\lambda T.$ (24)
 Proof.
@@
 Our results are consistent with the regret bound derived by [6].
-Compared to our regret bound $O(\frac{\log(T\Delta^{2})}{\Delta})$ with the bound $O(\frac{\log_{+}(T\Delta^{2})}{\Delta})$ in [18], we observe that our study introduces a more refined notion of the suboptimality gap $\Delta$.
+Compared to our regret bound $O(\frac{\log(T\Delta^{2})}{\Delta})$ with the bound $O(\frac{\log_{+}(T\Delta^{2})}{\Delta})$ in [18], although they seem very similar, we observe that our study introduces a more refined notion of the suboptimality gap $\Delta$.
 Specifically, in [18], $\Delta$ refers to the gap between pure NE and the action chosen when the opponent plays the action in equilibrium.
@@
 In this section, we extend the adaptive elimination algorithm by introducing a variant, ETC-TPZSG-AE-NUE.
-It uses a non-uniform exploration method The non-uniform exploration in these algorithms is based on players’ strategies, which indirectly reflect the probability of being NE.
+It uses a non-uniform exploration method, which is based on players’ strategies, which indirectly reflect the probability of being NE.
 The key difference from the previous algorithm is that Algorithm 3 explores action pairs in proportion to the strategies of the players derived from the estimated payoff matrix.
@@
 Since the true game matrix is assumed to have the pure NE, we expect these estimated strategies to converge to the pure strategies over time.
-The estimated strategies of the players serve as a basis to determine the exploration time in our proposed algorithms.
-Furthermore, the probability of selecting an action pair $(i,j)$ is estimated by $\hat{P}_{ij,t}=\hat{p}_{i,t}\hat{q}_{j,t}$, which satisfies $\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=1$.
+The estimated strategies of the players serve as a basis to determine the exploration time in Algorithm 3.
+The probability of selecting an action pair $(i,j)$ is estimated by $\hat{P}_{ij,t}=\hat{p}^{*}_{i,t}\hat{q}^{*}_{j,t}$, which satisfies $\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=1$, where $(\hat{p}_{t}^{*},\hat{q}_{t}^{*})$ is a mixed NE in round $t$.
+In other words, once optimal strategies are determined, we calculate the joint probability $\hat{P}_{ij,t}$, which represents the probability that the row player chooses action $i$ and the column player chooses action $j$ in round $t$.
 Consequently, in Algorithm 3, the exploration time for each action pair $(i,j)$ is proportional to $\hat{P}_{ij,t}$, allowing action pairs with higher estimated probabilities to be played more frequently.
-We incorporate the strategies of the players $\hat{p}_{t}$ and $\hat{q}_{t}$ together with $k_{t}$.
+We incorporate the strategies of the players $\hat{p}^{*}_{t}$ and $\hat{q}^{*}_{t}$ together with $k_{t}$.
 Specifically, in each round $t$, this involves estimating the optimal strategies for each player by solving a linear system using the estimated payoff matrix $\hat{A}$ as follows:
-$\begin{aligned} &\\ &\hat{P}_{t}=\hat{p}_{t}\hat{q}_{t}\\ &\hat{p}_{t}=\max_{p}\min_{q}p^{T}\hat{A}q\\ &\hat{q}_{t}=\min_{q}\max_{p}p^{T}\hat{A}q\\ \end{aligned}\qquad\begin{aligned} &\text{Subject to:}\\ &\sum_{i\in S_{x}}\hat{p}_{i,t}=1,\quad\hat{p}_{i,t}\geq 0,\ \forall i\in S_{x}\\ &\sum_{j\in S_{y}}\hat{q}_{j,t}=1,\quad\hat{q}_{j,t}\geq 0,\ \forall j\in S_{y}\\ &\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=\sum_{i,j}\hat{p}_{i,t}\hat{q}_{j,t}=1\end{aligned}$ (24)
+$\begin{aligned} &\\ &\hat{P}_{ij,t}=\hat{p}^{*}_{i,t}\hat{q}^{*}_{j,t}\\ &\hat{p}^{*}_{t}=\operatorname*{argmax}_{\hat{p}}\min_{\hat{q}}\hat{p}^{T}\hat{A}\hat{q}\\ &\hat{q}^{*}_{t}=\operatorname*{argmin}_{\hat{q}}\max_{\hat{p}}\hat{p}^{T}\hat{A}\hat{q}\\ \end{aligned}\qquad\begin{aligned} &\text{Subject to:}\\ &\sum_{i\in S_{x}}\hat{p}_{i,t}=1,\quad\hat{p}_{i,t}\geq 0,\ \forall i\in S_{x}\\ &\sum_{j\in S_{y}}\hat{q}_{j,t}=1,\quad\hat{q}_{j,t}\geq 0,\ \forall j\in S_{y}\\ &\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=\sum_{i,j}\hat{p}_{i,t}\hat{q}_{j,t}=1\end{aligned}$ (25)
 Thus, the exploration time is given by the following:
@@
 As a result, due to this extended exploration phase, the number of rounds is reduced by half.
-1:Input:
+Algorithm 3 ETC-TPZSG-AE-NUE
+1:Input: $S_{A},m,l,\sigma,T(1\leq ml\ll T)$,
+2:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$
+3:for $t=0,1,2,...,\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$ do
+4:  if $\mathbf{|S_{t}|>1}:$ then $\triangleright$ Exploration Phase
+5:   Explore each $(i,j)$ in $S_{t}$ $k_{ij,t}$ times, defined in (26)
+6:   Update $\hat{A}$ using (4)
 […]
-3:$m$: number of actions for row player
-4:$l$: number of actions for column player
-5:$T$: time horizon $\triangleright$ $1\leq ml\ll T$
-[…]
-7:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$, $\hat{A}=[0]^{m\times l}$ and $P_{ij,0}=\frac{1}{ml},\forall i,j$
-8:In round $t=0,1,2,...,\left\lfloor\frac{1}{4}\log_{2}\frac{T}{e}\right\rfloor$;
-9:    Action Pair Selection:
-10:     If $\mathbf{|S_{t}|>1}:$ $\triangleright$ Exploration Phase
-11:      Explore each $(i,j)$ in $S_{t}$ $k_{ij,t}$ times, defined in (25)
-12:      Update $\hat{A}$ using (4).
-13:     Else: $\triangleright$ Committing Phase
-14:      Play with a unique action pair in $S_{t}$ until step $T$
-15:    Action Pair Elimination:
-16:     Remove the action pairs $(i,j)$ NOT satisfying the following for $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}}$:
+8:   Play with a unique action pair in $S_{t}$ until step $T$
+9:  end if
+10:  Remove the action pairs $(i,j)$ from $S_{t}$ NOT satisfying the following for $\varepsilon_{t}$ in (22): $\triangleright$ Elimination Phase
 $\hskip 35.56593pt\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
-17:    Update:
-18:     $\hat{P}_{ij,t}$ by solving the linear system in (24).
-19:     $S_{t+1}$ by "Action Pair Elimination" step
-20:     $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
-Algorithm 3 ETC-TPZSG-AE-NUE
+11:  Solve the linear system in (25) to get $\hat{P}_{ij,t+1}$
+12:  $S_{t+1}\leftarrow S_{t}$
+13:  $\hat{\Delta}_{t+1}\leftarrow\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
+14:end for
 Theorem 4.
 The Nash regret of Algorithm 3 is upper bounded by
-$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}+\Delta_{ij}^{*}T\Big)$ (26)
+$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\Big)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (27)
 where $\lambda\geq 4\sigma\sqrt[4]{e/T}$ and, $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$ are two subsets of the action pairs.
@@
 The results align with the bound in Theorem 1.
-In Figures 3 and 3, we compare the proposed algorithms and Tsallis-INF [18], the only existing method with instance-dependent bounds for TPZSGs, based on cumulative regret, computed as the absolute difference between the game value and the payoff from the action played to avoid negative values, averaged over $10^{2}$ runs.
+In Figures 3 and 3, we compare the proposed algorithms and Tsallis-INF [18], very recent method with instance-dependent bounds for TPZSGs, based on cumulative regret, computed as the absolute difference between the game value and the payoff from the action played to avoid negative values, averaged over $10^{2}$ runs.
 We use a $2\times 2$ game matrix by Gaussian payoffs with variance $\sigma^{2}$ and $k$ exploration rounds per action pair for ETC-TPZSG is randomly selected from a predefined list.
@@
 The authors would like to thank Debabrota Basu, Thomas Kleine Buening and Ronald Ortner for fruitful discussions on this topic.
-Appendix A Subgaussian Properties
-This section collects some supplementary results from [20] that are used repeatedly throughout our proofs.
-These results concern standard properties of subgaussian random variables such as tail inequalities.
-They are included here for completeness and easy reference as they play a critical role in our analyses.
-The statements are not original, we include only the parts that are directly useful for our purposes.
-Theorem 5 ([20, Theorem 5.3]).
+Appendix A Useful Properties
+In this section, we present several properties that are used throughout the proofs.
+Moreover, this section collects some supplementary results from [20].
+These results concern the standard properties of subgaussian random variables.
+They are included here for completeness and easy reference, as they play a critical role in our regret analyses.
+Lemma 1.
+For any $a,b>0$, we can write the following:
+$\max\{a^{2},b^{2}\}\geq\frac{a^{2}+b^{2}}{2}$ (28)
+Proof.
+If we assume that their maximum is $a^{2}$, then we have $a^{2}\geq a^{2}+b^{2}/2$, which implies $a^{2}\geq b^{2}$.
+Similarly, if $\max\{a^{2},b^{2}\}=b^{2}$, it gives us $b^{2}\geq a^{2}$.
+They are true by assumptions, and thus we conclude the proof.
+∎
+Lemma 2.
+For any $a,b>0$, the following holds:
+$a^{2}+b^{2}\geq\frac{(a+b)^{2}}{2}$ (29)
+Proof.
+The proof follows from the fact that $(a-b)^{2}\geq 0$. ∎
+Lemma 3.
+For events $e_{1},e_{2},\cdots,e_{n}$, we have the following for any $i\in\{1,2,\cdots,n\}$:
+$\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{n})\leq\mathbb{P}(e_{i}).$ (30)
+Proof.
+For any $i\in\{1,2,\cdots,n\}$, we observe that $e_{1}\cap e_{2}\cap\cdots\cap e_{n}\subseteq e_{i}$.
+Thus, the proof comes from the fact that an intersection of events cannot be more likely than any of the events involved.
+∎
+Lemma 4 ([20, Theorem 5.3]).
 If $X$ is $\sigma$-subgaussian random variable, for any $\epsilon\geq 0$,
@@
 Appendix B Proof of Theorem 1
-To analyze the Nash regret of the Algorithm 1, we begin by decomposing the total expected regret into two phases: the exploration phase and the committing phase.
-During exploration, each action pair is played a fixed number of times $k$ to estimate their mean rewards, potentially incurring regret if suboptimal action pairs are played.
+To analyze the Nash regret of the Algorithm 1, we begin by decomposing the total expected regret into two phases: the exploration phase and the commit phase.
+During exploration, each action pair is played a fixed number of times $k$ to estimate their mean payoffs, potentially incurring regret if suboptimal action pairs are played.
 Once the algorithm commits to the empirically best action pair, which is the pure NE, the regret accumulates only if this action pair is not the true optimal one.
-Our goal is to bound the expected regret by quantifying the probability of committing to a suboptimal action pair, which we can call it as a misidentified NE, based on the estimates of their payoffs obtained during the exploration phase.
+Our goal is to bound the expected regret by quantifying the probability of committing to a suboptimal action pair, which we can call it a misidentified NE, based on the estimates of their payoffs obtained during the exploration phase.
 Let $(i^{\prime},j^{\prime})$ denote a misidentified NE, so we aim to identify an action pair $(i^{\prime},j^{\prime})$, which appears better than $(i^{*},j^{*})$ based on the estimated matrix.
@@
 $\mathbb{P}(E_{1}\cap E_{2})$ $=\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j),\forall i,j)$ (32)
-$\leq\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))$ (29)
-$=\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime}))\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))}$ (30)
-$=\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)}$ (31)
-$\leq\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})\leq\Delta_{i^{\prime}j^{\prime}}^{\text{max}})}$
-$\quad\times\sqrt{\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)\leq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})}$ (32)
+$\leq\mathbb{P}(\hat{A}(i_{1},j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j_{1}))$ (33)
+$\leq\sqrt{\mathbb{P}(\hat{A}(i_{1},j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime}))\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j_{1}))}$ (34)
+$=\sqrt{\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i_{1},j^{\prime})\geq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j_{1})-\hat{A}(i^{\prime},j^{\prime})\geq 0)}$ (35)
+$\leq\sqrt{\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})-(\hat{A}(i_{1},j^{\prime})-A(i_{1},j^{\prime}))\geq\Delta_{i^{\prime}j^{\prime}}^{\text{max}})}$
+$\quad\times\sqrt{\mathbb{P}(\hat{A}(i^{\prime},j_{1})-A(i^{\prime},j_{1})-(\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime}))\geq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})}$ (36)
 $\leq\sqrt{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\Big)}\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\Big)}}$ (37)
@@
 $=\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}})^{2}}{16\sigma^{2}}\Big)}$ (41)
-where $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}=\max_{i}A(i,j^{\prime})-A(i^{\prime},j^{\prime})$, $\Delta_{i^{\prime}j^{\prime}}^{\text{min}}=A(i^{\prime},j^{\prime})-\min_{j}A(i^{\prime},j)$ and $\Delta_{i^{\prime}j^{\prime}}=\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}}$.
-To get (30), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, that is, for any collection of events $e_{1},e_{2},\ldots,e_{n}$, we have $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{n})\leq\mathbb{P}(e_{i})$ for all $i\in{1,\ldots,n}$.
-The inequality (32) holds by using the followings:
-$\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)=\mathbb{P}(\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})\leq A(i,j^{\prime})-A(i^{\prime},j^{\prime}))$ (38)
-$\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)=\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)\leq A(i^{\prime},j^{\prime})-A(i^{\prime},j))$ (39)
-Then, we apply $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}\geq A(i,j^{\prime})-A(i^{\prime},j^{\prime}),\Delta_{i^{\prime}j^{\prime}}^{\text{min}}\geq A(i^{\prime},j^{\prime})-A(i^{\prime},j)$, which are true from their definitions.
-For step (33), we apply Theorem 5 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable.
-At this point, we need to show that the differences $\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})$ and $\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)$ are $\sqrt{2\sigma^{2}/k}$-subgaussian but it is enough to check that for any action pairs $(i,j)$ and $(x,y)$ in $S_{A}$, $\hat{A}(i,j)-\hat{A}(x,y)$ are $\sqrt{2\sigma^{2}/k}$-subgaussian.
+where we assume that $A(i_{1},j^{\prime})=\max_{i\in S_{x}}A(i,j^{\prime})$ and $A(i^{\prime},j_{1})=\min_{i\in S_{x}}A(i^{\prime},j)$.
+On the other hand, we note $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}=\max_{i}A(i,j^{\prime})-A(i^{\prime},j^{\prime})$, $\Delta_{i^{\prime}j^{\prime}}^{\text{min}}=A(i^{\prime},j^{\prime})-\min_{j}A(i^{\prime},j)$ and $\Delta_{i^{\prime}j^{\prime}}=\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}}$.
+To obtain (33), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, as stated in 3.
+For step (37), we apply Lemma 4 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable.
+At this point, we need to show that the differences $\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})-\hat{A}(i_{1},j^{\prime})+A(i_{1},j^{\prime})$ and $\hat{A}(i^{\prime},j_{1})-A(i^{\prime},j_{1})-\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})$ are $\sqrt{2\sigma^{2}/k}$-subgaussian.
 To establish this, we utilize Lemma 5, which provides key properties of subgaussian random variables.
-Specifically, both $\hat{A}(i,j)$ and $\hat{A}(x,y)$ are $\sqrt{\sigma^{2}/k}$-subgaussian as the average payoffs are computed by the equation (4) and during the exploration phase, each action pair is played $k$ times.
-By the lemma, the difference of two independent subgaussian random variables with parameters $\sqrt{\sigma^{2}/k}$ is subgaussian with parameter $\sqrt{2\sigma^{2}/k}$, which follows that $\hat{A}(i,j)-\hat{A}(x,y)$ is $\sqrt{2\sigma^{2}/k}$-subgaussian for any action pairs $(i,j)$ and $(x,y)$.
-To get the last inequality (35), we observe that
-$\frac{\Delta_{i^{\prime}j^{\prime}}^{2}}{2}=\frac{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{2}\leq(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}$
-where the inequality follows from the fact that $(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}-\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\geq 0$.
-As we mention before, there is an alternative approach to bound $\mathbb{P}(E_{1}\cap E_{2})$ to find a bound for the probability of a misidentified NE.
+Specifically, for any $(i,j)\in S_{A}$, $\hat{A}(i,j)-A(i,j)$ are $\sqrt{\sigma^{2}/k}$-subgaussian since during the exploration phase, each action pair is played $k$ times.
+Then by Lemma 5, the difference of two independent subgaussian random variables with parameters $\sqrt{\sigma^{2}/k}$ is subgaussian with parameter $\sqrt{2\sigma^{2}/k}$.
+To obtain the last inequality (39), we apply Lemma 2.
+As we mentioned earlier, there is an alternative approach to bound $\mathbb{P}(E_{1}\cap E_{2})$ to find a bound for the probability of a misidentified NE.
 It involves an inequality based on the minimum of individual probabilities.
@@
 $\mathbb{P}(E_{1}\cap E_{2})$ $=\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j),\forall i,j)$ (42)
-$\leq\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))$ (41)
-$=\min\{\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})),\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))\}$ (42)
-$=\min\{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)\}$ (43)
+$\leq\mathbb{P}(\hat{A}(i_{1},j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j_{1}))$ (43)
+$\leq\min\{\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i_{1},j^{\prime})\geq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j_{1})-\hat{A}(i^{\prime},j^{\prime})\geq 0)\}$ (44)
 $\leq\min\Big\{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\Big)},\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\Big)}\Big\}$ (45)
@@
 $\leq\exp{\Big(-\frac{k\Delta_{i^{\prime}j^{\prime}}^{2}}{16\sigma^{2}}\Big)}$ (49)
-where to get the inequality (46), we use the following:
-$\max\{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2},(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\}\geq\frac{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{2}$ (49)
-If we assume that their maximum is $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}$, then it will imply $(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}\geq(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}$.
-Similarly, if $\max\{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2},(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\}=(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}$, it will give us $(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\geq(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}$.
-They are true by assumptions, then, (49) holds.
+where to get the inequality (47), we use Lemma 1.
 Thus, both approaches provide the same upper bound for the probability of misidentified NE.
 In the regret analysis, we will focus on the loss due to playing with suboptimal (or not true NE) action pairs.
-Thus, after $T$ times playing, $1\leq Nk\leq T$ where $N=ml$ as the number of action pairs, we can write the Nash regret as
-$R_{T}^{*}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]$
-where $\mathbb{E}[n_{ij,T}]$ is the expected number of playing the action pair $(i,j)$.
+Thus, after playing $T$ times, $1\leq Nk\leq T$ where $N=ml$ is the number of action pairs, we can write the Nash regret as
+$R_{T}^{*}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]$ (50)
+where $\mathbb{E}[n_{ij,T}]$ is the expected number of plays for the action pair $(i,j)$.
 We already know that each action pair $(i,j)\in S_{A}$ is played at least $k$ times for exploration.
-After the exploration step, we will consider playing the misidentified NE $(i^{\prime},j^{\prime})$ instead of true one $(i^{*},j^{*})$ because we expect the players to select NE action pair as optimal after the exploration phase.
+After the exploration step, we will consider playing the misidentified NE $(i^{\prime},j^{\prime})$ instead of the true one $(i^{*},j^{*})$ because we expect the players to select the NE action pair as optimal after the exploration phase.
 Since we already have the probability of playing a misidentified NE from (41) or (49), we can write
-$\mathbb{E}[n_{ij,T}]\leq k+(T-Nk)\exp{\Big(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big)}.$
+$\mathbb{E}[n_{ij,T}]\leq k+(T-Nk)\exp{\Big(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big)}.$ (51)
 Hence, the proof is concluded. $\Box$
 Appendix C Proof of Theorem 2
-According to the algorithm, $4\sigma\sqrt{e/T}$ is a critical value for $\hat{\Delta}_{t}$ since it takes this value in round $\frac{1}{2}\log_{2}\frac{T}{e}$.
+According to the algorithm, $4\sigma\sqrt{e/T}$ is a critical value for $\hat{\Delta}_{t}$ since it takes this value in the round $\frac{1}{2}\log_{2}\frac{T}{e}$.
 In this way, let define a threshold parameter as $\lambda\geq 4\sigma\sqrt{e/T}$, which gives us two subsets of actions pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$.
-Since $\Delta_{ij}^{*}\leq\Delta_{ij}$ ensures the bounds are preserved, it makes sense to use these action pair sets in the analysis.
+Since $\Delta_{ij}^{*}\leq\Delta_{ij}$, it ensures that the bounds are preserved and it makes sense to use these action pair sets in the analysis.
 Thus, we can write the Nash regret as
-$R_{T}^{*}=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]+\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]$
-where $\Delta_{ij}^{*}=A(i^{*},j^{*})-A(i,j).$ and $n_{ij,T}$ is the total number of times to play action pair $(i,j)$ by time $T$.
+$R_{T}^{*}=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]+\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]$ (52)
+where $\Delta_{ij}^{*}=A(i^{*},j^{*})-A(i,j).$ and $n_{ij,T}$ is the total number of times to play the action pair $(i,j)$ by time $T$.
 To analyze the Nash regret of Algorithm 2, it is essential to consider various scenarios that may lead to suboptimal outcomes.
 Each of these scenarios contributes to the total Nash regret, as they either involve the incorrect elimination of the optimal action pair or the unnecessary selection of suboptimal pairs.
-The former results in discarding the optimal strategy because of insufficient evidence, while the latter leads to spend more time with action pairs that are unlikely to be part of any near optimal equilibrium.
+The former results in discarding the optimal strategy because of insufficient evidence, while the latter leads to spending more time with action pairs that are unlikely to be part of any near optimal equilibrium.
 The algorithm employs a scheduling mechanism for $\Delta_{ij}$, where $\hat{\Delta}_{t}$ is halved in each round.
@@
 As $\hat{\Delta}_{t}$ decreases, the algorithm allocates more exploration to action pairs that are closer to being part of a NE, because those that are clearly suboptimal have already been eliminated.
-This adaptive process enables a more precise identification of near-optimal strategies by $\varepsilon$-NE condition while minimizing regret.
+This adaptive process enables a more precise identification of near-optimal strategies by the $\varepsilon$-NE condition while minimizing regret.
 For each suboptimal action pair $(i,j)$, let $t_{ij}=\min\{t:\hat{\Delta}_{t}<\Delta_{ij}/2\}$ refer to the earliest round such that $\hat{\Delta}_{t}<\Delta_{ij}/2$.
+Specifically, we are interested in the first round where $\hat{\Delta}_{t}$ satisfies both $\hat{\Delta}_{t}<\Delta_{ij}^{\text{max}}$ and $\hat{\Delta}_{t}<\Delta_{ij}^{\text{min}}$, but combining these conditions results in the simplified condition $\hat{\Delta}_{t}<\Delta_{ij}/2$.
 Using the fact that $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ and the definition of $t_{ij}$, we can write the following:
 $\frac{1}{\hat{\Delta}_{t_{ij}}}\leq\frac{4}{\Delta_{ij}}<\frac{1}{\hat{\Delta}_{t_{ij}+1}}$ (53)
+Furthermore, we have the following:
+$\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\Big(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\Big)}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{{t_{ij}}+1}<\frac{\Delta_{ij}}{4}$ (54)
+where the exploration time $k_{t}$ is defined in (21), which satisfies $k_{t_{ij}}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}$.
 For simplicity of notation, we write $\hat{A}$ instead of $\hat{A}_{t}$ for the estimated payoff matrix in round $t$ throughout this section.
@@
 A suboptimal action pair $(i,j)$ is not eliminated in round $t_{ij}$ or earlier while the optimal action pair is in the set $S_{t_{ij}}$.
-Since neither some suboptimal nor the optimal action pair is not eliminated, the algorithm fails to discard suboptimal choices.
+Since neither some suboptimal nor the optimal action pair is eliminated, the algorithm fails to discard suboptimal choices.
 It might converge to an incorrect solution without eliminating suboptimal actions, but since it still maintains the optimal action pair within the action pair selection set.
@@
 Let explain briefly in which situations the algorithm eliminates or keeps an action pair.
-For action pair $(i,j)$, if $\varepsilon$-NE property does not hold in round $t=t_{ij}$ which implies that at least one of the inequalities fails, then it will be eliminated in round $t_{ij}$.
-Furthermore, we have $\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\Big(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\Big)}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{t_{ij}+1}<\frac{\Delta_{ij}}{4}$.
+For action pair $(i,j)$, if the $\varepsilon$-NE property does not hold in round $t=t_{ij}$, which implies that at least one of the inequalities fails, then it will be eliminated in round $t_{ij}$.
 In other words, if there exists $i^{\prime}\in S_{x}$ or $j^{\prime}\in S_{y}$, such that one of the following is true:
-$I_{1}:\hat{A}(i^{\prime},j)-\hat{A}(i,j)>\varepsilon_{t_{ij}},\qquad I_{2}:\hat{A}(i,j)-\hat{A}(i,j^{\prime})>\varepsilon_{t_{ij}},$
+$I_{1}:\hat{A}(i^{\prime},j)-\hat{A}(i,j)>\varepsilon_{t},\qquad I_{2}:\hat{A}(i,j)-\hat{A}(i,j^{\prime})>\varepsilon_{t},$ (55)
 then $(i,j)$ is eliminated.
-The inequality $I_{1}$ indicates that there exists an alternative action $i^{\prime}$ for the row player that offers a higher payoff than action $i$ against the column action $j$.
-Similarly, $I_{2}$ implies that there exists a better action $j^{\prime}$ for the column player than action $j$ when playing against the row action $i$.
-If either $I_{1}$ or $I_{2}$ holds, the action pair $(i,j)$ cannot be a part of the NE, as at least one player has an incentive to deviate.
+The inequality $I_{1}$ indicates that there exists an alternative action $i^{\prime}$ for the row player that offers a higher payoff than action $i$ against column action $j$.
+Similarly, $I_{2}$ implies that there exists a better action $j^{\prime}$ for the column player than the action $j$ when playing against the row action $i$.
+If $I_{1}$ or $I_{2}$ holds, the action pair $(i,j)$ cannot be part of the NE, as at least one player has an incentive to deviate.
 Hence, $(i,j)$ is not a NE with high probability.
@@
 there exists at least one $i^{\prime}$ such that $A(i^{\prime},j)>A(i,j)$, the probability of $E_{1}$ can be expressed as follows:
-$\mathbb{P}(E_{1})$ $=\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t},\forall i^{\prime}\in S_{x})$ (52)
-$\leq\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t})$ (53)
-$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{\frac{4\sigma^{2}}{k_{t}}}\Bigg)}$ (54)
-$=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}$ (55)
-where $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}.$ To derive inequality (53), we use the fact that the probability of the intersection of multiple events is always less than or equal to the probability of any of individual events.
-Specifically, since we have $m$ actions in $S_{x}$, for any collection of events $e_{1},e_{2},\cdots,e_{m}$, it holds that $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{m})\leq\mathbb{P}(e_{i})$ for all $i=1,2,\cdots,m$.
-In the last inequality (54), we apply Theorem 5 with $\sqrt{2\sigma^{2}/k_{t}}$-subgaussian random variable.
-The subgaussian parameter is derived by using Lemma 1 and we note that each action pair is played $k_{t}$ times.
+$\mathbb{P}(E_{1})$ $=\mathbb{P}(\hat{A}(i,j)-\hat{A}(i^{\prime},j)+\varepsilon_{t}\geq 0,\forall i^{\prime}\in S_{x})$ (57)
+$\leq\mathbb{P}(\hat{A}(i,j)-\hat{A}(i_{1},j)+\varepsilon_{t}\geq 0)$ (58)
+$=\mathbb{P}(\hat{A}(i,j)-A(i,j)-(\hat{A}(i_{1},j)-A(i_{1},j))\geq\Delta_{ij}^{\text{max}}-\varepsilon_{t})$ (59)
+$\leq\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}$ (60)
+where we assume $A(i_{1},j)=\max_{i\in S_{x}}A(i,j)$ and $A(i,j_{1})=\min_{j\in S_{y}}A(i,j)$.
+To derive the inequality (58), we use the fact that the probability of the intersection of multiple events is always less than or equal to the probability of any of the individual events.
+Specifically, since we have $m$ actions in $S_{x}$, for any collection of events $e_{1},e_{2},\cdots,e_{m}$, we have $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{m})\leq\mathbb{P}(e_{1})$, as in Lemma 3.
+Then, we see that $\hat{A}(i,j)-A(i,j)-(\hat{A}(i_{1},j)-A(i_{1},j)$ is $\sqrt{2\sigma^{2}/k_{t}}$-subgaussian using subgaussian properties.
+Thus, we apply Lemma 4 to get inequality (60) since $\Delta_{ij}^{\text{max}}-\varepsilon_{t}\geq 0$ where $t$ is round $t_{ij}$ or later.
 Similarly, in order to calculate the probability of the event $E_{2}$, we can write
-$\mathbb{P}(E_{2})=\mathbb{P}(\hat{A}(i,j)-\hat{A}(i,j^{\prime})\leq\varepsilon_{t},\forall j^{\prime}\in S_{y})\leq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (56)
-Therefore, we have the probability of keeping an action pair $(i,j)$ in any round $t$ as
-$\mathbb{P}(E_{1}\cap E_{2})\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (57)
-It means that the probability that a suboptimal action pair $(i,j)$ is not eliminated in round $t_{ij}$ or before is bounded by $\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}T}$.
+$\mathbb{P}(E_{2})$ $=\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i,j)+\varepsilon_{t}\geq 0,\forall i^{\prime}\in S_{x})$ (61)
+$\leq\mathbb{P}(\hat{A}(i,j_{1})-\hat{A}(i,j)+\varepsilon_{t}\geq 0)$ (62)
+$=\mathbb{P}(\hat{A}(i,j_{1})-A(i,j_{1})-(\hat{A}(i,j)-A(i,j))\geq\Delta_{ij}^{\text{min}}-\varepsilon_{t})$ (63)
+$\leq\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}$ (64)
+where $A(i,j_{1})=\min_{j\in S_{y}}A(i,j)$.
+We note that $\hat{A}(i,j_{1})-A(i,j_{1})-(\hat{A}(i,j)-A(i,j)$ is $\sqrt{2\sigma^{2}/k_{t}}$-subgaussian.
+Then, we apply Lemma 4, as we have $\Delta_{ij}^{\text{min}}-\varepsilon_{t}\geq 0$ for $t=t_{ij}$ or later rounds from the definition of $t_{ij}$.
+Therefore, we write the probability of keeping an action pair $(i,j)$ as
+$\mathbb{P}(E_{1}\cap E_{2})\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}$ $\leq\min\Bigg\{\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)},\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}\Bigg\}$ (65)
+$=\exp{\Big(-\frac{k_{t}\max\{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2},(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}\}}{4\sigma^{2}}\Big)}$ (66)
+$\leq\exp{\Big(-\frac{k_{t}((\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}+(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2})}{8\sigma^{2}}\Big)}$ (67)
+$\leq\exp{\Big(-\frac{k_{t}(\Delta_{ij}-2\varepsilon_{t})^{2}}{16\sigma^{2}}\Big)}$ (68)
+$<\exp{\Big(-\frac{k_{t}\varepsilon_{t}^{2}}{4\sigma^{2}}\Big)}$ (69)
+$=\exp{\Bigg(-\frac{k_{t}\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{4\sigma^{2}}\Bigg)}$ (70)
+$=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (71)
+For (67), we apply Lemma 1.
+Then , to write (68), we use Lemma 2.
+Using the definition of $t_{ij}$ and (54), $\varepsilon_{t}<\Delta_{ij}/4$ for rounds $t\geq t_{ij}$.
+Then, we derive the inequality (69).
+It is important to note that this probability bound holds only for rounds $t_{ij}$ or later.
+However, the probability of keeping a suboptimal action pair $(i,j)$ in round $t_{ij}$ also accounts for previous rounds, as it implies that the algorithm has not eliminated this pair in earlier rounds.
+Therefore, the probability that a suboptimal action pair $(i,j)$ is not eliminated in round $t_{ij}$ or before is bounded by $\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}T}$.
 Then, the regret contribution is simply bounded by a worst case which is $T\Delta_{ij}^{*}$ for any suboptimal action pair $(i,j)\in S_{A_{1}}$.
@@
 Thus, using (53), each action pair $(i,j)$ is played at most $k_{t_{ij}}$ times such that
-$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq\Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil.$
+$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq\Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil.$ (73)
 Then, the regret contribution is expressed by
@@
 The optimal action pair $(i^{*},j^{*})$ is eliminated by some suboptimal one $(i,j)$ in round $t_{*}$ such that $t_{ij}\geq t_{*}$.
-It leads to a misidentification of the NE, implying that an action pair $(i,j)$ should be better than the optimal one based on the current estimates. By $\varepsilon$-NE property, clearly we have
+It leads to a misidentification of the NE, implying that an action pair $(i,j)$ should be better than the optimal one based on the current estimates.
+By the $\varepsilon$-NE property, we clearly have the following:
 $\hat{A}(i^{*},j)-\varepsilon_{t_{*}}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{*})+\varepsilon_{t_{*}}.$
 However, it is not enough to keep a suboptimal action pair $(i,j)$.
-On the other hand, we note that if the optimal action pair $(i^{*},j^{*})$ is eliminated in round $t_{*}$, it must fail to satisfy $\varepsilon$-NE property as the following:
+If the algorithm keeps a suboptimal action pair $(i,j)$ up to round $t_{*}$, the events specified in (56) must be satisfied with the parameter $\varepsilon_{t_{*}}$.
+Furthermore, we note that if the optimal action pair $(i^{*},j^{*})$ is eliminated in round $t_{*}$, it must fail to satisfy the $\varepsilon$-NE property as the following:
 $\hat{A}(i,j^{*})-\hat{A}(i^{*},j^{*})>\varepsilon_{t_{*}}\quad\text{or/and}\quad\hat{A}(i^{*},j^{*})-\hat{A}(i^{*},j)>\varepsilon_{t_{*}}.$
 That is, under the estimated payoff matrix $\hat{A}$, there exists an action $i$ that is estimated to yield a higher payoff than $i^{*}$ for the row player, and an action $j$ that is estimated to be more favorable than $j^{*}$ for the column player.
-The optimal action pair can only be eliminated by a suboptimal action pair $(i,j)$ in round $t_{*}$ such that $t_{ij}\geq t_{*}$ since action pair $(i,j)$ must remain under consideration at the elimination round of the optimal action pair.
+We observe that all action pairs $(i,j)$ with $t_{ij}<t_{*}$ have already been eliminated in round $t_{ij}$ or earlier, which we already consider this condition in the previous case.
+Then, the optimal action pair can only be eliminated by a suboptimal action pair $(i,j)$ in round $t_{*}$ such that $t_{ij}\geq t_{*}$ since action pair $(i,j)$ must remain under consideration at the elimination round of the optimal action pair.
 This condition also implies that $\hat{\Delta}_{t_{ij}}\leq\hat{\Delta}_{t_{*}}$ because of the decreasing behavior of $\hat{\Delta}_{t}$ across rounds in the algorithm.
 Moreover, it is important to note that the algorithm keeps this action pair $(i,j)$ for at least one additional round following the elimination of the optimal pair as it is identified as a NE.
-We further observe that all action pairs $(i,j)$ with $t_{ij}<t_{*}$ have already been eliminated in round $t_{ij}$ or earlier, which we already consider this condition in the previous case.
 On the other hand, for the action pairs in $S_{A_{2}}$, we assume that $\hat{\Delta}_{t}<\frac{\lambda}{2}$, which implies that a suboptimal action pair remains in the game until the final round.
@@
 Consequently, these suboptimal action pairs contribute to the overall regret and the analysis accounts for their effect to establish accurate performance guarantees.
-If the algorithm keeps a suboptimal action pair $(i,j)$ up to round $t_{*}$, the events specified in (51) must hold with parameter $\varepsilon_{t_{*}}$.
-Consequently, by applying (57), the probability of keeping a suboptimal action pair can be bounded by $\frac{16\sigma^{2}}{\hat{\Delta}_{t_{*}}^{2}T}$.
+We observe that the optimal action pair is eliminated by $(i,j)$ in some round $t_{*}$ where $t_{*}$ ranges from 0 to $t_{ij}$.
+Once this elimination occurs, the action pair action pair $(i,j)$ must remain active for at least one additional round, since it is selected as a misidentified NE.
+Hence, the algorithm necessarily keeps $(i,j)$ until at least round $t_{*}+1$.
+Furthermore, because $t_{*}\leq t_{ij}$, probability of keeping a suboptimal action pair $(i,j)$ in round $t_{*}+1$ can be bounded by $\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}+1}^{2}T}$.
 Thus, we can write the regret contribution of this case as
-$\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{16\sigma^{2}}{\hat{\Delta}_{t_{*}}^{2}T}$ $<\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}+1}^{2}T}$ (62)
-$=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{64\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}T}$ (63)
+$\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}+1}^{2}T}$ $=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{64\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}T}$ (77)
 $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\frac{512\sigma^{2}}{\Delta_{ij}^{2}}+\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}$ (78)
-where $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ and we apply (50).
-We use $\hat{\Delta}_{t+1}$ because the optimal action pair might be eliminated at some round $t_{*}$ such that $t_{*}\in[0,\max_{ij}t_{ij}]$.
-This implies that the optimal action pair can be eliminated by any $(i,j)$ no later than the last round that $(i,j)$ remains in the game.
-However, we must keep $(i,j)$ at least until the next round, since it is not removed in round $t_{*}$ and is selected as a misidentified NE.
+where $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ and we apply (53) to get (78).
+We consider the action pairs in $S_{A_{2}}$ for this case, as this suboptimal action pair is identified as a NE.
+This is necessary because we need to account for regret from the commit phase in Algorithm 2 and the round $t_{ij}+1$ might be the round when the commit phase started.
 Case 4.
 A suboptimal action pair $(i,j)$ from the set $S_{A_{2}}$ is selected as the unique action pair in $S_{t_{ij}}$ for remaining rounds.
-This implies that a suboptimal action pair is played during the committing phase of Algorithm 2 up to step $T$. Thus, we need to account for an additional regret contribution term given by
-$\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T.$ (65)
+This implies that a suboptimal action pair is played during the commit phase of Algorithm 2 up to step $T$.
+Thus, we need to account for an additional regret contribution term given by a trivial upper bound $\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$.
 This regret contribution accounts for scenarios where the algorithm selects suboptimal action pair as the pure NE.
@@
 Then, if we keep the same events from the proof of Theorem 2, Appendix C, we have the following:
-$\mathbb{P}(E_{1})$ $=\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t},\forall i^{\prime}\in S_{x})$ (66)
-$\leq\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t})$ (67)
-$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2(\frac{\sigma^{2}}{k_{i^{\prime}j,t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)}$ (68)
-$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{i^{\prime}j,t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)}$ (69)
-$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2\sigma^{2}\Big(\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)}$ (70)
-$=\exp{\Bigg(-\frac{2(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}\Bigg)}$ (71)
-$\leq\exp{\Bigg(-\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}\Bigg)$ (72)
-$=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}$ (73)
-Since the term $\frac{2(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}$ is greater than one, the inequality (72) follows.
-In addition, since we obtain the same result for $\mathbb{P}(E_{2})$ as $\mathbb{P}(E_{2})$, which leads to the same probability of keeping an action pair as in the proof of Theorem 2 and also the same results for Case 1, Case 3.
+$\mathbb{P}(E_{1}\cap E_{2})$ $\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}$ (79)
+$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{i^{\prime}j,t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{ij^{\prime},t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)}\Bigg\}$ (80)
+$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{i^{\prime}j,t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij^{\prime},t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)}\Bigg\}$ (81)
+$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\sigma^{2}\Big(\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\sigma^{2}\Big(\frac{(\hat{P}_{ij^{\prime},t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{ij^{\prime},t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)}\Bigg\}$ (82)
+$\leq\min\Bigg\{\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)},\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}\Bigg\}$ (83)
+$\leq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (84)
+Since the terms $\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}$ and $\frac{(\hat{P}_{ij^{\prime},t}+\hat{P}_{ij,t}+2)}{(1+\hat{P}_{ij^{\prime},t})(1+\hat{P}_{ij,t})}$ are obviously less than 2 since $0\leq\hat{P}_{ij,t}\leq 1$ for any $(i,j)\in S_{A}$, the inequality (83) follows.
+Then, it leads to the same probability of keeping an action pair as in the proof of Theorem 2 and also to the same results for Case 1, Case 3.
 On the other hand, since we explore each action pair non-uniformly, we need to update Case 2 by using
-$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t_{ij}}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$
+$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t_{ij}}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$ (85)
 since $\hat{P}_{ij,t_{ij}}\leq 1$ for all $i,j$. Then, the regret contribution becomes

### v4

@@
 Abstract
-We study a two-player zero-sum game in which the row player aims to maximize their payoff against an adversarial column player, under an unknown payoff matrix estimated through bandit feedback.
-We propose three algorithms based on the Explore-Then-Commit (ETC) framework.
+We study a two-player zero-sum game in which the row player aims to maximize their payoff against a competing column player, under an unknown payoff matrix estimated through bandit feedback.
+We propose three algorithms based on the Explore-Then-Commit (ETC) and action pair elimination frameworks.
 The first adapts it to zero-sum games, the second incorporates adaptive elimination that leverages the $\varepsilon$-Nash Equilibrium property to efficiently select the optimal action pair, and the third extends the elimination algorithm by employing non-uniform exploration.
-Our objective is to demonstrate the applicability of ETC in a zero-sum game setting by focusing on learning pure strategy Nash Equilibria.
+Our objective is to demonstrate the applicability of ETC and action pair elimination algorithms in a zero-sum game setting by focusing on learning pure strategy Nash Equilibria.
 A key contribution of our work is a derivation of instance-dependent upper bounds on the expected regret of our proposed algorithms, which has received limited attention in the literature on zero-sum games.
-Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC in zero-sum game setting and $O(\log(T\Delta^{2})/\Delta)$ for the adaptive elimination algorithm and its variant with non-uniform exploration, where $\Delta$ denotes the suboptimality gap.
-Therefore, our results indicate that ETC-based algorithms perform effectively in zero-sum game settings, achieving regret bounds comparable to existing methods while providing insight through instance-dependent analysis.
-Keywords zero-sum games, regret analysis, bandit feedback
+Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC in zero-sum game setting and $O\left(\frac{\log(T\Delta^{2})}{\Delta}\right)$ for the adaptive elimination algorithm and its variant with non-uniform exploration, where $\Delta$ denotes the suboptimality gap.
+Therefore, our results indicate that the ETC and action pair elimination algorithms perform effectively in zero-sum game settings, achieving regret bounds comparable to existing methods while providing insight through instance-dependent analysis.
 1 Introduction
-This paper focuses on finite two-player zero-sum games (TPZSGs), where two players interact in an adversarial setting with strictly opposing objectives, as first analyzed in [33, 34].
+This paper focuses on finite two-player zero-sum games (TPZSGs), where two players interact in a setting with strictly opposing objectives, as first analyzed in [33, 34].
 These games play a foundational role in game theory and online learning, capturing adversarial interactions in various applications such as economic competition and adversarial machine learning.
@@
 Studies on learning in games typically use either full information feedback where players observe the opponent’s actions or the entire payoff matrix after each round or partial feedback, also called bandit feedback, where each player observes only their own payoffs without access to the opponent’s strategy or the full matrix [7].
-In this paper, we study a TPZSG setting with bandit feedback, where players observe each other’s actions, by learning pure strategies for both players through balancing of exploration and exploitation.
+Zero-sum games with bandit feedback setting, where players also observe each other’s actions, were introduced in [29].
+In this paper, we study a TPZSG setting under bandit feedback with centralized control over both players by learning pure strategies through balancing of exploration and exploitation.
 Several existing studies have advanced our understanding of how players learn to play optimally in zero-sum games with bandit feedback.
-The variant of UCB algorithm for adversarial games under bandit feedback has been analyzed using worst-case regret bounds in [29].
-In particular, worst-case analyses lead to all games as equally hard by ignoring the specific structure of the game.
+The variant of UCB algorithm for zero-sum games under bandit feedback has been analyzed using worst-case regret bounds in [29].
+However, worst-case ignores the specific structure of the game.
 Instance-dependent analysis adapts the regret guarantees to the specific properties of a game to provide more practical performance guarantees.
@@
 Contributions.
-This paper investigates the effectiveness of the Explore-Then-Commit (ETC) algorithm and extensions in the context of TPZSGs.
-While ETC has been widely studied in standard multi-armed bandit (MAB) problems, there has been no prior work specifically analyzing its performance to zero-sum game settings, which makes this the first work to provide its detailed analysis in such a setting.
-A key motivation for studying ETC is its algorithmic simplicity, which allows for a clear structure.
-Moreover, our main focus is to show whether ETC and action elimination algorithms is applicable to zero-sum games by offering a refined understanding of regret behavior in TPZSGs under bandit feedback.
-In particular, we propose three algorithms that adapt the ETC approach to the TPZSG setting and present their instance-dependent regret analyses.
-The first one follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure Nash Equilibria (NE), achieves an upper bound of $O(\Delta+\sqrt{T})$ .
-The second algorithm incorporates an adaptive elimination (AE), similar to [6], by sequentially eliminating actions that, with high probability, are not part of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), gives us an upper regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$.
-Intuitively, this design is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action sets of the players.
-Finally, the last algorithm extends the adaptive elimination approach through non-uniform exploration, which also achieves a regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$.
-Consequently, despite the simplicity of the ETC framework, our results demonstrate that ETC-based approaches can be highly effective in adversarial settings.
-Paper structure.
-Section 2 reviews related work.
-Section 3 introduces preliminaries and formalizes a two-player zero-sum game setting.
-In Section 4, we present the ETC in a TPZSG setting and analyze its regret.
-Section 5 introduces an adaptive elimination algorithm and provides its regret analysis.
-In Section 6, we extend the elimination algorithm by using the non-uniform exploration approach and analyze its regret upper bound.
-Section 7 offers empirical results to validate their theoretical performances.
-Finally, Section 8 discusses our conclusions and future research directions.
+This paper investigates the effectiveness of the Explore-Then-Commit (ETC) and action elimination algorithms in TPZSGs.
+While ETC has been widely studied in standard multi-armed bandit problems, there has been no prior work specifically analyzing its performance to zero-sum game settings.
+Thus, our work is the first to provide a detailed analysis of ETC in this context.
+A key motivation for studying ETC is its algorithmic simplicity, which enables a clear structure.
+Our primary objective is to determine whether ETC and action elimination algorithms are applicable to zero-sum games by offering a refined understanding of regret behavior.
+In particular, we propose three algorithms that adapt the ETC and action pair elimination approaches to the TPZSG setting and provide their instance-dependent regret analyses.
+The first one follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure Nash Equilibria (NE), achieving a regret upper bound of $O(\Delta+\sqrt{T})$ (Section 4).
+The second algorithm incorporates adaptive elimination (AE), similar to [6], by sequentially eliminating actions that, with high probability, are not part of an $\varepsilon$-Nash Equilibrium; this yields an upper regret bound of $O\left(\frac{\log(T\Delta^{2})}{\Delta}\right)$ (Section 5).
+Intuitively, this approach is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action sets of the players.
+Finally, the third algorithm extends the adaptive elimination approach through non-uniform exploration, also achieving the same regret bound (see Section 6).
+Consequently, our results demonstrate that ETC and elimination approaches can be highly effective in zero-sum game settings.
 2 Related Work
-The multi-armed bandit (MAB) problem, has been introduced by [32] and originally described by [31], has been widely explored and has become a crucial framework for online decision-making under uncertainty, as explored in works such as [3, 4, 5, 6, 8, 9].
-The MAB problem models scenarios in which a player must repeatedly choose from a set of actions to maximize their own rewards while balancing exploration-exploitation tradeoff, has been formalized by [19] and further explored in studies such as [2, 21].
+The multi-armed bandit (MAB) problem was introduced by [32] and originally described by [31].
+It has since become a crucial framework for online decision-making under uncertainty, as explored in works such as [3, 4, 5, 6, 8, 9].
+The MAB problem models scenarios in which a player must repeatedly choose from a set of actions to maximize their own rewards while balancing exploration-exploitation tradeoff.
+This formalization was established by [19] and further investigated in studies such as [2, 21].
 A central challenge is identifying the action with the highest expected reward while minimizing suboptimal choices, which require efficient exploration and statistical confidence.
@@
 Due to its simplicity and effectiveness, variants of the ETC algorithm are widely used in various decision-making scenarios [17, 36].
-Zero-sum games represent a core concept in game theory.
-They describe competitive scenarios in which the gain of one player is exactly equal to the loss of the other.
+Zero-sum games represent a core concept in game theory, which describe competitive scenarios in which the gain of one player is exactly equal to the loss of the other.
 First highlighted by [33, 34], their theory and applications have been extensively explored [8, 12, 35].
@@
 [15] uses this property to efficiently approximate equilibrium in two-player games.
-In a zero-sum game setting with bandit feedback, players engage in repeated interactions without full knowledge of the payoff matrix, which makes it essential to estimate the unknown payoffs by previous observations.
+In a zero-sum game setting with bandit feedback, players engage in repeated interactions without knowledge of the true payoff matrix.
 Various approaches, such as UCB [29], Tsallis-INF [18] and a novel algorithm [24], have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret.
-Learning the game matrix is crucial for converging to equilibrium strategies and ensuring optimal decision-making in adversarial environments.
-Additionally, [23] and [22] investigate instance-dependent sample complexity bounds for zero-sum games under bandit feedback while [11] examines the convergence rates to NE.
-While [29] focuses on the worst-case regret analysis and [24] analyzes instance-dependent regret only for a $2\times 2$ game matrix, [18] provides an instance-dependent regret analysis which achieves an upper bound of $O(\log_{+}(T\Delta^{2})/\Delta)$ in the case of a pure strategy equilibrium.
-Compared to [18], our work introduces more refined notions of suboptimality gaps for game instances, whereas their regret analysis focuses solely on the gap between the NE and the action chosen by a player when the opponent plays the equilibrium action.
-Since their regret bound involves the square of the sum of suboptimality gaps, our instance-dependent analysis gives a tighter bound.
+Learning the game matrix is crucial for converging to equilibrium strategies and ensuring optimal decision-making in such environments.
+In addition, [23] and [22] investigate instance-dependent sample complexity bounds while [11] examines the convergence rates to NE.
+While [29] focuses on the worst-case regret analysis, achieving $O(\sqrt{T})$, and [24] analyzes instance-dependent regret only for a $2\times 2$ game matrix, [18] provides an instance-dependent analysis achieving an upper bound of $O\left(\left(\sum_{i\neq i^{*}}\frac{1}{\Delta_{i}}+\sum_{j\neq j^{*}}\frac{1}{\Delta^{\prime}_{j}}\right)\log T\right)$ in the case of pure equilibria where $\Delta$ and $\Delta^{\prime}$ denote some suboptimality gaps.
+Consequently, developing a comprehensive understanding of instance-dependent regret in TPZSGs under bandit feedback remains insufficiently addressed.
+In contrast to [18], where both players independently follow the Tsallis-INF algorithm, our setting is closely related to [29] and [24], but we assume control over both players.
+Our goal is to demonstrate the applicability of ETC and action elimination algorithms to zero-sum games and to achieve comparable regret bounds.
 3 Preliminaries
@@
 In a TPZSG, the goals of the players are strictly opposed, which implies that any gain by one player corresponds to an equal loss by the other.
-This framework provides a fundamental model for adversarial interactions where strategic decision-making under competition is central.
-The outcome of such a game is fully determined by the strategies chosen by both players and the goal of each player is to maximize their own payoff while minimizing that of their opponent.
+The outcome of such a game is fully determined by the strategies chosen by players and the goal of each player is to maximize their own payoff while minimizing that of their opponent.
 3.1 Two-Player Zero-Sum Games
 In a two-player zero-sum game, consisting of a row $(x)$ player and a column $(y)$ player, the row player aims to maximize their own payoff while the column player attempts to take an action to minimize this payoff.
-$S_{x}$ and $S_{y}$ are the action sets for the row player and column player, respectively.
-Thus, the action space of the game is $S_{A}=S_{x}\times S_{y}$, so that the action pairs $(i,j)\in S_{A}$ where $i\in S_{x},j\in S_{y}$.
-We also define $m=|S_{x}|$ as the number of actions for the row player and $l=|S_{y}|$ as the number of actions for the column player, while finally $N=ml=|S_{A}|$ is the total number of action pairs.
-The game itself is defined through a game matrix $A$, therefore if the players choose the pair $(i,j)$ then the row player obtains a payoff of $A(i,j)$ and the column player A(i,j)-A(i,j).
-The row player has a strategy (i.e. a probability distribution) $p$ to select an action $i\in S_{x}$, while the column player selects an action $j\in S_{y}$ according to a strategy $q$.
+Let $S_{x}$ and $S_{y}$ denote the action sets of the row and column players, respectively, with the joint action space $S_{A}=S_{x}\times S_{y}$.
+We define $m=|S_{x}|$ and $l=|S_{y}|$ as the numbers of actions available to the row and column players, and let $N=ml=|S_{A}|$ be the total number of action pairs.
+The game is specified by a payoff matrix $A$; therefore, if $(i,j)$ is played, the row player receives a payoff of $A(i,j)$ and the column player receives A(i,j)-A(i,j).
+Moreover, the row player has a strategy (i.e.
+a probability distribution) $p$ to select an action $i\in S_{x}$, while the column player selects an action $j\in S_{y}$ according to a strategy $q$.
 Definition 1.
 A pair of mixed strategies $(p^{*},q^{*})$ is a Nash Equilibrium if for all strategies $p$ and $q$, it satisfies
-$V^{*}$ $\geq{p}^{\top}Aq^{*},$ $V^{*}$ $\leq{p^{*}}^{\top}Aq$ (1)
-where $V^{*}={p^{*}}^{\top}Aq^{*}$ is the value of the game [26]. This means that $p^{*}$ and $q^{*}$ are optimal strategies for row and column players, respectively.
+${p}^{\top}Aq^{*}\leq{p^{*}}^{\top}Aq^{*}\leq{p^{*}}^{\top}Aq$ (1)
+where ${p^{*}}^{\top}Aq^{*}$ is the value of the game [26].
 There always exists an optimal mixed strategy for both players that guaranties the best possible outcome according to the Minimax Theorem [33, 34].
@@
 Definition 2 (Pure Nash Equilibria).
-A pair $(i^{*},j^{*})$ is a unique pure NE if:
-$i^{*}=\operatorname*{argmax}_{i\in S_{x}}\min_{j\in S_{y}}A(i,j)\quad\text{ and }\quad j^{*}=\operatorname*{argmin}_{j\in S_{y}}\max_{i\in S_{x}}A(i,j),$ (2)
-or equivalently
-$A(i,j^{*})\leq A(i^{*},j^{*})\leq A(i^{*},j),\qquad\forall i\in S_{x},j\in S_{y}.$ (3)
-The condition in (3) ensures that no player has an incentive to deviate to another action. We use $V^{*}=A(i^{*},j^{*})$ to denote the value of the game.
+A pair $(i^{*},j^{*})$ is a unique pure NE if $i^{*}=\operatorname*{argmax}_{i\in S_{x}}\min_{j\in S_{y}}A(i,j)$ and $j^{*}=\operatorname*{argmin}_{j\in S_{y}}\max_{i\in S_{x}}A(i,j)$ or equivalently
+$A(i,j^{*})\leq A(i^{*},j^{*})\leq A(i^{*},j),\quad\forall i\in S_{x},j\in S_{y}.$ (2)
+The condition in (2) ensures that no player has an incentive to deviate to another action.
 3.2 Games with Bandit Feedback
 The game proceeds in rounds.
-At time $t$, the row player draws an action $i_{t}\in S_{x}$ from a strategy $p_{t}$, and the column player draws $j_{t}\in S_{y}$ from a strategy $q_{t}$.
-The players then observe a noisy payoff $r_{t}$ with expected value $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, where the maximizing player receives $r_{t}$ and the minimizing player receives $r_{t}$.
+At time $t$, the row player draws an action $i_{t}\in S_{x}$ from a strategy $p_{t}$, and the column player draws $j_{t}\in S_{y}$ from $q_{t}$.
+The players then observe a noisy payoff $r_{t}$ with $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, where the maximizing player receives $r_{t}$ and the minimizing player receives $r_{t}$.
 We assume that both players observe each other’s action.
+[29] introduced this setting with algorithms for row player against any column player; we instead assume centralized control over both players.
 Since the players do not know $A$, we use the notation $\hat{A}(i,j)$ to denote the estimated payoff when the action pair $(i,j)$ is played.
-We express the action pair played in round $t$ by $(i_{t},j_{t})$.
 In each round $t$, the estimated game matrix is updated by computing the average payoff for each action pair $(i,j)$ as follows:
@@
 $\Delta_{ij}=\Delta_{ij}^{\text{max}}+\Delta_{ij}^{\text{min}}$ (6)
-$\Delta_{ij}^{*}=A(i^{*},j^{*})-A(i,j)=V^{*}-A(i,j)$ (8)
+$\Delta_{ij}^{*}=A(i^{*},j^{*})-A(i,j)$ (7)
 We note that $\Delta_{ij}^{\text{max}}\geq 0$ and $\Delta_{ij}^{\text{min}}\geq 0$, therefore $\Delta_{ij}\geq 0$ for any action pair $(i,j)$. However, $\Delta_{ij}^{*}$ might be negative.
-Using the NE property in (3), for all $i\in S_{x}$ and $j\in S_{y}$ we can write the following:
-$A(i^{*},j^{*})\leq A(i^{*},j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)$
-which implies
+Lemma 1.
+For all $i\in S_{x}$ and $j\in S_{y}$, $\Delta_{ij}^{*}\leq\Delta_{ij}$.
+Proof.
+Using the NE property in (2), for all $i\in S_{x}$ and $j\in S_{y}$ we can write $A(i^{*},j^{*})\leq A(i^{*},j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)$, which implies
 $A(i^{*},j^{*})-A(i,j)\leq\max_{i^{\prime}\in S_{x}}A(i^{\prime},j)-A(i,j).$ (8)
-Moreover,
-$A(i^{*},j^{*})\geq A(i,j^{*})\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})$
-implies the following:
+Similary, since $A(i^{*},j^{*})\geq A(i,j^{*})\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})$ we have the following:
 $A(i^{*},j^{*})-A(i,j)\geq\min_{j^{\prime}\in S_{y}}A(i,j^{\prime})-A(i,j).$ (9)
@@
 Therefore, we have $\Delta_{ij}^{\text{min}}\leq\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}$.
-On the other hand, since $\Delta_{ij}^{\text{min}}\geq 0$, we can write
-$\Delta_{ij}^{*}\leq\Delta_{ij}^{\text{max}}+\Delta_{ij}^{\text{min}}=\Delta_{ij}.$ (11)
+Since $\Delta_{ij}^{\text{min}}\geq 0$, it concludes the proof.
+∎
 3.4 Regret Notions
@@
 Let begin with the following formulations, which are the external regret of the max and min players, respectively:
-$R_{T}^{\text{max}}=\max_{i\in S_{x}}\mathbb{E}\Big[\sum_{t=1}^{T}A(i,j_{t})-A(i_{t},j_{t})\Big]$
-$R_{T}^{\text{min}}=\max_{j\in S_{y}}\mathbb{E}\Big[\sum_{t=1}^{T}A(i_{t},j_{t})-A(i_{t},j)\Big]$
+$R_{T}^{\text{max}}=\max_{i\in S_{x}}\mathbb{E}\left[\sum_{t=1}^{T}A(i,j_{t})-A(i_{t},j_{t})\right]$
+$R_{T}^{\text{min}}=\max_{j\in S_{y}}\mathbb{E}\left[\sum_{t=1}^{T}A(i_{t},j_{t})-A(i_{t},j)\right]$
 where $i_{t}$ and $j_{t}$ are the actions selected by the row and column player in round $t$, respectively.
@@
 Definition 3 (External regret).
-The (combined) external regret is given by the following:
+The (combined) external regret, which quantifies how much worse the players have performed compared to their best actions in hindsight, is given by the following:
 $R_{T}=R_{T}^{\text{max}}+R_{T}^{\text{min}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}\mathbb{E}[n_{ij,T}]$ (10)
-External regret quantifies how much worse the players have performed compared to their best actions in hindsight.
-An alternative regret notion is Nash regret, which measures the deviation of value from the Nash equilibrium:
+An alternative regret notion is Nash regret, which measures the value of the deviation from the Nash equilibrium:
 Definition 4 (Nash regret).
@@
 Algorithm 1 presents the implementation of ETC-TPZSG, where each player only observes the payoffs of the action pair they choose during the exploration period, then in the commit phase, the algorithm selects an optimal action pair which refers to the pure NE.
-Applying ETC in a zero-sum game setting presents some challenges compared to the standard MAB scenario.
-In bandit problems, after the exploration phase, a player simply commits to the arm with the highest estimated reward, which is typically effective if exploration is sufficient.
-However, in a zero-sum game with pure NE, the objective is not to identify an action with high reward but to find a possibly true NE.
-Inadequate exploration can lead to poor estimates of payoffs, making it difficult to correctly identify such an equilibrium.
-As a result, the design of the ETC approach in the zero-sum game setting requires careful balance of the exploration phase.
+Theorem 1.
+For $1\leq Nk\leq T$, the Nash regret of Algorithm 1, interacting with $\sigma$-subgaussian payoffs, is upper bounded by
+$R_{T}^{*}\leq k\sum_{i,j}\Delta_{ij}^{*}+(T-Nk)\sum_{i,j}\Delta_{ij}^{*}\exp{\left(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\right)}$ (12)
+where $k$ is the exploration time per action pair $(i,j)\in S_{A}$ and $N$ is the total number of action pairs.
+Proof Sketch.
+The total regret consists of two components: the regret from exploration and the regret from exploitation.
+During the exploration phase, each action pair is played $k$ times.
+In the exploitation phase, the algorithm plays according to the estimated pure equilibria.
+In other words, in the exploitation phase, we analyze the loss incurred from playing with a misidentified NE.
+Thus, the expected number of plays for any suboptimal action pair $(i,j)$ can be written by $\mathbb{E}[n_{ij,T}]=k+(T-Nk)\ \mathbb{P}((i,j)\text{ is identified as NE})$.
+Once this probability is bounded, the proof is complete.
+(See Appendix B).
+∎
 Algorithm 1 ETC-TPZSG
-1:Input: $S_{A},m,l,k,\sigma,T(1\leq mlk\leq T)$
+1:Input: $S_{A},m,l,k,\sigma,T$
 2:Initialize: $\hat{A}=[0]^{m\times l}$
@@
 7:for $t=mlk+1$ to $T$ do $\triangleright$ Commit Phase
-8:  Play an equilibrium $(i^{*},j^{*})$ satisfying the following property:
+8:  Play an equilibrium $(i^{*},j^{*})$ satisfying:
 $\hat{A}(i,j^{*})\leq\hat{A}(i^{*},j^{*})\leq\hat{A}(i^{*},j),\quad\forall i\in S_{x},j\in S_{y}$
 9:end for
-Theorem 1.
-The Nash regret of Algorithm 1, when interacting with $\sigma$-subgaussian payoffs and $1\leq Nk\leq T$, is upper bounded as follows:
-$R_{T}^{*}\leq k\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}+(T-Nk)\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}\exp{\Big(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big)}$ (14)
-where $k$ is the exploration time per action pair and $N$ is the total number of action pairs.
-Proof Sketch.
-The total regret consists of two components: the regret from exploration and the regret from exploitation.
-During the exploration phase, each action pair is played $k$ times.
-In the exploitation phase, the algorithm plays according to the pure equilibria, where we account for the probability that the pure NE is not correctly identified.
-In other words, in the exploitation phase, we analyze the loss incurred from playing with a misidentified NE.
-Thus, the expected number of plays for any suboptimal action pair $(i,j)$ can be written by the following:
-$\mathbb{E}[n_{ij,T}]=k+(T-Nk)\ \mathbb{P}((i,j)\text{ is identified as NE})$
-Once this probability is bounded, the proof is complete. A detailed derivation is provided in Appendix B. ∎
 As mentioned above, the regret comprises two main components, the exploration phase and the loss due to misidentifying the NE.
@@
 Then, we have $\Delta=A(i_{1},j)-A(i_{2},j)$ and we can write the regret simply as
-$R_{T}^{*}$ $\leq k\Delta+(T-2k)\Delta\exp{\Big(-\frac{k\Delta^{2}}{16\sigma^{2}}\Big)}$ (15)
-$\leq k\Delta+T\Delta\exp{\Big(-\frac{k\Delta^{2}}{16\sigma^{2}}\Big)}.$ (16)
+$R_{T}^{*}\leq k\Delta+T\Delta\exp{\left(-\frac{k\Delta^{2}}{16\sigma^{2}}\right)}.$ (13)
 Taking the first derivative with respect to $k$ and solving it, we get the following for the exploration time:
-$k=\max\Big\{1,\Big\lceil\frac{16\sigma^{2}}{\Delta^{2}}\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\rceil\Big\}$ (17)
+$k=\max\left\{1,\left\lceil\frac{16\sigma^{2}}{\Delta^{2}}\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\right\rceil\right\}$ (14)
 If $\Delta$ is known, we can easily find the necessary exploration time. Therefore, when we put it into the regret bound, it becomes
-$R_{T}^{*}\leq\min\Big\{T\Delta,\Delta+\frac{16\sigma^{2}}{\Delta}\Big(1+\max\Big\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\}\Big)\Big\}$ (18)
+$R_{T}^{*}\leq\min\left\{T\Delta,\Delta+\frac{16\sigma^{2}}{\Delta}\left(1+\max\left\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\right\}\right)\right\}$ (15)
 where the first term $T\Delta$ is the worst-case regret and for the other, we combine the regret we obtain in (13) with the value of the exploration time in (14).
-Then, focusing on $\max\Big\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\}$, we obtain $\Delta\geq\frac{4\sigma}{\sqrt{T}}$.
+Then, focusing on $\max\left\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\right\}$, we obtain $\Delta\geq\frac{4\sigma}{\sqrt{T}}$.
 Thus, when we put it in the regret bound (15), we have the following:
@@
 If $\Delta\leq 1$, it becomes $R_{T}^{*}\leq 1+c\sqrt{T}$, which refers to an instance-independent bound since it depends only on the time horizon $T$ not on the game instance.
-In the zero-sum game setting, our analysis shows that we get a regret bound similar to the standard bandit setting in [20], indicating that the ETC-based algorithm performs effectively in adversarial environments.
+Our analysis shows that we obtain a regret bound comparable to the standard bandit setting in [20], indicating that the ETC-based algorithm performs effectively even in the zero-sum game setting.
 5 ETC-TPZSG-AE Algorithm
@@
 Definition 5.
-The action pair $(i^{*},j^{*})$ is an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE) if the following condition holds:
-$A(i,j^{*})-\varepsilon\leq A(i^{*},j^{*})\leq A(i^{*},j)+\varepsilon,\quad\forall i\in S_{x},j\in S_{y}.$ (20)
+The action pair $(i^{*},j^{*})$ is an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE) if the following condition holds for any $\varepsilon>0$ and for all $i\in S_{x}$ and $j\in S_{y}$:
+$A(i,j^{*})-\varepsilon\leq A(i^{*},j^{*})\leq A(i^{*},j)+\varepsilon.$ (17)
 For the exploration time of the algorithm, let just go back to the regret bound in (12) and assume that there exist some $\Delta^{*}$ and $\Delta$ such that $\Delta^{*}\geq\Delta_{ij}^{*}$ and $\Delta\leq\Delta_{ij}$ for all $(i,j)\in S_{A}$.
@@
 Thus, as the exploration time, we can clearly use the following:
-$k_{t}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}\Big\rceil$ (21)
+$k_{t}=\left\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\left(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\right)}\right\rceil$ (18)
 In order to perform updates over rounds, an initial estimate is required.
@@
 For the property of $\varepsilon$-NE, we define the tolerance parameter as follows:
-$\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}$ (22)
+$\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{\left(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\right)}}$ (19)
 where $\varepsilon_{t}$ reflects how much players are allowed to deviate from the equilibrium and the tolerance over time gradually decreases to ensure convergence towards the NE.
-Substituting condition $k_{t}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}$ into the equation, where $k_{t}$ is defined in (21), we obtain $\varepsilon_{t}\leq\hat{\Delta}_{t}/2$.
+Substituting condition $k_{t}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\left(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\right)}$ into the equation, where $k_{t}$ is defined in (18), we obtain $\varepsilon_{t}\leq\hat{\Delta}_{t}/2$.
 This bound indicates that as the number of rounds $t$ increases, the tolerance $\varepsilon_{t}$ becomes smaller, forcing the system to narrow the set of action pairs by progressively eliminating those that do not approximate a NE.
@@
 By reducing the set of action pairs to include only the pairs that meet this criterion, it ensures that the remaining action pairs are the only ones that could potentially be near the NE; thus, it enables to focus on the most relevant action pairs.
-Consequently, we expect that the algorithm converges to an accurate approximation of the NE, even in the presence of estimation errors in $\hat{A}$.
+Consequently, we expect that the algorithm converges to an accurate approximation of the NE.
 Theorem 2.
-The Nash regret of Algorithm 2, interacting with $\sigma$-subgaussian payoffs, is upper bounded by
-$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\Big)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (23)
+For $1\leq ml\ll T$, the Nash regret of Algorithm 2, interacting with $\sigma$-subgaussian payoffs, is upper bounded by
+$R_{T}^{*}\leq$ $\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\left(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right)+\sum_{(i,j)\in S_{A_{2}}}\left(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\right)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (20)
 where $\lambda\geq 4\sigma\sqrt{e/T}$, and $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$ are two subsets of the action pairs.
 Proof Sketch.
-To analyze the regret of Algorithm 2, at first we define $t_{ij}=\min\{t:\hat{\Delta}_{t}<\Delta_{ij}/2\}$ as the earliest round satisfying $\hat{\Delta}_{t}<\Delta_{ij}/2$ and then we consider four distinct cases.
+We first define $t_{ij}=\min\{t:\hat{\Delta}_{t}<\Delta_{ij}/2\}$ as the earliest round satisfying $\hat{\Delta}_{t}<\Delta_{ij}/2$.
+Thus, we consider four distinct cases.
 First, we analyze the suboptimal action pairs which are not eliminated in round $t_{ij}$ or before.
@@
 The third case examines the scenario where the optimal action pair is eliminated in $t_{ij}$.
-Finally, we consider the case in which a suboptimal action pair remains as a unique one in the game (discussed in detail in Appendix C).
-∎
+Finally, we consider the case in which a suboptimal action pair remains as a unique one in the game.
+(See Appendix C) ∎
+Moreover, if we consider another regret approach in (10), we obtain the following regret bound.
+Theorem 3.
+The external regret incurred by Algorithm 2 when interacting with $\sigma$-subgaussian payoffs is bounded as following:
+$R_{T}\leq$ $\sum_{(i,j)\in S_{A_{1}}}\left(\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right)+\sum_{(i,j)\in S_{A_{2}}}\left(\frac{512\sigma^{2}}{\lambda}\right)+\lambda T.$ (21)
+Proof.
+Although the regret notion differs, this does not affect the underlying regret analysis. Thus, the previous theoretical bounds can be extended using Lemma 1, which completes the proof. ∎
 Algorithm 2 ETC-TPZSG-AE
-1:Input: $S_{A},m,l,\sigma,T(1\leq ml\ll T)$
+1:Input: $S_{A},m,l,\sigma,T$
 2:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$
 3:for $t=0,1,2,...,\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$ do
-4:  if $\mathbf{|S_{t}|>1}:$ then $\triangleright$ Exploration Phase
-5:   Explore each $(i,j)$ in $S_{t}$ $k_{t}$ times, defined in (21)
+4:  if $\mathbf{|S_{t}|>1}:$ then
+5:   Explore each $(i,j)$ in $S_{t}$ $k_{t}$ times in (18)
 6:   Update $\hat{A}$ using (3)
-[…]
+7:  else
 8:   Play with a unique action pair in $S_{t}$ until step $T$
 9:  end if
-10:  Remove the action pairs $(i,j)$ from $S_{t}$ NOT satisfying the following for $\varepsilon_{t}$ in (22): $\triangleright$ Elimination Phase
-$\hskip 35.56593pt\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
+10:  Remove all $(i,j)$ from $S_{t}$ NOT satisfying:
+[…]
 11:  $S_{t+1}\leftarrow S_{t}$
@@
 13:end for
-Moreover, if we consider another regret approach in (12), we obtain the following regret bound.
-Theorem 3.
-The external regret incurred by Algorithm 2 when interacting with $\sigma$-subgaussian payoffs is bounded as following:
-$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Big(\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}\Big)+\lambda T.$ (24)
-Proof.
-Although we have a different notion of regret, this does not affect the underlying structure of the regret analysis.
-This ensures that the previous theoretical bounds can be extended using $\Delta_{ij}^{*}\leq\Delta_{ij}$, then the result follows.
-∎
-We simplify the regret bound to $O(\frac{\log(T\Delta^{2})}{\Delta})$, as the logarithmic term dominates since by setting $\lambda=\sqrt{16\sigma^{2}e/T}$, we ensure that the term $\lambda T$ is bounded by $\sqrt{16\sigma^{2}eT}$, which is at most $\frac{e}{\Delta_{ij}}$ when $\Delta_{ij}\leq\lambda$.
+We simplify the regret bound to $O\left(\frac{\log(T\Delta^{2})}{\Delta}\right)$, as the logarithmic term dominates since by setting $\lambda=\sqrt{16\sigma^{2}e/T}$, we ensure that the term $\lambda T$ is bounded by $\sqrt{16\sigma^{2}eT}$, which is at most $\frac{e}{\Delta_{ij}}$ when $\Delta_{ij}\leq\lambda$.
 Our results are consistent with the regret bound derived by [6].
-Compared to our regret bound $O(\frac{\log(T\Delta^{2})}{\Delta})$ with the bound $O(\frac{\log_{+}(T\Delta^{2})}{\Delta})$ in [18], although they seem very similar, we observe that our study introduces a more refined notion of the suboptimality gap $\Delta$.
-Specifically, in [18], $\Delta$ refers to the gap between pure NE and the action chosen when the opponent plays the action in equilibrium.
-Moreover, our regret bound is tighter, as their formulation involves the square of the sum of suboptimality gaps, while ours uses the sum of squares of suboptimality gaps within the logarithmic factor.
-Moreover, our elimination algorithm specifically incorporates the extended pure NE property through $\varepsilon$, which also allows us to explain why we achieve a tighter regret bound for a TPZSG setting where both players have a pure strategy.
+Compared to our result, [18] achieves a regret bound of $O\left(\left(\sum_{i\neq i^{*}}\frac{1}{\Delta_{i}}+\sum_{j\neq j^{*}}\frac{1}{\Delta^{\prime}_{j}}\right)\log T\right)$ where $\Delta_{i}=A(i^{*},j^{*})-A(i,j^{*})$ and $\Delta^{\prime}_{j}=A(i^{*},j)-A(i^{*},j^{*})$ denote the gaps between pure NE and the payoff obtained by deviating while the opponent plays its equilibrium action.
+Although the bound is not directly comparable to ours, as the settings and the definition of suboptimality gaps differ, it currently presents the only instance-dependent regret analysis for general $m\times l$ zero-sum matrix games.
+Our regret bound is tighter and our analysis relies on a more refined notion of the suboptimality gap $\Delta$, as it exhibits a sharper dependence on the gap parameter and avoids the cumulative sum of suboptimality gaps that appears in [18].
+Furthermore, our elimination algorithm specifically incorporates the extended pure NE property through $\varepsilon$ in a centralized setting, which explains why we achieve a tighter regret bound for a TPZSG setting where both players have a pure strategy; this is further supported by the simulation results in Section 7.
 As a result, we have $R_{T}^{*}\leq R_{T}$, where $R_{T}^{*}$ evaluates its performance against the NE, capturing the cumulative loss relative to the equilibrium strategy, while the regret analysis of $R_{T}$ allows us to consider the losses of both players with respect to their best individual responses in each round, providing a broader and more flexible notion of regret.
-This outcome aligns with the interpretation that the expected regret against the NE establishes a more specific comparison, and thus, it naturally provides a tighter upper bound on the expected regret.
+This outcome aligns with the interpretation that the expected regret against the NE establishes a more specific comparison, and thus, it naturally provides a tighter upper bound.
 6 ETC-TPZSG-AE-NUE Algorithm
@@
 Thus, it is employed to reduce the number of plays for suboptimal action pairs more effectively.
-As mentioned in Section 3, the strategies of the row and column players are defined as probability distributions $p$ and $q$, respectively, over their corresponding action sets.
-In round $t$, using the estimated game matrix, we denote the estimated strategy of the row player by $\hat{p}_{t}$ and the estimated strategy of the column player by $\hat{q}_{t}$.
-Since the true game matrix is assumed to have the pure NE, we expect these estimated strategies to converge to the pure strategies over time.
+In round $t$, we denote the estimated optimal strategy of the row player by $\hat{p}_{t}$ and the estimated optimal strategy of the column player by $\hat{q}_{t}$, where they are calculated trough the estimated game matrix $\hat{A}$.
+We expect these estimated strategies to converge to the pure strategies over time.
 The estimated strategies of the players serve as a basis to determine the exploration time in Algorithm 3.
-The probability of selecting an action pair $(i,j)$ is estimated by $\hat{P}_{ij,t}=\hat{p}^{*}_{i,t}\hat{q}^{*}_{j,t}$, which satisfies $\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=1$, where $(\hat{p}_{t}^{*},\hat{q}_{t}^{*})$ is a mixed NE in round $t$.
-In other words, once optimal strategies are determined, we calculate the joint probability $\hat{P}_{ij,t}$, which represents the probability that the row player chooses action $i$ and the column player chooses action $j$ in round $t$.
+The joint probability of selecting an action pair $(i,j)$ is estimated by $\hat{P}_{ij,t}=\hat{p}_{i,t}\hat{q}_{j,t}$, which satisfies $\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=1$, where $(\hat{p}_{t},\hat{q}_{t})$ is a mixed NE in round $t$.
 Consequently, in Algorithm 3, the exploration time for each action pair $(i,j)$ is proportional to $\hat{P}_{ij,t}$, allowing action pairs with higher estimated probabilities to be played more frequently.
-We incorporate the strategies of the players $\hat{p}^{*}_{t}$ and $\hat{q}^{*}_{t}$ together with $k_{t}$.
-Specifically, in each round $t$, this involves estimating the optimal strategies for each player by solving a linear system using the estimated payoff matrix $\hat{A}$ as follows:
-$\begin{aligned} &\\ &\hat{P}_{ij,t}=\hat{p}^{*}_{i,t}\hat{q}^{*}_{j,t}\\ &\hat{p}^{*}_{t}=\operatorname*{argmax}_{\hat{p}}\min_{\hat{q}}\hat{p}^{T}\hat{A}\hat{q}\\ &\hat{q}^{*}_{t}=\operatorname*{argmin}_{\hat{q}}\max_{\hat{p}}\hat{p}^{T}\hat{A}\hat{q}\\ \end{aligned}\qquad\begin{aligned} &\text{Subject to:}\\ &\sum_{i\in S_{x}}\hat{p}_{i,t}=1,\quad\hat{p}_{i,t}\geq 0,\ \forall i\in S_{x}\\ &\sum_{j\in S_{y}}\hat{q}_{j,t}=1,\quad\hat{q}_{j,t}\geq 0,\ \forall j\in S_{y}\\ &\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=\sum_{i,j}\hat{p}_{i,t}\hat{q}_{j,t}=1\end{aligned}$ (25)
+We incorporate the strategies of the players $\hat{p}_{t}$ and $\hat{q}_{t}$ together with $k_{t}$ in (18).
+Specifically, in each round $t$, this involves estimating the optimal strategies for each player by solving a linear system using the estimated payoff matrix $\hat{A}$.
 Thus, the exploration time is given by the following:
-$k_{ij,t}=k_{t}+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t}}{\hat{\Delta}_{t}^{2}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}\Big\rceil$ (26)
-where $k_{t}$ is defined in (21).
+$k_{ij,t}=k_{t}+\left\lceil\frac{16\sigma^{2}\hat{P}_{ij,t}}{\hat{\Delta}_{t}^{2}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}\right\rceil$ (22)
 Thus, the algorithm explores each action pair at least $k_{t}$ times.
@@
 Algorithm 3 ETC-TPZSG-AE-NUE
-1:Input: $S_{A},m,l,\sigma,T(1\leq ml\ll T)$,
+1:Input: $S_{A},m,l,\sigma,T$
 2:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$
-3:for $t=0,1,2,...,\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$ do
-4:  if $\mathbf{|S_{t}|>1}:$ then $\triangleright$ Exploration Phase
-5:   Explore each $(i,j)$ in $S_{t}$ $k_{ij,t}$ times, defined in (26)
-6:   Update $\hat{A}$ using (4)
-[…]
-8:   Play with a unique action pair in $S_{t}$ until step $T$
-9:  end if
-10:  Remove the action pairs $(i,j)$ from $S_{t}$ NOT satisfying the following for $\varepsilon_{t}$ in (22): $\triangleright$ Elimination Phase
-$\hskip 35.56593pt\hat{A}(i^{\prime},j)-\varepsilon_{t}\leq\hat{A}(i,j)\leq\hat{A}(i,j^{\prime})+\varepsilon_{t},\forall i^{\prime},\forall j^{\prime}$
-11:  Solve the linear system in (25) to get $\hat{P}_{ij,t+1}$
-12:  $S_{t+1}\leftarrow S_{t}$
-13:  $\hat{\Delta}_{t+1}\leftarrow\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
-14:end for
+3:for $t=0,1,2,...,\left\lfloor\frac{1}{4}\log_{2}\frac{T}{e}\right\rfloor$ do
+4:  Run Algorithm 2 with exploration time $k_{ij,t}$ in (22)
+5:  Calculate optimal strategies $\hat{p}_{t}$ and $\hat{q}_{t}$ based on $\hat{A}$
+6:  Update $\hat{P}_{ij,t+1}$
+7:end for
 Theorem 4.
-The Nash regret of Algorithm 3 is upper bounded by
-$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\Big)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (27)
+For $1\leq ml\ll T$, the Nash regret of Algorithm 3 is upper bounded by
+$R_{T}^{*}\leq$ $\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\left(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right)+\sum_{(i,j)\in S_{A_{2}}}\left(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\right)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (23)
 where $\lambda\geq 4\sigma\sqrt[4]{e/T}$ and, $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$ are two subsets of the action pairs.
@@
 We investigate a TPZSG setting with bandit feedback, where the true payoff matrix is unknown and must be learned through player interactions.
-This setting is challenging, as players must estimate payoffs while making strategic decisions in an adversarial environment.
+This setting is challenging, as players must estimate payoffs while making strategic decisions with strictly opposing objectives.
 We adopt the ETC algorithm due to its simplicity, widespread use and lack of prior analysis in this context.
@@
 If $X$ is $\sigma$-subgaussian random variable, for any $\epsilon\geq 0$,
-$\mathbb{P}(X\geq\epsilon)\leq\exp{\Big(-\frac{\epsilon^{2}}{2\sigma^{2}}\Big)}.$ (31)
+$\mathbb{P}(X\geq\epsilon)\leq\exp{\left(-\frac{\epsilon^{2}}{2\sigma^{2}}\right)}.$ (27)
 Lemma 6 ([20, Lemma 5.4]).
@@
 $\quad\times\sqrt{\mathbb{P}(\hat{A}(i^{\prime},j_{1})-A(i^{\prime},j_{1})-(\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime}))\geq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})}$ (32)
-$\leq\sqrt{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\Big)}\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\Big)}}$ (37)
-$=\sqrt{\exp{\Big(-\frac{k[(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}]}{4\sigma^{2}}\Big)}}$ (38)
-$\leq\sqrt{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{8\sigma^{2}}\Big)}}$ (39)
-$=\sqrt{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}})^{2}}{8\sigma^{2}}\Big)}}$ (40)
-$=\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}})^{2}}{16\sigma^{2}}\Big)}$ (41)
+$\leq\sqrt{\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\right)}\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\right)}}$ (33)
+$=\sqrt{\exp{\left(-\frac{k[(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}]}{4\sigma^{2}}\right)}}$ (34)
+$\leq\sqrt{\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{8\sigma^{2}}\right)}}$ (35)
+$=\sqrt{\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}})^{2}}{8\sigma^{2}}\right)}}$ (36)
+$=\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}})^{2}}{16\sigma^{2}}\right)}$ (37)
 where we assume that $A(i_{1},j^{\prime})=\max_{i\in S_{x}}A(i,j^{\prime})$ and $A(i^{\prime},j_{1})=\min_{i\in S_{x}}A(i^{\prime},j)$.
 On the other hand, we note $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}=\max_{i}A(i,j^{\prime})-A(i^{\prime},j^{\prime})$, $\Delta_{i^{\prime}j^{\prime}}^{\text{min}}=A(i^{\prime},j^{\prime})-\min_{j}A(i^{\prime},j)$ and $\Delta_{i^{\prime}j^{\prime}}=\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}}$.
-To obtain (33), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, as stated in 3.
+To obtain (29), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, as stated in 4.
 For step (33), we apply Lemma 5 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable.
@@
 $\leq\min\{\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i_{1},j^{\prime})\geq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j_{1})-\hat{A}(i^{\prime},j^{\prime})\geq 0)\}$ (40)
-$\leq\min\Big\{\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\Big)},\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\Big)}\Big\}$ (45)
-$=\exp{\Big(-\frac{k\max\{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2},(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\}}{4\sigma^{2}}\Big)}$ (46)
-$\leq\exp{\Big(-\frac{k[(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}]}{8\sigma^{2}}\Big)}$ (47)
-$\leq\exp{\Big(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{16\sigma^{2}}\Big)}$ (48)
-$\leq\exp{\Big(-\frac{k\Delta_{i^{\prime}j^{\prime}}^{2}}{16\sigma^{2}}\Big)}$ (49)
+$\leq\min\left\{\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}}{4\sigma^{2}}\right)},\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{4\sigma^{2}}\right)}\right\}$ (41)
+$=\exp{\left(-\frac{k\max\{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2},(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\}}{4\sigma^{2}}\right)}$ (42)
+$\leq\exp{\left(-\frac{k[(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}]}{8\sigma^{2}}\right)}$ (43)
+$\leq\exp{\left(-\frac{k(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{16\sigma^{2}}\right)}$ (44)
+$\leq\exp{\left(-\frac{k\Delta_{i^{\prime}j^{\prime}}^{2}}{16\sigma^{2}}\right)}$ (45)
 where to get the inequality (43), we use Lemma 2.
@@
 Since we already have the probability of playing a misidentified NE from (37) or (45), we can write
-$\mathbb{E}[n_{ij,T}]\leq k+(T-Nk)\exp{\Big(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big)}.$ (51)
+$\mathbb{E}[n_{ij,T}]\leq k+(T-Nk)\exp{\left(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\right)}.$ (47)
 Hence, the proof is concluded. $\Box$
@@
 In this way, let define a threshold parameter as $\lambda\geq 4\sigma\sqrt{e/T}$, which gives us two subsets of actions pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$.
-Since $\Delta_{ij}^{*}\leq\Delta_{ij}$, it ensures that the bounds are preserved and it makes sense to use these action pair sets in the analysis.
+Since $\Delta_{ij}^{*}\leq\Delta_{ij}$ from Lemma 1, it ensures that the bounds are preserved and it makes sense to use these action pair sets in the analysis.
 Thus, we can write the Nash regret as
@@
 Furthermore, we have the following:
-$\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\Big(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\Big)}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{{t_{ij}}+1}<\frac{\Delta_{ij}}{4}$ (54)
-where the exploration time $k_{t}$ is defined in (21), which satisfies $k_{t_{ij}}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}$.
+$\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\left(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\right)}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{{t_{ij}}+1}<\frac{\Delta_{ij}}{4}$ (50)
+where the exploration time $k_{t}$ is defined in (18), which satisfies $k_{t_{ij}}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{\left(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\right)}$.
 For simplicity of notation, we write $\hat{A}$ instead of $\hat{A}_{t}$ for the estimated payoff matrix in round $t$ throughout this section.
@@
 $=\mathbb{P}(\hat{A}(i,j)-A(i,j)-(\hat{A}(i_{1},j)-A(i_{1},j))\geq\Delta_{ij}^{\text{max}}-\varepsilon_{t})$ (55)
-$\leq\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}$ (60)
+$\leq\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)}$ (56)
 where we assume $A(i_{1},j)=\max_{i\in S_{x}}A(i,j)$ and $A(i,j_{1})=\min_{j\in S_{y}}A(i,j)$.
@@
 $=\mathbb{P}(\hat{A}(i,j_{1})-A(i,j_{1})-(\hat{A}(i,j)-A(i,j))\geq\Delta_{ij}^{\text{min}}-\varepsilon_{t})$ (59)
-$\leq\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}$ (64)
+$\leq\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)}$ (60)
 where $A(i,j_{1})=\min_{j\in S_{y}}A(i,j)$.
@@
 Therefore, we write the probability of keeping an action pair $(i,j)$ as
-$\mathbb{P}(E_{1}\cap E_{2})\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}$ $\leq\min\Bigg\{\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)},\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}\Bigg\}$ (65)
-$=\exp{\Big(-\frac{k_{t}\max\{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2},(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}\}}{4\sigma^{2}}\Big)}$ (66)
-$\leq\exp{\Big(-\frac{k_{t}((\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}+(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2})}{8\sigma^{2}}\Big)}$ (67)
-$\leq\exp{\Big(-\frac{k_{t}(\Delta_{ij}-2\varepsilon_{t})^{2}}{16\sigma^{2}}\Big)}$ (68)
-$<\exp{\Big(-\frac{k_{t}\varepsilon_{t}^{2}}{4\sigma^{2}}\Big)}$ (69)
-$=\exp{\Bigg(-\frac{k_{t}\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{4\sigma^{2}}\Bigg)}$ (70)
+$\mathbb{P}(E_{1}\cap E_{2})\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}$ $\leq\min\left\{\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)},\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)}\right\}$ (61)
+$=\exp{\left(-\frac{k_{t}\max\{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2},(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}\}}{4\sigma^{2}}\right)}$ (62)
+$\leq\exp{\left(-\frac{k_{t}((\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}+(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2})}{8\sigma^{2}}\right)}$ (63)
+$\leq\exp{\left(-\frac{k_{t}(\Delta_{ij}-2\varepsilon_{t})^{2}}{16\sigma^{2}}\right)}$ (64)
+$<\exp{\left(-\frac{k_{t}\varepsilon_{t}^{2}}{4\sigma^{2}}\right)}$ (65)
+$=\exp{\left(-\frac{k_{t}\frac{4\sigma^{2}}{k_{t}}\ln{\left(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\right)}}{4\sigma^{2}}\right)}$ (66)
 $=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (67)
@@
 Thus, using (49), each action pair $(i,j)$ is played at most $k_{t_{ij}}$ times such that
-$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq\Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil.$ (73)
+$k_{t_{ij}}=\left\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\right\rceil\leq\left\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right\rceil.$ (69)
 Then, the regret contribution is expressed by
-$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$ $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)$ (74)
-$=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}$ (75)
+$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\left\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right\rceil$ $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\left(1+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right)$ (70)
+$=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}$ (71)
 Case 3.
@@
 Since the action pair selection persists up to time step $T$, resulting in an additional term in the regret bound.
-We note that $\Delta_{ij}^{*}\leq\Delta_{ij}$, which ensures that the regret does not grow excessively, and this relationship allows us to analyze the regret in terms of the defined action sets.
+Lemma 1 ensures that the regret does not grow excessively, and this relationship allows us to analyze the regret in terms of the defined action sets.
 As a result, if we sum these regret contributions as mentioned, we can conclude the proof.
@@
 $\mathbb{P}(E_{1}\cap E_{2})$ $\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}$ (75)
-$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{i^{\prime}j,t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{ij^{\prime},t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)}\Bigg\}$ (80)
-$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{i^{\prime}j,t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij^{\prime},t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)}\Bigg\}$ (81)
-$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\sigma^{2}\Big(\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\sigma^{2}\Big(\frac{(\hat{P}_{ij^{\prime},t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{ij^{\prime},t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)}\Bigg\}$ (82)
-$\leq\min\Bigg\{\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)},\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}\Bigg\}$ (83)
+$\leq\min\left\{\exp{\left(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{i^{\prime}j,t}}+\frac{\sigma^{2}}{k_{ij,t}})}\right)},\exp{\left(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{ij^{\prime},t}}+\frac{\sigma^{2}}{k_{ij,t}})}\right)}\right\}$ (76)
+$\leq\min\left\{\exp{\left(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\left(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{i^{\prime}j,t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\right)}\right)},\exp{\left(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\left(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij^{\prime},t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\right)}\right)}\right\}$ (77)
+$\leq\min\left\{\exp{\left(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\sigma^{2}\left(\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}\right)}\right)},\exp{\left(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\sigma^{2}\left(\frac{(\hat{P}_{ij^{\prime},t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{ij^{\prime},t})(1+\hat{P}_{ij,t})}\right)}\right)}\right\}$ (78)
+$\leq\min\left\{\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)},\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)}\right\}$ (79)
 $\leq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (80)
@@
 On the other hand, since we explore each action pair non-uniformly, we need to update Case 2 by using
-$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t_{ij}}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$ (85)
+$k_{t_{ij}}=\left\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{\left(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\right)}\right\rceil+\left\lceil\frac{16\sigma^{2}\hat{P}_{ij,t_{ij}}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{\left(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\right)}\right\rceil\leq 2\ \left\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right\rceil$ (81)
 since $\hat{P}_{ij,t_{ij}}\leq 1$ for all $i,j$. Then, the regret contribution becomes
-$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\ 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$ $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}$ (86)
+$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\ 2\ \left\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right\rceil$ $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}$ (82)
 Then, it concludes the proof.
@@
 The external regret of Algorithm 3 is bounded by
-$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{512\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}+\lambda T\Big).$
-Since we have $\Delta_{ij}^{*}\leq\Delta_{ij}$, it concludes the proof.
+$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{512\sigma^{2}}{\Delta_{ij}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}+\sum_{(i,j)\in S_{A_{2}}}\left(\frac{512\sigma^{2}}{\lambda}+\lambda T\right).$
+Since we have $\Delta_{ij}^{*}\leq\Delta_{ij}$ from Lemma 1, it concludes the proof.
 Appendix E Experimental Details